### v1

# Recovering Parameters from Edge Fluctuations: Beta-Ensembles and Critically-Spiked Models

Abstract.
Let $\Lambda=\{\Lambda_{0},\Lambda_{1},\Lambda_{2},\ldots\}$ be the point process that describes the edge scaling limit of either (i) ”regular” beta-ensembles with inverse temperature $\beta>0$, or (ii) the top eigenvalues of Wishart or Gaussian invariant random matrices perturbed by $r_{0}\geq 1$ critical spikes. In other words, $\Lambda$ is the eigenvalue point process of one of the scalar or multivariate stochastic Airy operators. We prove that a single observation of $\Lambda$ suffices to recover (almost surely) either (i) $\beta$ in the case of beta-ensembles, or (ii) $r_{0}$ in the case of critically-spiked models. Our proof relies on the recently-developed semigroup theory for the multivariate stochastic Airy operators.
Going beyond these parameter-recovery applications, our results also (iii) refine our understanding of the rigidity properties of $\Lambda$, and (iv) shed new light on the equality (in distribution) of stochastic Airy spectra with different dimensions and the same Robin boundary conditions.
1. Introduction
1.1. Main Result
Let $\Theta$ be the set of parameters of the form $\theta=(r,\beta,w)$, where $r\in\mathbb{N}$ is any positive integer and $\beta$ and $w$ are defined as follows:
  1. (1)
    If $r=1$, then $\beta>0$ and $w\in(-\infty,\infty]$, and
  2. (2)
    if $r>1$, then $\beta\in\{1,2,4\}$ and $w=(w_{1},\ldots,w_{r})\in(-\infty,\infty]^{r}$.
Given $\theta\in\Theta$, let $\mathcal{H}_{\theta}$ be the corresponding stochastic Airy operator (SAO). That is:
  1. (1)
    If $r=1$, then we let $W:[0,\infty)\to\mathbb{R}$ be a standard Brownian motion, and let
    (1.1) $\mathcal{H}_{\theta}=-\tfrac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+x+\tfrac{2}{\beta^{1/2}}W^{\prime}(x),$
    which acts on functions $f\in L^{2}\big{(}[0,\infty),\mathbb{R}\big{)}$ subject to the boundary condition
    $\begin{cases}f(0)=0&\text{if }w=\infty\text{ (Dirichlet)},\\ f^{\prime}(0)=wf(0)&\text{if }w\in\mathbb{R}\text{ (Robin)}.\end{cases}$
  2. (2)
    Let $\mathbb{F}_{1}=\mathbb{R}$, $\mathbb{F}_{2}=\mathbb{C}$, and $\mathbb{F}_{4}=\mathbb{H}$. If $r>1$ and $\beta\in\{1,2,4\}$, then let $W_{\beta}$ be a standard $r$-dimensional matrix Brownian motion on $\mathbb{F}_{\beta}$ (i.e., with GOE, GUE, or GSE increments depending on whether $\beta=1,2,4$; see [8, Page 2730]), and define
    (1.2) $\mathcal{H}_{\theta}=-\tfrac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+rx+\sqrt{2}W_{\beta}^{\prime}(x).$
    This acts on $r$-dimensional vector-valued functions $f=\big{(}f(i,\cdot):1\leq i\leq r\big{)}$ whose components $f(i,\cdot)\in L^{2}\big{(}[0,\infty),\mathbb{F}_{\beta}\big{)}$ are subject to the boundary conditions
    (1.3) $\begin{cases}f(i,0)=0&\text{if }w_{i}=\infty\text{ (Dirichlet)},\\ f^{\prime}(i,0)=w_{i}f(i,0)&\text{if }w_{i}\in\mathbb{R}\text{ (Robin)},\end{cases}\qquad 1\leq i\leq r.$
The SAOs were introduced by Edelman-Sutton [24], Ramírez-Rider-Virág [53], and Bloemendal-Virág [7, 8] to describe the edge fluctuations of various point processes of high interest in mathematical physics and statistics. That is, depending on the choice of $\theta$, the eigenvalues of $\mathcal{H}_{\theta}$ either describe the edge fluctuations of a wide class of beta-ensembles with so-called ”regular” external potential [4, 5, 11, 19, 44, 49, 53, 56], or the soft-edge fluctuations of critically-spiked Wishart or Gaussian invariant ensembles with real, complex, or quaternion entries [2, 7, 8, 45, 50, 59].
In this paper, we are interested in the following question regarding the SAOs:
Question 1.1.
Let $\theta\in\Theta$ be fixed, and let $\Lambda^{\theta}=\big{\{}\Lambda^{\theta}_{0},\Lambda^{\theta}_{1},\Lambda^{\theta}_{2},\ldots\big{\}}$ denote the eigenvalues of $\mathcal{H}_{\theta}$ arranged in increasing order. What information about $\theta$ can be recovered almost-surely from a single realization of $\Lambda^{\theta}$?
Given $(r,\beta,w)\in\Theta$, define the parameter
(1.4) $r_{0}=r_{0}(w)=\sum_{i=1}^{r}\mathbf{1}_{\{w_{i}<\infty\}};$
if $r=1$, then this reduces to $r_{0}=\mathbf{1}_{\{w<\infty\}}$. In words, $r_{0}$ is the number of Robin boundary conditions imposed on $\mathcal{H}_{\theta}$’s domain (hence $r-r_{0}$ is the number of Dirichlet conditions). Our main result asserts that if either one of $\beta$ or $r_{0}$ is known, then the other parameter can be recovered almost surely from a single realization of $\Lambda^{\theta}$:
Theorem 1.2.
There exists a deterministic function $\mathcal{T}$ (which can be written as an explicit limit; see (3.5) for the details) such that for every $\theta\in\Theta$, one has
(1.5) $\mathcal{T}(\Lambda^{\theta})=r_{0}+1/\beta\qquad\text{almost surely.}$.
1.2. Applications
Our main motivation for proving Theorem 1.2 consists of two parameter-recovery results:
  1. (i)
    Corollary 2.4: The temperature of any sequence of ”regular” beta-ensembles can be recovered from a single observation of their edge fluctuations.
  2. (ii)
    Corollary 2.14: The number of signals ”close enough” to the critical threshold in sequences of spiked Wishart and Gaussian invariant models can be recovered from a single observation of their top edge fluctuations.
Part of the significance of these results is that, to the best of our knowledge, all previously-known techniques to recover these parameters use the asymptotics of the full configuration of the point process or random matrix eigenvalues in question. In sharp contrast to this, Corollaries 2.4 and 2.14 show that the same can be done using only the vanishing proportion of particles/eigenvalues that contribute to edge scaling limits. See Sections 2.1 and 2.2 for more details and background, as well as open problems motivated by these results (Open Problems 2.5 and 2.15).
In addition to the above, we present two applications of a more theoretical nature:
  1. (iii)
    In Corollaries 2.20 and 2.22, we explain how Theorem 1.2 refines our understanding of the rigidity properties of $\Lambda^{\theta}$ (notably number rigidity).
  2. (iv)
    In Corollary 2.27, we explain how our proof technique sheds new light on the mysterious invariance of SAO eigenvalues with respect to adding components with Dirichlet boundary conditions.
We refer once again to Sections 2.3 and 2.4 for more details, background, and open problems motivated by these applications.
1.3. Proof Technique
Our approach to prove Theorem 1.2 is inspired by the theory of inverse problems in spectral geometry. That is, the observation that an impressive amount of information about a differential operator’s domain can be inferred from its eigenvalues alone. One of the most well-known manifestation of this phenomenon is the Weyl law [60] and its various extensions: If $\big{\{}\lambda_{k}(M,b):k\geq 0\big{\}}$ are the eigenvalues of the Laplacian on some compact manifold $M$ subject to some boundary condition $b$ (if $\partial M\neq\varnothing$), then $M$’s dimension, $M$’s volume, $M$’s boundary area, and the boundary condition $b$ (at least Dirichlet versus Neumann) can in many cases be determined by the asymptotics of the eigenvalue counting function
$\lambda\mapsto\big{|}\big{\{}k\geq 0:\lambda_{k}(M,b)\leq\lambda\big{\}}\big{|}.$
In this context, Question 2.1 can be viewed as the inverse spectral problem for the SAOs, as the components of $\theta=(r,\beta,w)$ characterize $\mathcal{H}_{\theta}$’s domain and noise. Thus, the proof of Theorem 1.2 relies on the observation that $r_{0}+1/\beta$ can be recovered from precise asymptotics of the $\Lambda_{k}^{\theta}$’s as $k\to\infty$. That being said, instead of recovering $r_{0}+1/\beta$ using the asymptotics of the eigenvalue counting function, we do so via the exponential traces $\mathrm{Tr}\big{[}\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big{]}=\sum_{k=0}^{\infty}\mathrm{e}^{-t\Lambda^{\theta}_{k}/2}$ for small $t>0$ (which is also widely used in spectral geometry; see, e.g., [38, Chapter 3] and references therein). More specifically, our proof of Theorem 1.2 relies on showing that
(1.6) $\mathrm{Tr}\big{[}\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big{]}=\sqrt{\frac{2}{\pi}}t^{-3/2}+\frac{1}{2}\left(r_{0}+\frac{1}{\beta}\right)-\frac{1}{4}+\mathfrak{o}_{t}\qquad\text{for all }t\in(0,1],$
where the random remainder terms $\mathfrak{o}_{t}$ are such that
(1.7) $\mathbf{E}[\mathfrak{o}_{t}]=o(1)\qquad\text{ as }t\to 0^{+},$
and such that there exists a constant $C>0$ satisfying
(1.8) $\mathbf{Cov}[\mathfrak{o}_{s},\mathfrak{o}_{t}]\leq C\left(\frac{\min\{s,t\}}{\max\{s,t\}}\right)^{1/4}\qquad\text{ for every }s,t\in(0,1].$
In particular, the fact that any sequence $\mathfrak{o}_{t_{1}},\mathfrak{o}_{t_{2}},\mathfrak{o}_{t_{3}},\ldots$ such that $t_{n}\to 0^{+}$ has a bounded variance and vanishing covariance means that $r_{0}+1/\beta$ can be recovered from (1.6) by the law of large numbers, that is, by averaging the traces $\mathrm{Tr}\big{[}\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big{]}$ over a sequence of $t$’s that vanishes fast enough; see (3.5) for details.
The main technical input in the proof of (1.6) consists of the recently-developed semigroup theory for the multivariate SAOs [29], i.e., the case $r>1$ (which is essential for our application of Theorem 1.2 to the spiked models in Corollary 2.14). This development extended prior results on the semigroup theory of the scalar SAOs [32, 39], i.e., the case $r=1$.
1.4. Questions of Optimality
Since Theorem 1.2 does not allow to recover the parameter $\theta$ completely, it is natural to ask if our result is optimal. For instance, one might ask if, in addition to $r_{0}$, it is possible to recover information about the magnitudes of the components $w_{i}$ in (1.3) that are finite. That said, it can be shown that this cannot be achieved with the method used in this paper.
More specifically, it is known that for some cases of $\theta$ (notably, $\theta=(1,2,\infty)$; see [31, Proposition 2.27]), the random remainder term $\mathfrak{o}_{t}$ in (1.6) satisfies
$\liminf_{t\to 0^{+}}\,\mathbf{Var}[\mathfrak{o}_{t}]>0.$
Given these nontrivial fluctuations, the deterministic terms in (1.6) represent the full extent of the information that can be inferred with 100% accuracy from the small $t$ asymptotics of $\mathrm{Tr}\big{[}\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big{]}$. In particular, if it is in fact possible to recover more information about $\theta$ from $\Lambda^{\theta}$, then we expect that a different approach than the one used in this paper would be required.
1.5. Organization
The remainder of this paper is organized as follows: In Section 2, we provide more details on the applications and open problems coming from Theorem 1.2. Then, in Section 3, we prove our main result.
1.6. Acknowledgements
The author thanks Promit Ghosal and Sumit Mukherjee for insightful conversations and references regarding the consistent estimation of inverse temperature in spin glasses and more general Gibbs point processes.
2. Applications and Open Problems
2.1. Temperature Recovery in Beta-Ensembles
2.1.1. Background
In statistical physics, the beta-ensembles model particle systems on $\mathbb{R}$ subjected to an external potential energy and pairwise logarithmic repulsion. More specifically, let $V:\mathbb{R}\to\mathbb{R}$ be a fixed function with sufficient growth, and for every $n\in\mathbb{N}$, define the $n$-particle Hamiltonian
(2.1) $H_{n}(x)=\frac{1}{2}\sum_{i=1}^{n}V(x_{i})-\frac{1}{n}\sum_{1\leq i<j\leq n}\log|x_{j}-x_{i}|,\qquad x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}.$
For every $\beta>0$ and $n\in\mathbb{N}$, we define the $n$-dimensional beta-ensemble with external potential $V$ and inverse temperature $\beta$ as the point process $\mathcal{B}^{\beta,n}=\big{\{}\mathcal{B}^{\beta,n}_{1},\ldots,\mathcal{B}^{\beta,n}_{n}\big{\}}$ sampled according to the Gibbs measure with density
(2.2) $\tfrac{1}{Z_{\beta,n}}\mathrm{e}^{-\beta nH_{n}(x)}~{}\mathrm{d}x,\qquad x\in\mathbb{R}^{n}.$
We adopt the convention that the points in $\mathcal{B}^{\beta,n}$ are in increasing order. The constant $Z_{\beta,n}>0$, called the partition function, ensures that (2.2) integrates to one.
The question of whether the temperature of infinite Gibbs point processes can be recovered from a single sample has attracted some attention in the literature; see, e.g., [20, Section 5.4] and [22] and references therein. In the case of a finite configuration, such as (2.2), this is of course not possible due to the mutual absolute continuity of the particles’ joint densities with different $\beta$’s. However, the problem becomes interesting if one asks about recovering $\beta$ as the number of particles grows to infinity. For instance, several prior works have studied the consistency of maximum likelihood and pseudolikelihood estimators (MLEs and MPLEs) of temperature in spin glasses in the large-dimensional limit; see, e.g., [6, 15, 17, 33] and references therein. In view of extending this to beta-ensembles, one may then ask:
Question 2.1.
Let $V:\mathbb{R}\to\mathbb{R}$ in (2.1) be fixed. Are there asymptotics (as $n\to\infty$) of the particles $\mathcal{B}^{\beta,n}_{1}\leq\cdots\leq\mathcal{B}^{\beta,n}_{n}$ from which $\beta$ can be recovered almost surely? If so, what is the smallest number of particles needed to recover $\beta$?
To the best of our knowledge, this has not been studied in the literature. That being said, one of the main hurdles in applying MLEs to Gibbs point processes consists of computing the partition function (e.g., [61, Section 1]). Therefore, given that $Z_{\beta,n}$ is very well understood for beta-ensembles—see [9, 10] and references therein—we expect that a similar strategy could be used to solve Question 2.1. To make this more concrete, consider the following heuristic, (which is informal as the asymptotics therein were calculated using Wolfram Mathematica; see Appendix A):
Heuristic 2.2 (Informal).
Consider the particle configuration’s total energy
(2.3) $H_{n}(\mathcal{B}^{\beta,n})=\frac{1}{2}\sum_{i=1}^{n}V(\mathcal{B}^{\beta,n}_{i})-\frac{1}{n}\sum_{1\leq i<j\leq n}\log\big{|}\mathcal{B}^{\beta,n}_{j}-\mathcal{B}^{\beta,n}_{i}\big{|}.$
If $V(x)=x^{2}/2$, then
(2.4) $\lim_{n\to\infty}-4\big{(}H_{n}(\mathcal{B}^{\beta,n})-\tfrac{3}{8}n+\tfrac{1}{2}\log n\big{)}-1=\log(\beta^{2}/4)-2\tfrac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}\qquad\text{in probability},$
where, $\Gamma(z)=\int_{0}^{\infty}t^{z-1}\mathrm{e}^{-t}~{}\mathrm{d}t$ denotes the Gamma function. Given that $\beta\mapsto\log(\beta^{2}/4)-2\tfrac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}$ is invertible for $\beta\in(0,\infty)$ (see Appendix A), we conclude that $\beta$ can, in principle, be recovered from the limit (2.4).
As argued in Appendix A, the limit (2.4) relies on an asymptotic analysis of the partition function. Thus, we expect that (2.4) could be extended to more general $V$’s using the asymptotic expansions of general beta-ensemble partition functions in [9, (1-3)] and [10, (1.14)].
2.1.2. New Result
A common feature of the statistics discussed above (i.e., MLEs, MPLEs, or the total energy in (2.3)) is that they use every point in the configuration $\mathcal{B}^{\beta,n}_{1}\leq\cdots\leq\mathcal{B}^{\beta,n}_{n}$ as an input. In sharp contrast to this, Theorem 1.2 implies that the temperature of many beta-ensembles can be recovered from a much more parsimonious asymptotic, which only depends on the vanishing proportion of particles that contribute to edge scaling limits (as the vast majority of particles instead contribute to the bulk scaling limits; e.g., [57, 58]). This is based on the edge universality of beta-ensembles, part of which can be informally summarized as follows:
Theorem 2.3 (Informal).
Let $V$ be a regular one-cut potential. That is, $V$ has an equilibrium measure $\mu_{V}$ supported on a compact interval $[\mathfrak{a},\mathfrak{b}]$, and $\mu_{V}$’s density vanishes like a square root at $\mathfrak{a}$ and $\mathfrak{b}$. Under additional technical conditions on $V$, for any edge $\mathfrak{e}\in\{\mathfrak{a},\mathfrak{b}\}$, there exists scaling constants $\mathfrak{s}_{n}=O(n^{2/3})$ such that
(2.5) $\lim_{n\to\infty}\mathfrak{s}_{n}\big{(}\mathfrak{e}-\mathcal{B}^{\beta,n}\big{)}=\Lambda^{(1,\beta,\infty)}\qquad\text{in distribution}.$
See [4, 11, 19, 44, 49, 53, 56] for the details, as well as [5] for a multi-cut extension of Theorem 2.3, whose statement we omit for simplicity. Given that $r_{0}=0$ when $\theta=(1,\beta,\infty)$, we thus obtain the following corollary from Theorem 1.2:
Corollary 2.4.
Suppose that (2.5) holds, and let $\mathcal{T}$ be the deterministic function in (1.5)/(3.5). For every $\beta>0$, one has
$\mathcal{T}\left(\lim_{n\to\infty}\mathfrak{s}_{n}\big{(}\mathfrak{e}-\mathcal{B}^{\beta,n}\big{)}\right)=1/\beta\qquad\text{almost surely},$
where the limit above is in distribution.
Similarly to Theorem 2.3, there is an obvious multi-cut extension to Corollary 2.4, which we omit for simplicity.
2.1.3. Open Problem
Following-up on Corollary 2.4, it is natural to wonder if it is possible to construct consistent estimators of $\beta$ that (1) only depend on $\mathcal{B}^{\beta,n}$’s edge fluctuations and (2) can be applied to finite-dimensional models. More specifically:
Open Problem 2.5.
Let $\mathfrak{e}$ and $\mathfrak{s}_{n}$ be as in (2.5). Does there exist a sequence of measurable functions $\{\mathcal{T}_{n}:n\in\mathbb{N}\}$ such that, as $n\to\infty$,
  1. (1)
    $\mathcal{T}_{n}$ only depends on the particles $\mathcal{B}^{\beta,n}_{k}\in\mathcal{B}^{\beta,n}$ such that $|\mathcal{B}^{\beta,n}_{k}-\mathfrak{e}|=o(1)$; and
  2. (2)
    $\mathcal{T}_{n}(\mathcal{B}^{\beta,n})\to 1/\beta$ almost surely?
If so, can one also characterize the fluctuations of $\mathcal{T}_{n}(\mathcal{B}^{\beta,n})-1/\beta$ as $n\to\infty$?
Thanks to Corollary 2.4 and (3.5), we expect that a positive resolution to Open Problem 2.5 could potentially rely on establishing a statement of the form
(2.6) $\lim_{n\to\infty}\left|\sum_{k\geq 0\text{ s.t. }|\mathcal{B}^{\beta,n}_{k}-\mathfrak{e}|\leq\varepsilon_{n}}\mathrm{e}^{-t_{n}\,\mathfrak{s}_{n}(\mathfrak{e}-\mathcal{B}^{\beta,n}_{k})/2}-\sum_{k=0}^{\infty}\mathrm{e}^{-t_{n}\Lambda^{(1,\beta,\infty)}_{k}/2}\right|=0$
for $\varepsilon_{n},t_{n}=o(1)$. A similar result already exists in the special case where $V(x)=x^{2}/2$ when $t_{n}$ is replaced by a fixed $t>0$; see the proof of [39, Corollary 2.10].
2.2. Detecting Critical Signals in Spiked Models
2.2.1. Background
In the early 2000’s, Johnstone introduced the spiked Wishart ensemble [42, Pages 301–304] as a means of modeling the effect of low-rank signals on the spectral statistics of covariance estimators in high dimension. Some years later, Péché [50] introduced a closely related additive spiked ensemble to study low-rank perturbations of self-adjoint matrices. In their simplest incarnation (i.e., rotationally-symmetric noise), these models are defined as follows: Let $r\in\mathbb{N}$, $\beta=\{1,2,4\}$, and $\ell=(\ell_{r-1},\ldots,\ell_{0})\in[0,\infty)^{r}$ be fixed, assuming that $\ell_{r-1}\leq\cdots\leq\ell_{0}$.
  1. (1)
    For any integers $n\geq 0$ and $p\geq r$, let $D_{\beta,n,p}$ be a $n\times p$ matrix whose entries are i.i.d. standard Gaussians in $\mathbb{F}_{\beta}$, and let $\Sigma_{\ell,p}$ be the $p\times p$ diagonal matrix with eigenvalues $1,\ldots,1,1+\ell_{r-1},\ldots,1+\ell_{0}$. We define the $(n,p)$-dimensional spiked Wishart ensemble as follows ($\cdot^{*}$ denotes the conjugate transpose):
    $W_{\beta,\ell,n,p}=\tfrac{1}{n}D_{\beta,n,p}\Sigma_{\ell,p}D_{\beta,n,p}^{*}.$
  2. (2)
    For any integer $n\geq r,$ let $X_{\beta,n}=(D_{\beta,n}^{*}+D_{\beta,n})/\sqrt{2}$, where $D_{\beta,n}$ is an $n\times n$ matrix whose entries are i.i.d. standard Gaussians in $\mathbb{F}_{\beta}$ (in other words, $X_{\beta,n}$ is a GOE, GUE, or GSE matrix, depending on whether $\beta=1$, $2$, or $4$). Then, let $P_{\ell,n}$ be the $n\times n$ diagonal matrix with eigenvalues $0,\ldots,0,\ell_{r-1},\ldots,\ell_{0}$. We define the $n$-dimensional spiked Gaussian invariant ensemble as
    $Y_{\beta,\ell,n}=\tfrac{1}{\sqrt{n}}X_{\beta,n}+P_{\ell,n}.$
The main problem involving these models consists of detecting the signal $\ell$ using the noisy observations provided by the spectra of $W_{\beta,\ell,n,p}$ or $Y_{\beta,\ell,n}$ in the high-dimensional setting $n\asymp p$. As a first step toward this goal, one may consider the asymptotic version of the question:
Question 2.6.
Let $r\in\mathbb{N}$ and $\beta=\{1,2,4\}$ be fixed. In the case of the Wishart ensemble, we assume that $p=p(n)$ depends on $n$ in such a way that
(2.7) $p(n)=\gamma\big{(}1+o(1)\big{)}\qquad\text{as }n\to\infty$
for some $\gamma\in(0,\infty)$. Let $\ell=\ell(n)\in[0,\infty)^{r}$ be either fixed or dependent on $n$, and let $\mathcal{S}^{\beta,\ell,n}=\big{\{}\mathcal{S}^{\beta,\ell,n}_{1},\ldots,\mathcal{S}^{\beta,\ell,n}_{n}\big{\}}$ be the eigenvalues of either $W_{\beta,\ell,n,p(n)}$ or $Y_{\beta,\ell,n}$, in increasing order. Are there asymptotics of $\mathcal{S}^{\beta,\ell,n}$ (as $n\to\infty$) from which any information about the signal $\ell$ can be recovered almost surely?
The first papers in this direction were by Baik-Ben Arous-Péché [2] and Péché [50]. Among other things, these results showed that the almost-sure asymptotic detectability of each signal $\ell_{i}$ depends on its size relative to the threshold
$\tau=\begin{cases}\sqrt{\gamma}&\text{(Wishart)},\\ 1&\text{(Gaussian invariant)}.\end{cases}$
That is, the analysis of Question 2.6 is typically split into three distinct cases:
  1. (1)
    The supercritical case, where $\ell_{i}>\tau$ for at least one $10\leq i\leq r-$ 1;
  2. (2)
    the subcritical case, where $\ell_{r-1}\leq\cdots\leq\ell_{0}<\tau$; and
  3. (3)
    the critical case, where $\ell_{r_{\mathrm{c}}-1},\ldots,\ell_{0}=\tau+o(1)$ for some $r_{\mathrm{c}}\geq 1$.
To the best of our knowledge, the current state of the art for this theory can be briefly summarized as follows. Firstly, we state a result for the supercritical case:
Theorem 2.7 ([2, 3, 14, 50]).
Let $\mathfrak{e}$ be the top edge of the limiting empirical spectral distribution (ESD) of the unperturbed Wishart or Gaussian invariant model, i.e.,
(2.8) $\mathfrak{e}=\begin{cases}(1+\sqrt{\gamma})^{2}&\text{(Wishart)},\\ 2&\text{(Gaussian invariant)}.\end{cases}$
Define the function
$\mathcal{O}(x)=\begin{cases}(1+x)(1+\gamma/x)&\text{(Wishart)},\\ x+1/x&\text{(Gaussian invariant)}.\end{cases}$
For any fixed $\ell\in[0,\infty)^{r}$ and $i\geq 0$, one has
(2.9) $\lim_{n\to\infty}\mathcal{S}^{\ell,\beta,n}_{n-i}=\begin{cases}\mathcal{O}(\ell_{i})&\text{if $i\leq r-1$ and $\ell_{i}>\tau$},\\ \mathfrak{e}&\text{if $i\leq r-1$ and $\ell_{i}\leq\tau$, or if $i>r-1$}\end{cases}\qquad\text{almost surely}.$
Remark 2.8.
$\mathcal{O}(x)>\mathfrak{e}$ is strictly increasing for $x\in(\tau,\infty)$. Thus, (2.9) implies that the number and magnitude of all supercritical signals $\ell_{i}>\tau$ can be recovered from the asymptotics of a finite number of the top eigenvalues of $W_{\beta,\ell,n,p}$ and $Y_{\beta,\ell,n}$.
Secondly, we state a result regarding the impossibility of detecting subcritical signals almost-surely using the eigenvalues $\mathcal{S}^{\beta,\ell,n}$ alone:
Theorem 2.9 ([41, 46, 47, 48]).
Let $\beta=1$ (i.e., $\mathbb{F}_{\beta}=\mathbb{R}$). For any fixed $\ell\in[0,\infty)^{r}$ such that $\ell_{r-1}\leq\cdots\leq\ell_{0}<\tau$, the sequence of spiked eigenvalues $(\mathcal{S}^{\beta,\ell,n}:n\geq 1)$ is contiguous with respect to the sequence of unperturbed eigenvalues $(\mathcal{S}^{\beta,0,n}:n\geq 1)$. That is, for any sequence of measurable maps $\{\mathcal{T}_{n}:n\geq 1\}$,
$\lim_{n\to\infty}\mathbf{P}\big{[}\mathcal{T}_{n}(\mathcal{S}^{\beta,0,n})=0\big{]}=1\qquad\text{implies that}\qquad\lim_{n\to\infty}\mathbf{P}\big{[}\mathcal{T}_{n}(\mathcal{S}^{\beta,\ell,n})=0\big{]}=1.$
Remark 2.10.
Despite the impossibility result in Theorem 2.9, one could ask if there nevertheless exist tests that distinguish between the hypotheses
(2.10) $\boldsymbol{H}_{0}\text{ : "no perturbation"}\qquad\text{versus}\qquad\boldsymbol{H}_{\ell}\text{ : "perturbation $=\ell$"}$ ”
with nontrivial statistical power when $\ell\neq 0$ is subcritical. The short answer to this question is yes; see, e.g., [25, 41, 43, 47, 48] and references therein for a non-exhaustive sample of this very sizeable literature. A unifying feature of all these results is that the statistical tests that achieve optimal or close-to-optimal power in distinguishing (2.10) (e.g., likelihood ratios or linear spectral statistics) all use the full spectrum $\mathcal{S}^{\beta,\ell,n}_{1}\leq\cdots\leq\mathcal{S}^{\beta,\ell,n}_{n}$ as an input. In the words of Dobriban [23]:
All eigenvalues matter to achieve sharp detection of weak principal components in high-dimensional data.
To give a simple example of how such results look like, consider the following:
Theorem 2.11 ([25, 41]).
Suppose that $\beta=1$, and that we consider the Gaussian invariant model $Y_{\beta,\ell,n}$ (hence $\tau=1$). Let $\ell\in[0,\infty)^{r}$ be such that $\ell_{r-1}=\cdots=\ell_{0}=\lambda$ for some fixed $\lambda\in[0,1)$. Let $\mathcal{L}_{n}$ denote the output of the likelihood ratio test (LRT) for (2.10) in the $n$-dimensional model, where $\mathcal{L}_{n}(\mathcal{S}^{\beta,\ell,n})=0$ means accepting $\boldsymbol{H}_{0}$ and $\mathcal{L}_{n}(\mathcal{S}^{\beta,\ell,n})=1$ means rejecting $\boldsymbol{H}_{0}$. Define the error function
(2.11) $\mathcal{E}_{n}(\lambda)=\mathbf{P}\big{[}\mathcal{L}_{n}(\mathcal{S}^{\beta,\ell,n})=1~{}\big{|}~{}\boldsymbol{H}_{0}\text{ is true}\big{]}+\mathbf{P}\big{[}\mathcal{L}_{n}(\mathcal{S}^{\beta,\ell,n})=0~{}\big{|}~{}\boldsymbol{H}_{\ell}\text{ is true}\big{]}.$
The LRT minimizes (2.11) over all measurable functions of $\mathcal{S}^{\beta,\ell,n}$. Moreover,
(2.12) $\lim_{n\to\infty}\mathcal{E}_{n}(\ell)=\mathrm{erfc}\left(\tfrac{r}{4}\sqrt{-\log(1-\lambda)}\right)\in(0,1),$
where we use $\mathrm{erfc}(z)=\frac{2}{\sqrt{\pi}}\int_{z}^{\infty}\mathrm{e}^{-t^{2}}~{}\mathrm{d}t$ to denote the complementary error function.
Thirdly, we did not find any works in the literature that directly address Question 2.6 in the critical case $\ell_{r_{\mathrm{c}}-1},\ldots,\ell_{0}=\tau+o(1)$. However, as pointed out in [51, Appendix A], consistent estimators of $\ell$ in this case can be proved to exist by examining the asymptotic power of subcritical tests when $\ell_{i}\to\tau^{-}$ (e.g., (2.12) goes to zero when $\ell_{i}=\lambda\to 1^{-}$). Thus, the answer to Question 2.6 is positive in the critical case. However, to the best of our knowledge, since this conclusion relies on the optimal subcritical estimators, it follows from Remark 2.10 that all eigenvalues are used in the currently-known methods to consistently recover critical signals.
2.2.2. New Result
Remarks 2.8 and 2.10 highlight a very sharp dichotomy between the number of eigenvalues in $\mathcal{S}^{\beta,\ell,n}$ that are required to optimally infer supercritical and subcritical signals (respectively, a finite number of top eigenvalues versus all eigenvalues). Given that critical signals interpolate between these two extremes, it is thus natural to ask:
Question 2.12.
In the context of Question 2.6, what is the minimal number of eigenvalues that are needed in order to detect critical signals with probability one?
According to (2.9), it is impossible to detect critical signals using any finite collection of top eigenvalues. In this view, the main result of this paper implies that critical signals ”close enough” (see (2.15)) to the threshold $\tau$ can be recovered from the vanishing proportion of top eigenvalues that contribute to edge scaling limits. This is based on the following:
Theorem 2.13 ([7, 8]; see also [2, 45, 50, 59]).
Let $r\in\mathbb{N}$ and $\beta=\{1,2,4\}$ be fixed, let $\gamma$ and $p=p(n)$ be as in (2.7) in the Wishart case, and suppose that $\ell=\ell(n)\in[0,\infty)^{r}$ is either fixed or depends on $n$. Suppose that for every $1\leq i\leq r$, the limit
(2.13) $w_{i}=\begin{cases}\displaystyle\tfrac{1}{\sqrt{\gamma}}\left(1+\tfrac{1}{\sqrt{\gamma}}\right)^{-2/3}\lim_{n\to\infty}n^{1/3}\big{(}\sqrt{\gamma}-\ell_{i-1}\big{)}&\text{(Wishart)},\\ \displaystyle\lim_{n\to\infty}n^{1/3}(1-\ell_{i-1})&\text{(Gaussian invariant)},\end{cases}$
exists and is an element of (,]($\infty$,$\infty$]( - ∞ , ∞ ]; then let $w=(w_{1},\ldots,w_{r})$. If we define
$\mathfrak{s}_{n}=\begin{cases}\tfrac{1}{\sqrt{\gamma}}\left(1+\tfrac{1}{\sqrt{\gamma}}\right)^{-4/3}n^{2/3}&\text{(Wishart)},\\ n^{2/3}&\text{(Gaussian invariant)},\end{cases}$
and we let $\mathfrak{e}$ be the top edge of the limiting ESD defined in (2.8), then it holds that
(2.14) $\lim_{n\to\infty}\mathfrak{s}_{n}\big{(}\mathfrak{e}-\mathcal{S}^{\beta,\ell,n}\big{)}=\Lambda^{(r,\beta,w)}\qquad\text{in distribution}.$
Thanks to (2.13), we see that $w_{i}\in\mathbb{R}$ corresponds to
(2.15) $\ell_{i-1}=\tau+O(n^{-1/3}),$
whereas $w_{i}=\infty$ corresponds to $\ell_{i-1}$ being smaller than $\tau$ by a difference of greater order than $n^{-1/3}$. Thus, in the present context, the parameter $r_{0}$ introduced in (1.4) counts the number of critical signals that are no farther than $O(n^{-1/3})$ from the threshold $\tau$. Given that one would presumably know whether $\beta=1,2,4$ if presented with a realization of the random matrices $W_{\beta,\ell,n,p}$ or $Y_{\beta,\ell,n}$, Theorem 1.2 then implies that the number of critical signals counted by $r_{0}$ can be determined from the top edge fluctuations of the random matrices:
Corollary 2.14.
Let $\mathcal{T}$ be as in (1.5)/(3.5). Under the assumptions of Theorem 2.13,
$\mathcal{T}\left(\lim_{n\to\infty}\mathfrak{s}_{n}\big{(}\mathfrak{e}-\mathcal{S}^{\beta,\ell,n}\big{)}\right)-1/\beta=r_{0}\qquad\text{almost surely},$
where the limit above is in distribution.
2.2.3. Open Problem
Similarly to our result on beta-ensembles, the natural follow-up to Corollary 2.14 is the following:
Open Problem 2.15.
Let $\mathfrak{e}$ and $\mathfrak{s}_{n}$ be as in (2.14). Does there exist a sequence of measurable functions $\{\mathcal{T}_{n}:n\in\mathbb{N}\}$ such that, as $n\to\infty$,
  1. (1)
    $\mathcal{T}_{n}$ only depends on the particles $\mathcal{S}^{\beta,\ell,n}_{k}\in\mathcal{S}^{\beta,\ell,n}$ such that $|\mathcal{S}^{\beta,\ell,n}_{k}-\mathfrak{e}|=o(1)$; and
  2. (2)
    $\mathcal{T}_{n}(\mathcal{S}^{\beta,\ell,n})\to r_{0}$ almost surely?
If so, can one also characterize the fluctuations of $\mathcal{T}_{n}(\mathcal{S}^{\beta,\ell,n})-r_{0}$ as $n\to\infty$?
That being said, the convergence result analogous to (2.6) that one would presumably need to prove Open Problem 2.15 using (1.5)/(3.5) is much more difficult in the case of spiked models. To illustrate this, we note that the limit
$\lim_{n\to\infty}\sum_{k=1}^{n}\mathrm{e}^{-t\,\mathfrak{s}_{n}(\mathfrak{e}-\mathcal{S}^{\beta,\ell,n})/2}=\sum_{k=0}^{\infty}\mathrm{e}^{-t\Lambda^{(r,\beta,w)}/2}$
for fixed $t>0$ has not been proved for any spiked model when $r_{0}>0$, even in the simplest case of the scalar SAO (i.e., $r=r_{0}=1$). Indeed, using current techniques, such a result would first require proving a seemingly difficult strong invariance principle for reflected random walks; see [27, Conjectures 2.23 and 6.11].
2.3. Rigidity
2.3.1. Background
In the last 15 years, a sizeable literature on ”rigidity” properties of point processes has been developing, starting with pioneering works by Ghosh and Peres [34, 37]. We refer to [18, 35] and references therein for a survey. In this literature, rigidity refers to the observation that for certain strongly-correlated point processes, some features of the point configuration inside a bounded set is almost-surely determined by the configuration of points outside that set. For instance:
Definition 2.16.
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a point process. We say that $\mathcal{X}$ is number rigid if for every bounded Borel set $B\subset\mathbb{R}$, there exists a measurable map $\mathcal{M}_{B}$ such that
$|\mathcal{X}\cap B|=\mathcal{M}_{B}(\mathcal{X}\cap B^{c})\qquad\text{almost surely}.$
(In words, the points outside $B$ determine the number of points inside $B$.) More generally, given any integer $k\geq 0$, we say that $\mathcal{X}$ is $k$-moment rigid if for every bounded Borel set $B\subset\mathbb{R}$, there exists a measurable map $\mathcal{M}_{B}$ such that
$\sum_{x\in\mathcal{X}\cap B}x^{k}=\mathcal{M}_{B}(\mathcal{X}\cap B^{c})\qquad\text{almost surely}.$
(In words, the points outside $B$ determine the $k^{\mathrm{th}}$ moment of points inside $B$.)
Examples of nontrivial rigid point processes include the following:
Example 2.17.
The Ginibre point process was shown to be number rigid in [37].
Example 2.18.
The $\alpha$-Gaussian analytic function ($\alpha$-GAF) zero point process (i.e., the zeros of the random entire function $z\mapsto\sum_{k\geq 0}\frac{g_{k}}{(k!)^{\alpha/2}}z^{k}$, where $g_{k}$ are i.i.d. complex Gaussians) were shown to be $k$-moment rigid for all $0\leq k\leq\lfloor 1/\alpha\rfloor$ in [36, 37].
Remark 2.19.
Rigidity was in part introduced to help distinguish structured point processes that are otherwise difficult to tell apart. As stated in [37, Page 1794]:
While a simple visual inspection suffices to (heuristically) distinguish a sample of the Poisson process from that of either the Ginibre or the GAF zero process (of the same intensity), the latter two are hard to set apart between themselves. It is therefore an interesting question to devise mathematical statistics that distinguish them.
From this point of view, the notion of rigidity becomes especially interesting once we note that the results outlined in Examples 2.17 and 2.18 are optimal, and thus serve as a means of distinguishing the point processes described therein:
  1. (1)
    For the Ginibre process, it was shown in [37] that the number of points is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
  2. (2)
    For the $\alpha$-GAF zeros, it was shown in [26, 37] that the $k^{\mathrm{th}}$ moments for $0\leq k\leq\lfloor 1/\alpha\rfloor$ is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
A formal version of these optimality results can be stated using the notion of tolerance; see, e.g., [36, Definitions 2 and 3].
Going back to our main topic, the number rigidity of $\Lambda^{\theta}$ was proved in a sequence of three papers. Namely, first for $\theta=(1,2,\infty)$ in [12], then for $\theta=(1,\beta,w)$ with any choice of $\beta>0$ and $w\in(-\infty,\infty]$ in [30], and finally for all $\theta\in\Theta$ in [29].
2.3.2. New Results
Theorem 1.2 refines current results on $\Lambda^{\theta}$’s rigidity in two directions: On the one hand, while the results of [29, 30] proved the existence of a single function $\mathcal{M}$ such that
(2.16) $\Lambda^{\theta}\cap B=\mathcal{M}(\Lambda^{\theta}\cap B^{c})\qquad\text{almost surely}$
for every bounded Borel set $B$, they did not provide an explicit construction of $\mathcal{M}$. In this view, Theorem 1.2 improves on the latter by identifying the exact mechanism that creates number rigidity: If we restrict the trace in (1.6) to the eigenvalues outside of some bounded set, then the fact that $\mathrm{e}^{-t\Lambda^{\theta}_{k}/2}=1+o(1)$ almost surely as $t\to 0$ for all fixed $k\geq 0$ implies the following:
Corollary 2.20.
For every bounded Borel set $B\subset\mathbb{R}$, it holds that
(2.17) $\sum_{k\geq 0\text{ s.t. }\Lambda^{\theta}_{k}\in B^{c}}\mathrm{e}^{-t\Lambda^{\theta}_{k}/2}=\sqrt{\frac{2}{\pi}}t^{-3/2}+\frac{1}{2}\left(r_{0}+\frac{1}{\beta}\right)-\frac{1}{4}-|\Lambda^{\theta}\cap B|+\mathfrak{o}_{t},\qquad t\in(0,1],$
where the random remainder terms $\mathfrak{o}_{t}$ satisfy (1.7) and (1.8). This provides a clear blueprint to construct the function $\mathcal{M}$ in (2.16) explicitly; see Remark 3.4 for details.
On the other hand, while the statement that $\Lambda^{\theta}$ is number rigid has intrinsic interest—as it says something nontrivial about correlations within SAO spectra—it nevertheless falls short of the ambition for the theory of rigidity laid out in Remark 2.19: Given that SAO eigenvalue point processes remain very similar as we vary the parameter $\theta$, we do not expect that different instances of $\Lambda^{\theta}$ satisfy different notions of rigidity, such as the rigidity/tolerance hierarchy observed for the Ginibre process and the $\alpha$-GAF zeroes in Examples 2.17 and 2.18.
Remark 2.21.
To the best of our knowledge, the optimality of number rigidity for $\Lambda^{\theta}$ (i.e., tolerance) has not been proved. That said, we expect that such a result could be proved using the technology developed in [26]. (The mutual absolute continuity of Palm measures of the Airy process—i.e., $\Lambda^{(1,2,\infty)}$—with the same number of fixed components proved in [13] serves as further evidence of this point.)
In this context, our main result can be viewed as a first step toward developing statistics that distinguish between different instances of SAO spectra, which goes beyond the rigidity/tolerance hierarchy:
Corollary 2.22.
If $\theta=(r,\beta,w)$ and $\tilde{\theta}=(\tilde{r},\tilde{\beta},\tilde{w})$ are such that $r_{0}+1/\beta\neq\tilde{r}_{0}+1/\tilde{\beta}$, then $\Lambda^{\theta}$ and $\Lambda^{\tilde{\theta}}$ are mutually singular.
2.3.3. Open Problems
Following-up on Corollary 2.22, it is natural to ask if there are other examples of parameter-dependent point processes that are mutually singular despite being known/expected to satisfy the same notions of rigidity. As a first step in this direction, we propose the following:
Open Problem 2.23.
[…]
Lastly, one can ask about the optimality of Corollary 2.22:
Open Problem 2.24.
Find a function $\Psi$ on the parameter space $\Theta$ such that $\Lambda^{\theta}$ and $\Lambda^{\tilde{\theta}}$ are mutually singular if and only if $\Psi(\theta)\neq\Psi(\tilde{\theta})$. Once $\Psi$ is found, characterize the possible relationships between the laws of $\Lambda^{\theta}$ and $\Lambda^{\tilde{\theta}}$ when $\Psi(\theta)=\Psi(\tilde{\theta})$ (e.g., equality, mutual absolute continuity, etc.).
2.4. Cancellation of Subcritical Coordinates
2.4.1. Background
In [8, Page 2734], Bloemendal and Virág made the curious observation that any given spiked Wishart or Gaussian invariant model converges to an infinite number of SAOs with different parameters: Looking back at Theorem 2.13, if we add a trivial component to the spike $\ell=(\ell_{r-1},\ldots,\ell)$, thus transforming it into ~$\tilde{\ell}=(0,\ell_{r-1},\ldots,\ell_{0})$, then the matrix model remains unchanged. However, the parameter $r$ in the limit (2.14) increases by one, and the added component has a Dirichlet boundary condition (since $\ell_{r}=0$ corresponds to $w_{r+1}=\infty$ in (2.13)). If we combine this observation with the fact that the SAOs are obviously invariant with respect to permutations of their components (as their noises are induced by GOE, GUE, and GSE matrices), then we obtain the following from Theorem 2.13:
Corollary 2.25 ([8]).
Let $\beta\in\{1,2,4\}$, and let $\theta=(r,\beta,w),\tilde{\theta}=(\tilde{r},\beta,\tilde{w})\in\Theta$ be such that
(2.18) $r_{0}=\tilde{r}_{0}$
(i.e., $\mathcal{H}_{\theta}$ and $\mathcal{H}_{\tilde{\theta}}$ have the same number of Robin boundary conditions) and
(2.19) $\{w_{1},\ldots,w_{r}\}\cap\mathbb{R}=\{\tilde{w}_{1},\ldots,\tilde{w}_{\tilde{r}}\}\cap\mathbb{R}$
(i.e., the Robin boundary conditions match, up to permutation). Then, $\Lambda^{\theta}\stackrel{{\scriptstyle\mathrm{distr.}}}{{=}}\Lambda^{\tilde{\theta}}$.
Given the somewhat indirect nature of the proof of this result, it is natural to ask:
Question 2.26.
Can one prove Corollary 2.25 using the intrinsic properties of the SAOs only (i.e., without reference to the matrix models that converge to the SAOs)?
2.4.2. New Result
While our results do not provide a complete answer to Question 2.26, they nevertheless shed new light on some of the delicate relationships that must exist between $\mathcal{H}_{\theta}$’s parameters in order for Corollary 2.25 to hold. To see this, in what follows we let $\theta=(r,\beta,w)$ with $r\geq 1$, $\beta\in\{1,2,4\}$, and $w\in(-\infty,\infty]^{r}$ be fixed, and we let $r_{0}$ be defined as in (1.4). Given $\kappa,\sigma,\upsilon>0$, define the parameter $\eta=(\kappa,\sigma,\upsilon)$, and define the Schrödinger operator
$\widehat{H}_{\theta,\eta}=-\tfrac{1}{2}\tfrac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+\kappa x+\xi,$
which acts on vector-valued $f\in L^{2}\big{(}[0,\infty),\mathbb{F}_{\beta}\big{)}^{r}$ with boundary conditions
$\begin{cases}f(i,0)=0&\text{if }w_{i}=\infty\text{ (Dirichlet)},\\ f^{\prime}(i,0)=w_{i}f(i,0)&\text{if }w_{i}\in\mathbb{R}\text{ (Robin)},\end{cases}\qquad 1\leq i\leq r,$
and where $\xi$ is a $r\times r$ matrix white noise whose diagonal components $\xi_{i,i}$ have variance $\sigma^{2}$, and whose off-diagonal components $\xi_{i,j}=\xi_{j,i}^{*}$ have variance $\upsilon^{2}$. The operators $\widehat{H}_{\theta,\eta}$ can be viewed as a generalization of the SAOs, since if we choose the parameter $\eta$ in such a way that $\kappa=\kappa(r)=\frac{r}{2}$, $\sigma^{2}=\frac{1}{\beta}$, and $\upsilon^{2}=\frac{1}{2}$, then $2\widehat{H}_{\theta,\eta}=\mathcal{H}_{\theta}$.
In this context, our new insight concerning Corollary 2.25 is that the trace asymptotic for $\mathcal{H}_{\theta}$ in (1.6) is a special case of the more general asymptotic
(2.20) $\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}=\tfrac{r}{\sqrt{2\pi}\kappa}t^{-3/2}+\tfrac{1}{4}\left(2r_{0}-r+\tfrac{r\sigma^{2}}{\kappa}+\tfrac{r(r-1)\upsilon^{2}}{\kappa}\right)+\mathfrak{o}_{t}\qquad\text{for all }t\in(0,1],$
where $\mathfrak{o}_{t}$ satisfy (1.7) and (1.8). With this in hand, we can obtain necessary conditions on the parameter $\eta$ in order for the cancellations in Corollary 2.25 to occur.
More specifically, let $\tilde{\theta}=(\tilde{r},\beta,\tilde{w})$ be related to $\theta$ in such a way that (2.18) and (2.19) hold. Then, consider $\eta=(\kappa,\sigma,\upsilon)$ such that, similarly to the SAOs,
(2.21) $\kappa=\kappa(r)\text{ possibly depends on }r,$
and conversely for the parameters that determine the noise’s variance,
(2.22) $\sigma\text{ and }\upsilon\text{ are independent of }r.$
In order for the operators $\widehat{H}_{\theta,\eta}$ and $\widehat{H}_{\tilde{\theta},\eta}$ to generate the same eigenvalue point process, it is necessary that every term in the asymptotic (2.20) be independent of $r$ (as this could be different from $\tilde{r}$), and instead only depend on the common parameter $r_{0}=\tilde{r}_{0}$. In order for the leading order term in (2.20) to be independent of $r$, it is necessary that $\kappa=\kappa(r)=\kappa_{0}r$ for some $\kappa_{0}>0$ independent of $r$. If such is the case, then (2.20) becomes
(2.23) $\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}=\tfrac{1}{\sqrt{2\pi}\kappa_{0}}t^{-3/2}+\tfrac{1}{4}\left(2r_{0}-r+\tfrac{\sigma^{2}}{\kappa_{0}}+\tfrac{(r-1)\upsilon^{2}}{\kappa_{0}}\right)+\mathfrak{o}_{t}.$
Once this is applied, the only $r$-dependence in (2.23) lies in the second order term, via the function $r\mapsto(\upsilon^{2}/\kappa_{0}-1)r$. Since $\upsilon$ is assumed independent of $r$, the only way for this to not depend on $r$ is to have $\upsilon^{2}=\kappa_{0}$, in which case (2.23) becomes
(2.24) $\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}=\tfrac{1}{\sqrt{2\pi}\kappa_{0}}t^{-3/2}+\tfrac{1}{4}\left(2r_{0}-1+\tfrac{\sigma^{2}}{\kappa_{0}}\right)+\mathfrak{o}_{t}.$
We thus obtain the following result:
Corollary 2.27.
Suppose that $\theta$ and $\tilde{\theta}$ satisfy (2.18) and (2.19), and that $\eta$ satisfies (2.21) and (2.22). If $\widehat{H}_{\theta,\eta}$ and $\widehat{H}_{\tilde{\theta},\eta}$ generate the same eigenvalue point process, then $\kappa(r)=r\upsilon^{2}$. In particular, if $\upsilon^{2}=\frac{1}{2}$ (as is the case for the SAOs), then the cancellation in Corollary 2.25 can only occur if $\kappa(r)=r/2$.
2.4.3. Open Problem
Based on this partial result, we expect that a more detailed study of the joint moments of $\mathrm{Tr}\big{[}\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big{]}$ (over multiple $t$’s) could be used to completely answer Question 2.26. As this lies outside the scope of the main applications of this paper, we leave this question open for future investigations.
3. Trace Asymptotics
We now embark on the proof of our trace asymptotic. The most general incarnation of the asymptotic is (2.20). For this purpose, we recall that, given $\theta=(r,\beta,w)\in\Theta$ and $\eta=(\kappa,\sigma,\upsilon)\in(0,\infty)^{3}$, we denote the operator
$\widehat{H}_{\theta,\eta}=-\tfrac{1}{2}\tfrac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+\kappa x+\xi,$
which acts on vector-valued $f\in L^{2}\big{(}[0,\infty),\mathbb{F}_{\beta}\big{)}^{r}$ with boundary conditions
$\begin{cases}f(i,0)=0&\text{if }w_{i}=\infty\text{ (Dirichlet)},\\ f^{\prime}(i,0)=w_{i}f(i,0)&\text{if }w_{i}\in\mathbb{R}\text{ (Robin)},\end{cases}\qquad 1\leq i\leq r,$
and where $\xi$ is a $r\times r$ matrix white noise whose diagonal components $\xi_{i,i}$ have variance $\sigma^{2}$, and whose off-diagonal components $\xi_{i,j}=\xi_{j,i}^{*}$ have variance $\upsilon^{2}$. The statement in (2.20) can be split into two claims:
Theorem 3.1.
For every $\theta\in\Theta$ and $\eta\in(0,\infty)^{3}$, one has
(3.1) $\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]=\tfrac{r}{\sqrt{2\pi}\kappa}t^{-3/2}+\tfrac{1}{4}\left(2r_{0}-r+\tfrac{r\sigma^{2}}{\kappa}+\tfrac{r(r-1)\upsilon^{2}}{\kappa}\right)+o(1)\qquad\text{as }t\to 0^{+},$
and there exists a constant $C>0$ such that
(3.2) $\mathbf{Cov}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-s\widehat{H}_{\theta,\eta}}\big{]},\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]\leq C\left(\frac{\min\{s,t\}}{\max\{s,t\}}\right)^{1/4}\qquad\text{ for every }s,t\in(0,1].$
The covariance bound in Theorem 3.1 was proved in [29, (8.4)]. Consequently, we only need to prove (3.1). The remainder of this section is organized as follows: Before we prove (3.1), in Section 3.1 we show how Theorem 3.1 is used to construct the function $\mathcal{T}$ in (1.5), and thus prove Theorem 1.2. Then, in Section 3.2, we state a Feynman-Kac formula in proved in [29] for the expected trace $\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]$. Finally, in Sections 3.3 and 3.4, we use the Feynman-Kac formula in question to prove the asymptotics in (3.1).
3.1. Proof of Theorem 1.2
We begin by defining $\mathcal{T}$:
Definition 3.2.
Let $c_{1},c_{2}>0$ be any fixed constants. Define the vanishing sequence of positive numbers $01\geq t_{1}>t_{2}>\cdots>$ 0 as
(3.3) $t_{n}=\mathrm{e}^{-c_{1}n},\qquad n\geq 1,$
and let $1\leq N_{1}<N_{2}<\cdots$ be the sequence of integers
(3.4) $N_{m}=\lceil m^{1+c_{2}}\rceil,\qquad m\geq 1.$
Given any point configuration $\mathcal{X}=\{x_{0},x_{1},x_{2},\ldots\}$ on $\mathbb{R}$, we define
(3.5) $\mathcal{T}(\mathcal{X})=\begin{cases}\displaystyle\frac{1}{2}+\lim_{m\to\infty}\frac{2}{N_{m}}\sum_{n=1}^{N_{m}}\left(\sum_{k=0}^{\infty}\mathrm{e}^{-t_{n}x_{k}/2}-\sqrt{\tfrac{2}{\pi}}t_{n}^{-3/2}\right)&\text{if the limit exists,}\\ 0&\text{otherwise.}\end{cases}$
Remark 3.3.
There is some degree of arbitrariness in how $t_{n}$ and $N_{m}$ were chosen in (3.3) and (3.4). As will be made clear by the proof in this section, any sequence of $t_{n}$’s and $N_{m}$’s that respectively vanish and diverge ”fast enough” suffice to obtain (1.5). Our specific choices in (3.3) and (3.4) are made for simplicity and to emphasize that it is possible to be completely explicit in our definition of the map $\mathcal{T}$.
Recall that if $\eta=(\frac{r}{2},\frac{1}{\beta},\frac{1}{2})$, then $2\widehat{H}_{\theta,\eta}=\mathcal{H}_{\theta}$. Thus, (3.1) implies that
$\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big{]}\right]=\sqrt{\tfrac{2}{\pi}}t^{-3/2}+\tfrac{r_{0}+1/\beta}{2}-\tfrac{1}{4}+o(1)\qquad\text{as }t\to 0^{+}.$
By a trivial rearrangement of the terms in (3.5), we therefore obtain that $\mathcal{T}(\Lambda^{\theta})=r_{0}+1/\beta$ almost surely if we can prove that
$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}-\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}\right]\right)=0\qquad\text{almost surely.}$.
Toward this end, by Markov’s inequality, we get that for every $m\in\mathbb{N}$ and $\varepsilon>0$,
$\mathbf{P}\left[\left|\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}-\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}\right]\right)\right|\geq m^{-\varepsilon/2}\right]$
$\leq$ $m^{\varepsilon}~{}\mathbf{E}\left[\frac{1}{N_{m}^{2}}\left(\sum_{n=1}^{N_{m}}\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}-\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}\right]\right)^{2}\right]$
(3.6) $=$ $\frac{m^{\epsilon}}{N_{m}^{2}}\sum_{n_{1},n_{2}=1}^{N_{m}}\mathbf{Cov}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n_{1}}\mathcal{H}_{\theta}/2}\big{]},\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n_{2}}\mathcal{H}_{\theta}/2}\big{]}\right].$
By the Borel-Cantelli lemma, it suffices to show that (3.6) is summable in $m$.
By (3.2), there is a constant $C>0$ independent of $N$ such that
$\frac{m^{\epsilon}}{N_{m}^{2}}\sum_{n_{1},n_{2}=1}^{N_{m}}\mathbf{Cov}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n_{1}}\mathcal{H}_{\theta}/2}\big{]},\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n_{2}}\mathcal{H}_{\theta}/2}\big{]}\right]\leq\frac{Cm^{\varepsilon}}{N_{m}^{2}}\sum_{n_{1},n_{2}=1}^{N_{m}}\left(\frac{\min\{t_{n_{1}},t_{n_{2}}\}}{\max\{t_{n_{1}},t_{n_{2}}\}}\right)^{1/4}.$
With $t_{n}$ defined as in (3.3), this simplifies to
$\frac{Cm^{\varepsilon}}{N_{m}^{2}}\sum_{n_{1},n_{2}=1}^{N_{m}}\left(\frac{\min\{t_{n_{1}},t_{n_{2}}\}}{\max\{t_{n_{1}},t_{n_{2}}\}}\right)^{1/4}=\frac{Cm^{\varepsilon}}{N_{m}^{2}}\sum_{n_{1},n_{2}=1}^{N_{m}}\mathrm{e}^{-c_{1}|n_{1}-n_{2}|/4}.$
We can split the above sum as follows:
$\frac{Cm^{\varepsilon}}{N_{m}^{2}}\sum_{\begin{subarray}{c}1\leq n_{1},n_{2}\leq N_{m}\\ |n_{1}-n_{2}|\leq(4/c_{1})\log N_{m}\end{subarray}}\mathrm{e}^{-c_{1}|n_{1}-n_{2}|/4}+\frac{Cm^{\varepsilon}}{N_{m}^{2}}\sum_{\begin{subarray}{c}1\leq n_{1},n_{2}\leq N_{m}\\ |n_{1}-n_{2}|>(4/c_{1})\log N_{m}\end{subarray}}\mathrm{e}^{-c_{1}|n_{1}-n_{2}|/4}.$
Given that $m^{1+c_{2}}\leq N_{m}\leq m^{1+c_{2}}+1$ (as per (3.4)) and $\mathrm{e}^{-c_{1}|n_{1}-n_{2}|/4}\leq 1$, we get
$\frac{Cm^{\varepsilon}}{N_{m}^{2}}\sum_{\begin{subarray}{c}1\leq n_{1},n_{2}\leq N_{m}\\ |n_{1}-n_{2}|\leq(4/c_{1})\log N_{m}\end{subarray}}\mathrm{e}^{-c_{1}|n_{1}-n_{2}|/4}=O\left(\frac{m^{\varepsilon}}{N_{m}^{2}}\cdot N_{m}\log N_{m}\right)=O(m^{-1-(c_{2}-\varepsilon)}\log m)$
as $m\to\infty$. Moreover,
$\frac{Cm^{\varepsilon}}{N_{m}^{2}}\sum_{\begin{subarray}{c}1\leq n_{1},n_{2}\leq N_{m}\\ |n_{1}-n_{2}|>(4/c_{1})\log N_{m}\end{subarray}}\mathrm{e}^{-c_{1}|n_{1}-n_{2}|/4}\\ \leq\frac{Cm^{\varepsilon}\mathrm{e}^{-\log N_{m}}}{N_{m}^{2}}\sum_{n_{1},n_{2}=1}^{N_{m}}1=O(m^{\varepsilon}N_{m}^{-1})=O(m^{-1-(c_{2}-\varepsilon)})$
as $m\to\infty$. If we choose $\varepsilon<c_{2}$, then this implies that (3.6) is summable, which concludes the proof that $\mathcal{T}(\Lambda^{\theta})=r_{0}+1/\beta$ almost surely.
Remark 3.4.
Following-up on the calculations performed in this section, we see that
$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\sqrt{\tfrac{2}{\pi}}t_{n}^{-3/2}+\tfrac{1}{2}\left(r_{0}+\tfrac{1}{\beta}\right)-\tfrac{1}{4}-\sum_{k=0}^{\infty}\mathrm{e}^{-t_{n}\Lambda^{\theta}_{k}/2}\right)=0\qquad\text{almost surely},$
where $t_{n}$ and $N_{m}$ are defined as in (3.3) and (3.4) respectively. Moreover, for any $k\geq 0$, one has
$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\mathrm{e}^{-t_{n}\Lambda^{\theta}_{k}/2}=1\qquad\text{almost surely.}$.
Consequently, if we define the map
$\mathcal{M}(\mathcal{X})=\begin{cases}\displaystyle\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\sqrt{\tfrac{2}{\pi}}t_{n}^{-3/2}+\tfrac{1}{2}\left(r_{0}+\tfrac{1}{\beta}\right)-\tfrac{1}{4}-\sum_{k=0}^{\infty}\mathrm{e}^{-t_{n}x_{k}/2}\right)&\text{if the limit exists,}\\ 0&\text{otherwise,}\end{cases}$
for any point configuration $\mathcal{X}=\{x_{0},x_{1},x_{2},\ldots\}\subset\mathbb{R}$, then (2.16) holds. Indeed, for any bounded set $B\subset\mathbb{R}$, we can write
(3.7) $\mathcal{M}(\Lambda^{\theta}\cap B^{c})=\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\sqrt{\tfrac{2}{\pi}}t_{n}^{-3/2}+\tfrac{1}{2}\left(r_{0}+\tfrac{1}{\beta}\right)-\tfrac{1}{4}-\sum_{k:\Lambda^{\theta}_{k}\in B^{c}}^{\infty}\mathrm{e}^{-t_{n}\Lambda^{\theta}_{k}/2}\right)\\ =\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\sqrt{\tfrac{2}{\pi}}t_{n}^{-3/2}+\tfrac{1}{2}\left(r_{0}+\tfrac{1}{\beta}\right)-\tfrac{1}{4}-\sum_{k=0}^{\infty}\mathrm{e}^{-t_{n}\Lambda^{\theta}_{k}/2}\right)+\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\sum_{k:\Lambda^{\theta}_{k}\in B}\mathrm{e}^{-t_{n}\Lambda^{\theta}_{k}/2};$
the first limit on the second line of (3.7) is almost-surely zero, and the second is almost-surely $|\Lambda^{\theta}\cap B|$.
3.2. Proof of (3.1) Step 1. Expected Trace Formula
The first step in our proof of (3.1) is to state the Feynman-Kac formula for $\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]$ obtained in [29, Theorem 4.9], which is what we will use to calculate the leading orders of the expected trace. This is stated as Lemma 3.13 below. Due to the complexity of the formula in question, we must set up a sizeable amount of notation to state it. First, we introduce the main stochastic process that drives the Feynman-Kac formula, which is the reflected Brownian bridge and its local times:
Definition 3.5.
Let $B$ be a standard Brownian motion on $\mathbb{R}$, and denote $B$’s transition kernel by
$\Pi_{B}(t;x,y)=\frac{\mathrm{e}^{-(x-y)^{2}/2t}}{\sqrt{2\pi t}},\qquad t>0,~{}x,y\in\mathbb{R}.$
We denote $B^{x}=(B|B(0)=x)$ and $B^{x,y}_{t}=(B|B(0)=x\text{ and }B(t)=y)$. Then, let $X=|B|$ be a reflected standard Brownian motion on $[0,\infty)$, and denote $X$’s transition kernel by
$\Pi_{X}(t;x,y)=\frac{\mathrm{e}^{-(x-y)^{2}/2t}}{\sqrt{2\pi t}}+\frac{\mathrm{e}^{-(x+y)^{2}/2t}}{\sqrt{2\pi t}},\qquad t>0,~{}x,y\geq 0.$
We similarly denote $X^{x}=(X|X(0)=x)$ and $X^{x,y}_{t}=(X|X(0)=x\text{ and }X(t)=y)$.
Definition 3.6.
Fix $t>0$, and let $Z$ be either $B$, $X$, or one of its conditioned version (i.e., $B^{x}$, $X^{x}$, $B^{x,y}_{t}$, or $X^{x,y}_{t}$). Given $0\leq u<v\leq t$, we let $\mathfrak{L}_{[u,v)}^{0}(Z)$ denote $Z$’s boundary local time collected on the time interval $[u,v)$, that is,
$\mathfrak{L}^{0}_{[u,v)}(Z)=\lim_{\varepsilon\to 0}\frac{1}{2\varepsilon}\int_{u}^{v}\mathbf{1}_{\{-\varepsilon<Z(s)<\varepsilon\}}~{}\mathrm{d}s\qquad\text{almost surely}.$
Moreover, we let $\big{(}L^{y}_{[u,v)}(Z):y\in[0,\infty)\big{)}$ denote the continuous version of $Z$’s local time process on $[u,v)$, that is the process such that
$\int_{u}^{v}f\big{(}Z(s)\big{)}~{}\mathrm{d}s=\int_{0}^{\infty}L^{y}_{[u,v)}(Z)f(y)~{}\mathrm{d}y\qquad\text{for every measurable }f:\mathbb{R}\to\mathbb{R}.$
Finally, we use the shorthands $\mathfrak{L}_{t}^{0}(Z)=\mathfrak{L}_{[0,t)}^{0}(Z)$ and $L_{t}^{y}(Z)=L_{[0,t)}^{y}(Z)$.
If we were only dealing with scalar-valued Schrödinger operators on $[0,\infty)$, then these would be the only processes that are needed to state the Feynman-Kac formula. However, since we are dealing with the vector-valued setting, we need more tools. First, we set up some notation to describe the function spaces related to $\widehat{H}_{\theta,\eta}$:
Definition 3.7.
Given $f,g:[0,\infty)\to\mathbb{F}_{\beta}$, we denote
$\langle f,g\rangle=\int_{0}^{\infty}f(x)g(x)~{}\mathrm{d}x\qquad\text{and}\qquad\|f\|_{2}=\sqrt{\langle f,f\rangle}.$
We denote $\mathcal{A}=\{1,\ldots,r\}\times[0,\infty)$, and we let $\mu$ denote the measure on $\mathcal{A}$ obtained by taking the product of the counting measure on $\{1,\ldots,r\}$ and the Lebesgue measure on $[0,\infty)$. Given $f:\mathcal{A}\to\mathbb{F}_{\beta}$, we denote
$\|f\|_{\mu}=\sqrt{\int_{\mathcal{A}}|f(a)|^{2}~{}\mathrm{d}\mu(a)}=\sqrt{\sum_{i=1}^{r}\|f(i,\cdot)\|_{2}^{2}}.$
Next, the vector-valued setting requires the introduction of a jump process on $\{1,\ldots,r\}$ which depends on $X$’s self-intersections as follows:
Definition 3.8.
Let $\mathcal{P}_{0}=\varnothing$ be the empty set, and given an even integer $n\geq 2$, let $\mathcal{P}_{n}$ denote the set of perfect pair matchings of $\{1,\ldots,n\}$. That is, $\mathcal{P}_{n}$ is the set
$\Big{\{}\big{\{}\{\ell_{1},\ell_{2}\},\{\ell_{3},\ell_{4}\},\ldots,\{\ell_{n-1},\ell_{n}\}\big{\}}:\ell_{i}\in\{1,\ldots,n\}\text{ and }\{\ell_{i},\ell_{i+1}\}\cap\{\ell_{j},\ell_{j+1}\}=\varnothing\text{ if }j\neq i\Big{\}}.$
Definition 3.9.
Let $t>0$ and $x\geq 0$, let $n\geq 2$ be an even integer, and let $q\in\mathcal{P}_{n}$. Given a realization of $X^{x,x}_{t}$, we let $\mathrm{si}_{t,n,q,X^{x,x}_{t}}$ denote the unique random Borel probability measure on $[0,t)^{n}$ such that for every $s_{1},t_{1},\ldots,s_{n},t_{n}\in[0,t)$ with $s_{k}<t_{k}$ for $k=1,\ldots,n$, one has
(3.8) $\mathrm{si}_{t,n,q,X^{x,x}_{t}}\big{(}[s_{1},t_{1})\times[s_{2},t_{2})\times\cdots\times[s_{n},t_{n})\big{)}\\ =\frac{1}{\|L_{t}(X^{x,x}_{t})\|_{2}^{2n}}\prod_{\{\ell_{1},\ell_{2}\}\in q}\left\langle L_{[s_{\ell_{1}},t_{\ell_{1}})}(X^{x,x}_{t}),L_{[s_{\ell_{2}},t_{\ell_{2}})}(X^{x,x}_{t})\right\rangle$
(this uniquely determines the measure since product rectangles are a determining class for the Borel $\sigma$-algebra).
Definition 3.10.
Let $t>0$ and $x\geq 0$ be fixed, and suppose that we are given a realization of $X^{x,x}_{t}$. For any $1\leq i\leq r$, we generate the jump process $\widehat{U}^{i}_{t}:[0,t]\to\{1,\ldots,r\}$ using the following step-by-step procedure:
  1. $\bullet$
    Given $X^{x,x}_{t}$, sample $\widehat{N}(t)$ according to the distribution
    (3.9) $\mathbf{P}\big{[}\widehat{N}(t)=2m\big{]}=\left(\frac{(r-1)^{2}\|L_{t}(X^{x,x}_{t})\|^{2}_{2}}{2}\right)^{m}\frac{1}{m!}\mathrm{e}^{-\frac{(r-1)^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}},\qquad m=0,1,2,\ldots,$
    i.e., $\widehat{N}(t)/2$ is Poisson with parameter $\frac{(r-1)^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}$.
  2. $\bullet$
    Given $\widehat{N}(t)$, sample $\widehat{q}$ uniformly at random in $\mathcal{P}_{\widehat{N}(t)}$.
  3. $\bullet$
    Given $X^{x,x}_{t}$, $\widehat{N}(t)$, and $\widehat{q}$, draw $\boldsymbol{T}=(T_{1},\ldots,T_{\widehat{N}(t)})\in[0,t]^{\widehat{N}(t)}$ according to the self-intersection measure $\mathrm{si}_{t,\widehat{N}(t),\widehat{q},X^{x,x}_{t}}$. Then, we let $\widehat{\boldsymbol{\tau}}=(\widehat{\tau}_{1},\ldots,\widehat{\tau}_{\widehat{N}(t)})$ be the tuple $\boldsymbol{T}$ rearranged in increasing order.
  4. $\bullet$
    Given $\widehat{N}(t)$, $\boldsymbol{T}$, and $\widehat{\boldsymbol{\tau}}$, we let $\widehat{\pi}$ be the random permutation of $\{1,\ldots,\widehat{N}(t)\}$ such that $\widehat{\tau}_{\widehat{\pi}(\ell)}=T_{\ell}$ for all $\ell\leq\widehat{N}(t)$, and we let $\widehat{p}\in\mathcal{P}_{\widehat{N}(t)}$ be the random perfect matching such that $\{\ell_{1},\ell_{2}\}\in\widehat{q}$ if and only if $\{\widehat{\pi}(\ell_{1}),\widehat{\pi}(\ell_{2})\}\in\widehat{p}$. In words, $\widehat{p}$ ensures that each pair of times $T_{k}$ that were matched by $\widehat{q}$ are still matched once arranged in nondecreasing order $\widehat{\tau}_{k}$.
  5. $\bullet$
    Let $M^{i}=(M^{i}_{0},M^{i}_{1},M^{i}_{2},\ldots)$ be a uniform random walk on the complete graph (without self-edges) on $\{1,\ldots,r\}$ with starting point $M^{i}_{0}=i$, i.e.,
    (3.10) $\mathbf{P}\big{[}M^{i}_{k}=i_{2}~{}\big{|}~{}M^{i}_{k-1}=i_{1}\big{]}=\frac{\mathbf{1}_{\{i_{2}\neq i_{1}\}}}{r-1},\qquad k\geq 1,~{}i_{1},i_{2}\in\{1,\ldots,r\}.$
    We assume $M^{i}$ is independent of $X^{x,x}_{t}$, $\widehat{N}(t)$, and $\widehat{\boldsymbol{\tau}}$.
  6. $\bullet$
    Combining all of the above, we then let $\big{(}\widehat{U}^{i}_{t}(s):1\leq s\leq t\big{)}$ be the càdlàg path with jump times $\widehat{\boldsymbol{\tau}}$ and jumps $M^{i}$:
    (3.11) $\widehat{U}^{i}_{t}(s)=i\mathbf{1}_{[0,\widehat{\tau}_{1})}(s)+\sum_{k=1}^{\widehat{N}(t)-1}M^{i}_{k}\mathbf{1}_{[\widehat{\tau}_{k},\widehat{\tau}_{k+1})}(s)+M_{\widehat{N}(t)}\mathbf{1}_{[\widehat{\tau}_{\widehat{N}(t)},t]}(s),\qquad 0\leq s\leq t.$
    In particular, if $\widehat{N}(t)=0$, then $\widehat{U}^{i}_{t}(t)=i$ is constant.
Then, we define local times of the combined processes $(\widehat{U}^{i}_{t},X^{x,x}_{t})$ as follows: For any $t>0$ and $(j,y)\in\mathcal{A}=\{1,\ldots,r\}\times[0,\infty)$, we let
(3.12) $\mathfrak{L}^{(j,0)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=\sum_{k\geq 1:~{}U^{i}_{t}(\widehat{\tau}_{k-1})=j}\mathfrak{L}_{[\widehat{\tau}_{k-1},\widehat{\tau}_{k})\cap[0,t]}^{0}(X^{x,x}_{t})$
and
(3.13) $L^{(j,y)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=\sum_{k\geq 1:~{}U^{i}_{t}(\widehat{\tau}_{k-1})=j}L_{[\widehat{\tau}_{k-1},\widehat{\tau}_{k})\cap[0,t]}^{y}(X^{x,x}_{t})$
respectively denote $X^{x,x}_{t}$’s boundary local time at zero and local time collected on the intervals of $[0,t]$ where the jump process $\widehat{U}^{i}_{t}$ is equal to $j$. We note that since $[u,v)\mapsto\mathfrak{L}^{0}_{[u,v)}(X^{x,x}_{t})$ and $[u,v)\mapsto L^{y}_{[u,v)}(X^{x,x}_{t})$ induce measures on $[0,\infty)$, it is clear that
(3.14) $\sum_{j=1}^{r}\mathfrak{L}^{(j,0)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=\mathfrak{L}_{t}(X^{x,x}_{t})\quad\text{and}\quad\sum_{j=1}^{r}L^{(j,y)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=L^{y}_{t}(X^{x,x}_{t}).$
Finally, the computation of the moment $\mathbf{E}\Big{[}\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\Big{]}$ requires introducing the following combinatorial constant:
Definition 3.11.
Given $n\in\mathbb{N}$, let $\mathcal{B}_{n}$ denote the set of binary sequences with $n$ steps, namely, sequences of the form $m=(m_{0},m_{1},\ldots,m_{n})\in\{0,1\}^{n+1}$. Given $h,l\in\{0,1\}$, we let $\mathcal{B}^{h,l}_{n}$ denote the set of sequences $m\in\mathcal{B}_{n}$ such that $m_{0}=h$ and $m_{n}=l$.
Definition 3.12.
Let $t>0$ and $1\leq i\leq r$ be fixed. Suppose we are given a realization of $\widehat{N}(t)\geq 2$, the matching $\widehat{p}\in\mathcal{P}_{\widehat{N}(t)}$, and the process $\widehat{U}^{i}_{t}$ in Definition 3.10. Letting $M^{i}=(M^{i}_{0},M^{i}_{1},\ldots)$ be as in Definition 3.10 (i.e., the process that determines the steps taken by $\widehat{U}^{i}_{t}$ in $\{1,\ldots,r\}$), we denote the jumps
$J_{k}=\big{(}M^{i}_{k-1},M^{i}_{k}\big{)},\qquad k\geq 1.$
Then, we let $J_{k}^{*}=\big{(}M^{i}_{k},M^{i}_{k-1}\big{)}$ denote these same jumps in reverse order. We define the combinatorial constant $\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})$ as follows:
  1.   (a)
    If $\mathbb{F}_{\beta}=\mathbb{R}$ (i.e., $\beta=1$), then
    $\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})=\begin{cases}1&\text{if either $J_{\ell_{1}}=J_{\ell_{2}}$ or $J_{\ell_{1}}=J_{\ell_{2}}^{*}$ for all $\{\ell_{1},\ell_{2}\}\in\widehat{p}$},\\ 0&\text{otherwise}.\end{cases}$
  2.   (b)
    If $\mathbb{F}_{\beta}=\mathbb{C}$ (i.e., $\beta=2$), then
    $\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})=\begin{cases}1&\text{if $J_{\ell_{1}}=J_{\ell_{2}}^{*}$ for all $\{\ell_{1},\ell_{2}\}\in\widehat{p}$},\\ 0&\text{otherwise}.\end{cases}$
  3.   (c)
    If $\mathbb{F}_{\beta}=\mathbb{H}$ (i.e., $\beta=4$), then
    $\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})=\begin{cases}\mathfrak{D}_{t}(\widehat{p},\widehat{U}^{i}_{t})&\text{if either $J_{\ell_{1}}=J_{\ell_{2}}$or $J_{\ell_{1}}=J_{\ell_{2}}^{*}$ for all $\{\ell_{1},\ell_{2}\}\in\widehat{p}$},\\ 0&\text{otherwise}.\end{cases}$
    Here, we define $\mathfrak{D}_{t}(\widehat{p},\widehat{U}^{i}_{t})$ as follows: A binary sequence $m\in\mathcal{B}^{0,0}_{N(t)}$ is said to respect $(\widehat{p},J)$ if for every pair $\{\ell_{1},\ell_{2}\}\in\widehat{p}$, the following holds:
    1. (c.1)
      if $J_{\ell_{1}}=J_{\ell_{2}}$, then $\big{(}(m_{\ell_{1}-1},m_{\ell_{1}}),(m_{\ell_{2}-1},m_{\ell_{2}})\big{)}$ are equal to one of the following pairs: $\big{(}(0,0),(1,1)\big{)}$, $\big{(}(1,1),(0,0)\big{)}$, $\big{(}(0,1),(1,0)\big{)}$, or $\big{(}(1,0),(0,1)\big{)}$.
    2. (c.2)
      if $J_{\ell_{1}}=J_{\ell_{2}}^{*}$, then $\big{(}(m_{\ell_{1}-1},m_{\ell_{1}}),(m_{\ell_{2}-1},m_{\ell_{2}})\big{)}$ are equal to one of the following pairs: $\big{(}(0,0),(0,0)\big{)}$, $\big{(}(1,1),(1,1)\big{)}$, $\big{(}(0,1),(1,0)\big{)}$, or $\big{(}(1,0),(0,1)\big{)}$.
    Lastly, we call any pair $\{\ell_{1},\ell_{2}\}\in p$ such that $J_{\ell_{1}}=J_{\ell_{2}}$ and
    $\big{(}(m_{\ell_{1}-1},m_{\ell_{1}}),(m_{\ell_{2}-1},m_{\ell_{2}})\big{)}=\big{(}(0,1),(1,0)\big{)}\text{ or }\big{(}(1,0),(0,1)\big{)}$
    a flip, and we let $f(m,\widehat{p},J)$ denote the total number of flips that occur in the combination of $m,\widehat{p},$ and $J$. With all this in hand, we finally define
    (3.15) $\mathfrak{D}_{t}(\widehat{p},\widehat{U}^{i}_{t})=2^{-\widehat{N}(t)/2}\sum_{m\in\mathcal{B}^{0,0}_{N(t)},~{}m\text{ respects }(\widehat{p},J)}(-1)^{f(m,\widehat{p},J)}.$
We are now finally in a position to state our Feynman-Kac formula for the trace expectation, which is a special case of [29, Theorem 4.9]:
Lemma 3.13.
For every $t>0$, $x\geq 0$, and $1\leq i\leq r$, define
(3.16) $\mathfrak{M}_{t}(\widehat{p},\widehat{U}^{i}_{t})=\begin{cases}1&\text{if }\widehat{N}(t)=0,\\ \upsilon^{\widehat{N}(t)}\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})&\text{otherwise},\end{cases}$
and
(3.17) $\mathfrak{H}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=-\kappa\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s\\ +\frac{(r-1)^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}+\frac{\sigma^{2}}{2}\|L_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})\|_{\mu}^{2}-\sum_{j=1}^{r}w_{j}\mathfrak{L}^{(j,0)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t}).$
For every $t>0$, one has
(3.18) $\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]=\int_{\mathcal{A}}\Pi_{X}(t;x,x)\mathbf{E}\left[\mathfrak{M}_{t}(\widehat{p},\widehat{U}^{i}_{t})\mathrm{e}^{\mathfrak{H}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})}\mathbf{1}_{\{\widehat{U}^{i}_{t}(t)=i\}}\right]~{}\mathrm{d}\mu(i,x).$
Remark 3.14.
Theorem 4.9 in [29] provides a formula for any joint moment of the form $\mathbf{E}\big{[}\mathrm{Tr}\big{[}\mathrm{e}^{-t_{1}\widehat{H}}\big{]}\cdots\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\widehat{H}}\big{]}\big{]}$ for general vector-valued Schrödinger operators $\widehat{H}=-\frac{1}{2}\Delta+V+\xi$ on any one-dimensional domain $I\subset\mathbb{R}$. The statement in Lemma 3.13 can be recovered from the latter by taking $n=1$ and $t=t_{1}$, then considering what is called ”Case 2” in [29] (i.e., an operator on the half-line), and then taking the potential function $V(i,x)=\kappa x$.
Remark 3.15.
The case where $w_{j}=\infty$ leads to some potential ambiguity in the definition of $\mathfrak{H}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})$ in (3.17). Using the convention $\mathrm{e}^{-\infty\cdot c}=\mathbf{1}_{\{c=0\}}$ for $c\geq 0$, we interpret the contribution of $\mathrm{e}^{-\infty\cdot\mathfrak{L}^{(j,0)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})}$ in (3.18) as follows:
$\mathrm{e}^{-\infty\cdot\mathfrak{L}^{(j,0)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})}=\mathbf{1}_{\{\mathfrak{L}^{(j,0)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=0\}}=\mathbf{1}_{\{X^{x,x}_{t}(s)\neq 0\text{ whenever }\widehat{U}^{i}_{t}(s)=j\}}.$
3.3. Proof of (3.1) Step 2. Simplifications
While (3.18) provides an exact formula for the expected trace, it is quite difficult to parse due to its complexity. Thus, our goal in this section is to reformulate (3.18) in a way that is amenable to computation, at least for the purposes of extracting the first- and second-order asymptotics.
For this purpose, we observe that most of the complexity in (3.18) comes from the contribution of the path $\widehat{U}^{i}_{t}$ to the combined local times (3.12) and (3.13) and the combinatorics of $\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})$. However, these terms are only difficult to understand when $\widehat{U}^{i}_{t}$’s path has numerous jumps. In this context, our ability to calculate the asymptotic (3.1) despite the complexity of (3.18) comes from the fact that the leading order terms in the latter only involve realizations of $\widehat{U}^{i}_{t}$ that have no jump or two jumps, thus simplifying the analysis.
In fact, the unifying theme of the simplifications performed in this section (as evidenced by Proposition 3.16 below) is that the exact calculation of leading terms and the estimate of remainder terms in (3.18) can both be reduced to the small-time asymptotics of three well-understood functionals of $X^{x,x}_{t}$ (neither or which involve $\widehat{U}^{i}_{t}$), namely, $\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s$, $\|L_{t}(X^{x,x}_{t})\|_{2}$, and $\mathfrak{L}_{t}^{0}(X^{x,x}_{t})$. We postpone the proofs of these asymptotics to Section 3.4.
To begin making this precise, we note that we can use (3.18) to split
(3.19) $\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]=T_{0}(t)+T_{2}(t)+T_{\geq 4}(t)$
into a sum of expectations with an indicator restricting the value of $\widehat{N}(t)$ as follows:
(3.20) $T_{0}(t)$ $=\int_{\mathcal{A}}\Pi_{X}(t;x,x)\mathbf{E}\left[\mathfrak{M}_{t}(\widehat{p},\widehat{U}^{i}_{t})\mathrm{e}^{\mathfrak{H}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})}\mathbf{1}_{\{\widehat{U}^{i}_{t}(t)=i\}\cap\{\widehat{N}(t)=0\}}\right]~{}\mathrm{d}\mu(i,x),$
(3.21) $T_{2}(t)$ $=\int_{\mathcal{A}}\Pi_{X}(t;x,x)\mathbf{E}\left[\mathfrak{M}_{t}(\widehat{p},\widehat{U}^{i}_{t})\mathrm{e}^{\mathfrak{H}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})}\mathbf{1}_{\{\widehat{U}^{i}_{t}(t)=i\}\cap\{\widehat{N}(t)=2\}}\right]~{}\mathrm{d}\mu(i,x),$
(3.22) $T_{\geq 4}(t)$ $=\int_{\mathcal{A}}\Pi_{X}(t;x,x)\mathbf{E}\left[\mathfrak{M}_{t}(\widehat{p},\widehat{U}^{i}_{t})\mathrm{e}^{\mathfrak{H}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})}\mathbf{1}_{\{\widehat{U}^{i}_{t}(t)=i\}\cap\{\widehat{N}(t)\geq 4\}}\right]~{}\mathrm{d}\mu(i,x).$
We discuss the contribution of each of these three terms to the $t\to 0^{+}$ asymptotic in Sections 3.3.13.3.3, and then summarize our findings in Section 3.3.4.
3.3.1. The Constant Path
We begin by simplifying $T_{0}(t)$ in (3.20). This term is the easiest to deal with, since if $\widehat{N}(t)=0$, then $\widehat{U}^{i}_{t}=i$ is constant on $[0,t]$. Therefore, we get the following three immediate simplifications when $\widehat{N}(t)=0$:
  1. (1)
    $\mathfrak{M}_{t}(\widehat{p},\widehat{U}^{i}_{t})=1$ by definition in (3.16).
  2. (2)
    Recalling (3.12) and (3.13), we see that
    $\mathfrak{L}^{(j,0)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=\begin{cases}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})&\text{if }j=i,\\ 0&\text{if }j\neq i,\end{cases}\quad and\quad L^{(j,y)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=\begin{cases}L^{y}_{t}(X^{x,x}_{t})&\text{if }j=i,\\ 0&\text{if }j\neq i;\end{cases}$
    hence (3.17) simplifies to
    $\mathfrak{H}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=-\kappa\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s+\frac{(r-1)^{2}+\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}-w_{i}\mathfrak{L}^{0}_{t}(X^{x,x}_{t}).$
  3. (3)
    $\mathbf{1}_{\{\widehat{U}^{i}_{t}(t)=i\}}=1$.
If we combine all of these simplifications into (3.20), then we get
(3.23) $T_{0}(t)=\sum_{i=1}^{r}\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathbf{E}\left[\mathrm{e}^{-\kappa\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s+\frac{(r-1)^{2}+\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}-w_{i}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})}\mathbf{1}_{\{\widehat{N}(t)=0\}}\right]~{}\mathrm{d}x.$
By Definition 3.10 (more specifically, the definition of $\widehat{N}(t)$ in (3.9) therein),
$\mathbf{E}\big{[}\mathbf{1}_{\{\widehat{N}(t)=0\}}\big{|}X^{x,x}_{t}\big{]}=\mathbf{P}\big{[}\widehat{N}(t)=0\big{|}X^{x,x}_{t}\big{]}=\mathrm{e}^{-\frac{(r-1)^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}}.$
Thus, by the tower property (conditioning on $X^{x,x}_{t}$), (3.23) simplifies to
(3.24) $T_{0}(t)=\sum_{i=1}^{r}\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathbf{E}\left[\mathrm{e}^{-\kappa\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s+\frac{\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}-w_{i}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})}\right]~{}\mathrm{d}x.$
In other words, if for $\mathsf{w}\in(-\infty,\infty]$ we denote
(3.25) $T_{0}^{(\mathsf{w})}(t)=\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathbf{E}\left[\mathrm{e}^{-\kappa\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s+\frac{\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}-\mathsf{w}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})}\right]~{}\mathrm{d}x,$
then we have that
(3.26) $T_{0}(t)=T_{0}^{(w_{1})}(t)+\cdots+T_{0}^{(w_{r})}(t).$
We now focus on understanding the asymptotics of each $T_{0}^{(w_{i})}(t)$ individually.
Toward this end, we first note that we expect
(3.27) $-\kappa\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s\approx-\kappa tx\qquad\text{as $t\to 0^{+}$ for fixed }x>0.$
Using this as a guide, we write
$T_{0}^{(\mathsf{w})}(t)=\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\left[\mathrm{e}^{-\kappa\int_{0}^{t}(X^{x,x}_{t}(s)-x)~{}\mathrm{d}s+\frac{\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}-\mathsf{w}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})}\right]~{}\mathrm{d}x.$
Next, by a Taylor expansion, we can write
(3.28) $\mathrm{e}^{z}=1+z+\mathsf{R}_{2}(z),\qquad\text{where }|\mathsf{R}_{2}(z)|\leq z^{2}\mathrm{e}^{|z|}.$
Thus, we obtain
(3.29) $T^{(\mathsf{w})}_{0}(t)=\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\Bigg{[}1+\tfrac{\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}\\ -\kappa{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s-\mathsf{w}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\\ +\mathsf{R}_{2}\Big{(}-\kappa{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s+\tfrac{\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}-\mathsf{w}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\Big{)}\Bigg{]}~{}\mathrm{d}x$
when $\mathsf{w}\in\mathbb{R}$, and, using Remark 3.15 to interpret $\mathrm{e}^{-\infty\cdot\mathfrak{L}^{0}_{t}(X^{x,x}_{t})}=\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}$, we get
(3.30) $T^{(\infty)}_{0}(t)=\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\Bigg{[}\left(1+\tfrac{\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}\right)\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\\ -\kappa\left({\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\right)\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\\ +\mathsf{R}_{2}\Big{(}-\kappa{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s+\tfrac{\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}\Big{)}\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\Bigg{]}~{}\mathrm{d}x.$
With this in hand, we now state a number of asymptotics of the functionals of $X^{x,x}_{t}$ appearing in (3.29) and (3.30), whose proofs we postpone to Section 3.4 below:
Proposition 3.16.
As $t\to 0^{+}$, we have the leading order contributions:
(3.31) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}~{}\mathrm{d}x=\tfrac{1}{2\pi\kappa}t^{-3/2}+\tfrac{1}{4}+o(1),$
(3.32) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}\big{]}~{}\mathrm{d}x=\tfrac{1}{2\kappa}+o(1)$
for the case $\mathsf{w}\in\mathbb{R}$, and
(3.33) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\big{]}~{}\mathrm{d}x=\tfrac{1}{2\pi\kappa}t^{-3/2}-\tfrac{1}{4}+o(1),$
(3.34) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\big{]}~{}\mathrm{d}x=\tfrac{1}{2\kappa}+o(1)$
for $\mathsf{w}=\infty$. Moreover, as $t\to 0^{+}$, we have the remainder terms
(3.35) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\Big{[}{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\Big{]}~{}\mathrm{d}x=o(1),$
(3.36) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\big{]}~{}\mathrm{d}x=o(1),$
and for every $\mathfrak{c}>0$,
(3.37) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\Bigg{[}\Big{(}\big{|}{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(X^{x,x}_{t})\|_{2}+\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\Big{)}^{2}\\ \cdot\exp\Bigg{\{}\mathfrak{c}\Big{(}\big{|}{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(X^{x,x}_{t})\|_{2}^{2}+\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\Big{)}\Bigg{\}}\Bigg{]}~{}\mathrm{d}x=o(1).$
On the one hand, if we apply (3.31) and (3.32) to the first line of (3.29), then apply (3.35) and (3.36) to the second line of (3.29), and finally apply (3.37) to the third line of (3.29) (as (3.37) matches the upper bound on $\mathsf{R}_{2}$ in (3.28)), then we get
$T_{0}^{(\mathsf{w})}(t)=\tfrac{1}{2\pi\kappa}t^{-3/2}+\tfrac{1}{4}\left(\tfrac{\sigma^{2}}{\kappa}+1\right)+o(1)\qquad\text{as }t\to 0^{+}$
when $\mathsf{w}\in\mathbb{R}$. On the other hand, if we apply (3.33) and (3.34) to the first line of (3.30), and then apply (3.35) and (3.37) to the second and third lines of (3.30) respectively (using the trivial bound $\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\leq 1$ in both cases), then we get
$T_{0}^{(\infty)}(t)=\tfrac{1}{2\pi\kappa}t^{-3/2}+\tfrac{1}{4}\left(\tfrac{\sigma^{2}}{\kappa}-1\right)+o(1)\qquad\text{as }t\to 0^{+}.$
Given that $r_{0}$ corresponds to the number of components $w_{i}$ that are finite (hence $r-r_{0}$ are the number of infinite components), then we conclude from (3.26) that
(3.38) $T_{0}(t)=\tfrac{r}{2\pi\kappa}t^{-3/2}+\tfrac{1}{4}\big{(}\tfrac{r\sigma^{2}}{\kappa}+2r_{0}-r\big{)}+o(1)\qquad\text{as }t\to 0^{+}.$
3.3.2. Two Jumps
We now simplify $T_{2}(t)$. We begin by arguing that if $\widehat{U}^{i}_{t}(t)=i$ and $\widehat{N}(t)=2$, then $\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})=1$, and therefore it follows from (3.16) that
(3.39) $\mathfrak{M}_{t}(\widehat{p},\widehat{U}^{i}_{t})=\upsilon^{2}.$
To see this, we note that when $\widehat{U}^{i}_{t}(t)=i$ and $\widehat{N}(t)=2$, the matching $\widehat{p}$ can only be $\big{\{}\{1,2\}\big{\}}$, and the two jumps in $\widehat{U}^{i}_{t}(t)$’s path must be of the form $J_{1}=(i,j)$ and $J_{2}=J_{1}^{*}=(j,i)$ for some $j\neq i$. When $\beta=1,2$, this immediately implies that $\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})=1$ by Definition 3.12-(a) and -(b). When $\beta=4$, Definition 3.12-(c) states that we must additionally argue that $\mathfrak{D}_{t}(\widehat{p},\widehat{U}^{i}_{t})=1$. For this, we note that since $J_{1}=J_{2}^{*}$, the only two binary paths in $\mathcal{B}^{0,0}_{2}$ that respect $J_{1}$ and $J_{2}$ are $(0,0,0)$ and $(0,1,0)$, and $\{1,2\}$ cannot be a flip. We then get from (3.15) that $\mathfrak{D}_{t}(\widehat{p},\widehat{U}^{i}_{t})=2^{-\widehat{N}(t)/2}\cdot 2=1$, as desired.
Next, we note that by (3.9), (3.10), and the independence of $M^{i}_{k}$ and $\{\widehat{N}(t),X^{x,x}_{t}\}$,
$\mathbf{P}\big{[}\big{\{}\widehat{U}^{i}_{t}(t)=i\big{\}}\cap\big{\{}\widehat{N}(t)=2\big{\}}~{}\big{|}~{}X^{x,x}_{t}\big{]}$ $=\mathbf{P}\big{[}\big{\{}M^{i}_{2}=i\big{\}}\cap\big{\{}\widehat{N}(t)=2\big{\}}~{}\big{|}~{}X^{x,x}_{t}\big{]}$
$=\mathbf{P}\big{[}M^{i}_{2}=i\big{]}\mathbf{P}\big{[}\widehat{N}(t)=2~{}\big{|}~{}X^{x,x}_{t}\big{]}$
$=\frac{1}{(r-1)}\cdot\frac{(r-1)^{2}\|L_{t}(X^{x,x}_{t})\|^{2}_{2}}{2}\mathrm{e}^{-\frac{(r-1)^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}}$
$=\frac{(r-1)\|L_{t}(X^{x,x}_{t})\|^{2}_{2}}{2}\mathrm{e}^{-\frac{(r-1)^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}}.$
Thus, if we denote the two-jump random walk bridge
$\tilde{U}^{i,i}_{t}=\Big{(}\widehat{U}^{i}_{t}~{}\Big{|}~{}\widehat{U}^{i}_{t}(t)=i,~{}\widehat{N}(t)=2\Big{)},$
then an application of the tower property (conditional on $X^{x,x}_{t}$) in (3.21) together with (3.39) yields
(3.40) $T_{2}(t)=\frac{(r-1)\upsilon^{2}}{2}\sum_{i=1}^{r}\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathbf{E}\Bigg{[}\|L_{t}(X^{x,x}_{t})\|^{2}_{2}\\ \cdot\exp\left(-\kappa\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s+\frac{\sigma^{2}}{2}\|L_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\|_{\mu}^{2}-\sum_{j=1}^{r}w_{j}\mathfrak{L}^{(j,0)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\right)\Bigg{]}~{}\mathrm{d}x.$
Without loss of generality (see Corollary 2.25), we can assume that
(3.41) $w_{1}<\cdots<w_{r_{0}}<\infty=w_{r_{0}+1}=\cdots=w_{r},$
keeping in mind that all or none of the $w_{i}$ could be equal to $\infty$ if $r_{0}=r$ or $r_{0}=0$. If we combine this with the heuristic (3.27), then we can rewrite (3.40) as
(3.42) $T_{2}(t)=\frac{(r-1)\upsilon^{2}}{2}\sum_{i=1}^{r}\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\Bigg{[}\|L_{t}(X^{x,x}_{t})\|^{2}_{2}\\ \cdot\exp\left(-\kappa\int_{0}^{t}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s+\frac{\sigma^{2}}{2}\|L_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\|_{\mu}^{2}-\sum_{j\leq r_{0}}w_{j}\mathfrak{L}^{(j,0)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\right)\\ \cdot\mathbf{1}_{\{\mathcal{L}^{(j,0)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})=0,~{}\forall j>r_{0}\}}\Bigg{]}~{}\mathrm{d}x.$
At this point, if we expand the exponential in the second line of (3.42) using
(3.43) $\mathrm{e}^{z}=1+\mathsf{R}_{1}(z),\qquad\text{where }|\mathsf{R}_{1}(z)|\leq|z|\mathrm{e}^{|z|},$
then we see that
(3.44) $T_{2}(t)=\frac{(r-1)\upsilon^{2}}{2}\sum_{i=1}^{r}\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\Big{[}\mathfrak{I}^{(1)}_{i,t}(x)\Big{(}1+\mathfrak{I}^{(2)}_{i,t}(x)\Big{)}\Big{]}~{}\mathrm{d}x,$
where we define
$\mathfrak{I}^{(1)}_{i,t}(x)$ $=\|L_{t}(X^{x,x}_{t})\|^{2}_{2}\mathbf{1}_{\{\mathfrak{L}^{(j,0)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})=0,~{}\forall j>r_{0}\}},$
$\mathfrak{I}^{(2)}_{i,t}(x)$ $=\mathsf{R}_{1}\left(-\kappa\int_{0}^{t}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s+\frac{\sigma^{2}}{2}\|L_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\|_{\mu}^{2}-\sum_{j\leq r_{0}}w_{j}\mathfrak{L}^{(j,0)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\right).$
We now analyze the contribution of each of these summands to $T_{2}(t)$’s asymptotics:
We begin with the contribution coming from $\mathfrak{I}^{(1)}_{i,t}(x)$. For this, we notice that
$\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\leq\mathbf{1}_{\{\mathfrak{L}^{(j,0)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})=0,~{}\forall j>r_{0}\}}\leq 1,$
where the first inequality follows from (3.14). If we combine this with both (3.32) and (3.34), then we obtain the $t\to 0^{+}$ asymptotic
(3.45) $\frac{(r-1)\upsilon^{2}}{2}\sum_{i=1}^{r}\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\Big{[}\mathfrak{I}^{(1)}_{i,t}(x)\Big{]}~{}\mathrm{d}x=\frac{r(r-1)\upsilon^{2}}{4\kappa}+o(1).$
Next, we analyze the contribution of $\mathfrak{I}^{(1)}_{i,t}(x)\mathfrak{I}^{(2)}_{i,t}(x)$. On the one hand, we note that
(3.46) $\|L_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\|_{\mu}^{2}=\sum_{j=1}^{r}\int_{0}^{\infty}\left(L^{(j,y)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\right)^{2}~{}\mathrm{d}y\\ \leq\int_{0}^{\infty}\left(\sum_{j=1}^{r}L^{(j,y)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\right)^{2}~{}\mathrm{d}y=\|L_{t}(X^{x,x}_{t})\|_{2}^{2},$
where the inequality follows from $z_{1}^{2}+\cdots+z_{n}^{2}\leq(z_{1}+\cdots+z_{n})^{2}$ whenever $z_{i}\geq 0$, and the last equality follows from (3.14). On the other hand, if
(3.47) $w_{\min}=\min\{w_{1},\ldots,w_{r}\},$
then we get from (3.14) that
(3.48) $-\sum_{j\leq r_{0}}w_{j}\mathfrak{L}^{(j,0)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})\leq-w_{\min}\sum_{j=1}^{r}\mathfrak{L}^{(j,0)}_{t}(\tilde{U}^{i,i}_{t},X^{x,x}_{t})=-w_{\min}\mathfrak{L}^{0}_{t}(X^{x,x}_{t}).$
Combining this with the trivial bound $\mathbf{1}_{E}\leq 1$ for any event $E$, we get that
(3.49) $\left|\mathfrak{I}^{(1)}_{i,t}(x)\mathfrak{I}^{(2)}_{i,t}(x)\right|\\ \leq C\|L_{t}(X^{x,x}_{t})\|^{2}_{2}\mathsf{R}_{1}\Bigg{(}\mathfrak{c}\Big{(}\big{|}{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(X^{x,x}_{t})\|_{2}^{2}+\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\Big{)}\Bigg{)}$
for some constants $C,\mathfrak{c}>0$. Given that
$\|L_{t}(X^{x,x}_{t})\|^{2}_{2}\leq\big{|}{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(X^{x,x}_{t})\|_{2}^{2}+\mathfrak{L}^{0}_{t}(X^{x,x}_{t}),$
if we combine (3.49) with the upper bound on $\mathsf{R}_{1}$ in (3.43), then we get
$\left|\mathfrak{I}^{(1)}_{i,t}(x)\mathfrak{I}^{(2)}_{i,t}(x)\right|\leq C\mathfrak{c}\Big{(}\big{|}{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(X^{x,x}_{t})\|_{2}+\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\Big{)}^{2}\\ \cdot\exp\Bigg{\{}\mathfrak{c}\Big{(}\big{|}{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(X^{x,x}_{t})\|_{2}^{2}+\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\Big{)}\Bigg{\}}.$
Thus, an application of (3.37) yields
(3.50) $\frac{(r-1)\upsilon^{2}}{2}\sum_{i=1}^{r}\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\Big{[}\mathfrak{I}^{(1)}_{i,t}(x)\mathfrak{I}^{(2)}_{i,t}(x)\Big{]}~{}\mathrm{d}x=o(1).$
Summarizing the argument for $T_{2}(t)$, if we combine (3.45) and (3.50) with (3.44), then we finally get
(3.51) $T_{2}(t)=\tfrac{r(r-1)\upsilon^{2}}{4\kappa}+o(1)\qquad\text{as }t\to 0^{+}.$
3.3.3. Four or More Jumps
We now discuss $T_{\geq 4}(t)$. For this purpose, we begin with the observation that the combinatorial constant introduced in Definition 3.12 has a very simple upper bound (see [29, (4.2)]):
$|\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})|\leq 1.$
Using essentially the same argument as in (3.46) and (3.48), we get the following:
(3.52) $\|L_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})\|_{\mu}^{2}\leq\|L_{t}(X^{x,x}_{t})\|_{2}^{2},$
(3.53) $-\sum_{j=1}^{r}w_{j}\mathfrak{L}^{(j,0)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})\leq-w_{\min}\mathfrak{L}^{0}_{t}(X^{x,x}_{t}),$
where we recall that $w_{\min}$ is defined in (3.47). Combining all of this into (3.22), we get from Jensen’s inequality (taking the absolute value inside the $\mathrm{d}x$ integral) and $|\mathfrak{M}_{t}(\widehat{p},\widehat{U}^{i}_{t})|=\upsilon^{\widehat{N}(t)}|\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})|$ that
(3.54) $|T_{\geq 4}(t)|\leq\int_{\mathcal{A}}\Pi_{X}(t;x,x)\mathbf{E}\Bigg{[}\mathrm{e}^{-\kappa\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s+\frac{(r-1)^{2}+\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}-w_{\min}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})}\\ \cdot\upsilon^{\widehat{N}(t)}\mathbf{1}_{\{\widehat{U}^{i}_{t}(t)=i\}\cap\{\widehat{N}(t)\geq 4\}}\Bigg{]}~{}\mathrm{d}\mu(i,x).$
By (3.10) and the independence of $M^{i}_{k}$ and $\{\widehat{N}(t),X^{x,x}_{t}\}$, we note that
$\mathbf{E}\left[\upsilon^{\widehat{N}(t)}\mathbf{1}_{\{\widehat{U}^{i}_{t}(t)=i\}\cap\{\widehat{N}(t)\geq 4\}}\bigg{|}X^{x,x}_{t}\right]=\frac{1}{1-r}\mathbf{E}\left[\upsilon^{\widehat{N}(t)}\mathbf{1}_{\{\widehat{N}(t)\geq 4\}}\bigg{|}X^{x,x}_{t}\right].$
Then, by (3.9),
$\frac{1}{r-1}\mathbf{E}\left[\upsilon^{\widehat{N}(t)}\mathbf{1}_{\{\widehat{N}(t)\geq 4\}}\bigg{|}X^{x,x}_{t}\right]$
$=\frac{1}{r-1}\sum_{m=2}^{\infty}\upsilon^{2m}\left(\frac{(r-1)^{2}\|L_{t}(X^{x,x}_{t})\|^{2}_{2}}{2}\right)^{m}\frac{1}{m!}\mathrm{e}^{-\frac{(r-1)^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}}$
$=\frac{\mathrm{e}^{-\frac{(r-1)^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}}}{r-1}\left(\mathrm{e}^{\frac{(r-1)^{2}\upsilon^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}}-1-\frac{(r-1)^{2}\upsilon^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}\right),$
where the last line follows from $\mathrm{e}^{z}$’s Taylor expansion. Given that $\mathrm{e}^{z}-1-z\leq z^{2}\mathrm{e}^{|z|}$ when $z\geq 0$, this simplifies to
(3.55) $\mathbf{E}\left[\upsilon^{\widehat{N}(t)}\mathbf{1}_{\{\widehat{U}^{i}_{t}(t)=i\}\cap\{\widehat{N}(t)\geq 4\}}\bigg{|}X^{x,x}_{t}\right]\leq\frac{(r-1)^{3}\upsilon^{4}\|L_{t}(X^{x,x}_{t})\|_{2}^{4}}{4}\mathrm{e}^{\frac{(\upsilon^{2}-1)(r-1)^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}}.$
Therefore, an application of the tower property in the left-hand side of (3.54) yields
(3.56) $|T_{\geq 4}(t)|\leq\frac{r(r-1)^{3}\upsilon^{4}}{4}\int_{\mathcal{A}}\Pi_{X}(t;x,x)\mathbf{E}\Bigg{[}\|L_{t}(X^{x,x}_{t})\|_{2}^{4}\\ \cdot\mathrm{e}^{-\kappa\int_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s+\frac{\upsilon^{2}(r-1)^{2}+\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}-w_{\min}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})}\Bigg{]}~{}\mathrm{d}\mu(i,x).$
On the one hand, we note that there exists a large enough constant $\mathfrak{c}>0$ such that
$\left(-\kappa{\textstyle{\int}_{0}^{t}X^{x,x}_{t}(s)~{}\mathrm{d}s+\frac{\upsilon^{2}(r-1)^{2}+\sigma^{2}}{2}\|L_{t}(X^{x,x}_{t})\|_{2}^{2}-w_{\min}\mathfrak{L}^{0}_{t}(X^{x,x}_{t})}\right)\\ \leq\mathfrak{c}\Big{(}\big{|}{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(X^{x,x}_{t})\|_{2}^{2}+\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\Big{)}.$
On the other hand,
$\|L_{t}(X^{x,x}_{t})\|_{2}^{4}\leq\Big{(}\big{|}{\textstyle\int_{0}^{t}}\big{(}X^{x,x}_{t}(s)-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(X^{x,x}_{t})\|_{2}^{2}+\mathfrak{L}^{0}_{t}(X^{x,x}_{t})\Big{)}^{2}.$
If we plug both of these estimates into (3.56), then we get from (3.37) that
(3.57) $T_{\geq 4}(t)=o(1)\qquad\text{as }t\to 0^{+}.$
3.3.4. Summary
Before moving on, we briefly mention that the combination of (3.38) (3.51), and (3.57) implies the asymptotic (3.1). Consequently, Theorem 3.1 now follows from Proposition 3.16. We now take on this final task.
3.4. Proof of (3.1) Step 3. Reflected Brownian Bridge Asymptotics
We now conclude this paper by proving Proposition 3.16.
3.4.1. Preliminary: Brownian Motion Coupling
Reflected Brownian motion bridges are more difficult to work with than Brownian bridges. Thus, we begin the proof of Proposition 3.16 by examining how one can write (or estimate) expectations of functions of $X^{x,x}_{t}$ as expectations of functions $B^{x,x}_{t}$ and/or $B^{x,-x}_{t}$ instead. For this purpose, we recall that in Definition 3.5, we couple $X$ with a standard Brownian motion $B$ via $X=|B|$. Under this coupling, we note that $X^{x,x}_{t}$ and a conditioned version of $B^{x}$ are related as follows:
$X^{x,x}_{t}(s)=\big{(}|B^{x}(s)|~{}\big{|}~{}B^{x}(t)\in\{-x,x\}\big{)},\qquad s\in[0,t].$
Given that
$\mathbf{P}\big{[}B^{x}(t)=\pm x~{}\big{|}~{}B^{x}(t)\in\{-x,x\}\big{]}=\frac{\Pi_{B}(t;x,\pm x)}{\Pi_{B}(t;x,x)+\Pi_{B}(t;x,-x)}=\frac{\Pi_{B}(t;x,\pm x)}{\Pi_{X}(t;x,x)}$
and $\Pi_{B}(t;x,x)=\frac{1}{\sqrt{2\pi t}}$, $\Pi_{B}(t;x,-x)=\frac{\mathrm{e}^{-2x^{2}/t}}{\sqrt{2\pi t}}$, for any functional $F$, we can write
(3.58) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}F\big{(}X^{x,x}_{t}\big{)}\big{]}~{}\mathrm{d}x\\ =\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}\mathbf{E}\big{[}F\big{(}|B^{x,x}_{t}|\big{)}\big{]}~{}\mathrm{d}x+\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx-2x^{2}/t}}{\sqrt{2\pi t}}\mathbf{E}\big{[}F\big{(}|B^{x,-x}_{t}|\big{)}\big{]}~{}\mathrm{d}x.$
If we replace $F(X^{x,x}_{t})$ by $F(X^{x,x}_{t})\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}$, then (3.58) can be further simplified. Indeed, given that $B^{x,-x}_{t}$ must obviously reach zero as it travels from $x$ to $-x$, one has $\mathbf{1}_{\{\mathfrak{L}^{0}(|B^{x,-x}_{t}|)=0\}}=0$. Thus, in that case (3.58) simplifies to
(3.59) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}F\big{(}X^{x,x}_{t}\big{)}\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\big{]}~{}\mathrm{d}x\\ =\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}\mathbf{E}\left[F\big{(}|B^{x,x}_{t}|\big{)}\mathbf{1}_{\{\mathfrak{L}^{0}(|B^{x,x}_{t}|)=0\}}\right]~{}\mathrm{d}x.$
Next, if we write
$\mathbf{1}_{\{\mathfrak{L}^{0}(|B^{x,x}_{t}|)=0\}}=\mathbf{1}_{\{\forall s\leq t,~{}B^{x,x}_{t}(s)\neq 0\}}=1-\mathbf{1}_{\{\exists s\leq t:~{}B^{x,x}_{t}(s)=0\}},$
then we can write the right-hand side of (3.59) as
(3.60) $\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}\mathbf{E}\left[F\big{(}|B^{x,x}_{t}|\big{)}\right]~{}\mathrm{d}x-\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}\mathbf{E}\left[F\big{(}|B^{x,x}_{t}|\big{)}\mathbf{1}_{\{\exists s\leq t:~{}B^{x,x}_{t}(s)=0\}}\right]~{}\mathrm{d}x.$
In order to simplify the second term in (3.60), we combine the following observations:
  1. (1)
    By the strong Markov property and the symmetry of Brownian motion about 00, we notice that
    $\big{(}|B^{x,x}_{t}|~{}\big{|}~{}\exists s\leq t:~{}B^{x,x}_{t}(s)=0\big{)}\stackrel{{\scriptstyle\mathrm{distr.}}}{{=}}|B^{x,-x}_{t}|$
    (i.e., we can couple the two processes above by reflecting the path of $B^{x,x}_{t}$ after its first passage to zero, thus making it end at $-x$ instead of $x$).
  2. (2)
    By using the joint density of the running minimum and the current value of a Brownian motion (e.g., [54, Chapter III, Exercise 3.14]), we obtain that
    $\mathbf{P}\big{[}\exists s\leq t:~{}B^{x,x}_{t}(s)=0\big{]}=\mathrm{e}^{-2x^{2}/t}.$
Thus, if we turn the second expectation in (3.60) into a conditional expectation given $\{\exists s\leq t:~{}B^{x,x}_{t}(s)=0\}$ (by multiplying and dividing by the probability of that event, i.e., $\mathrm{e}^{-2x^{2}/t}$), then we obtain the following counterpart to (3.58)
(3.61) $\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}F\big{(}X^{x,x}_{t}\big{)}\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\big{]}~{}\mathrm{d}x\\ =\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}\mathbf{E}\big{[}F\big{(}|B^{x,x}_{t}|\big{)}\big{]}~{}\mathrm{d}x-\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx-2x^{2}/t}}{\sqrt{2\pi t}}\mathbf{E}\big{[}F\big{(}|B^{x,-x}_{t}|\big{)}\big{]}~{}\mathrm{d}x.$
Finally, while (3.58) and (3.61) are useful for exact calculations (and thus necessary to obtain the explicit leading order terms in our asymptotics), simpler expressions can be used when we are only interested in upper estimates. For this purpose, we invoke [28, (5.9)], which states that for any nonnegative functional $F$, one has
(3.62) $\mathbf{E}\big{[}F(X^{x,x}_{t})\big{]}\leq 2\mathbf{E}\big{[}F(|B^{x,x}_{t}|)\big{]}.$
We are now prepared to undertake the proofs of our technical results.
3.4.2. Proof of (3.31) and (3.33)
By (3.58) and (3.61) with $F=1$,
$\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}~{}\mathrm{d}x=\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}~{}\mathrm{d}x+\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx-2x^{2}/t}}{\sqrt{2\pi t}}~{}\mathrm{d}x,$
and
$\int_{0}^{\infty}\Pi_{X}(t;x,x)\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}\big{]}~{}\mathrm{d}x=\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}~{}\mathrm{d}x-\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx-2x^{2}/t}}{\sqrt{2\pi t}}~{}\mathrm{d}x.$
We then get (3.31) and (3.33) as follows: On the one hand,
$\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}~{}\mathrm{d}x=\frac{1}{\sqrt{2\pi}\kappa}t^{-3/2}.$
On the other hand, by the change of variables $x\mapsto t^{1/2}x$ and dominated convergence,
$\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx-2x^{2}/t}}{\sqrt{2\pi t}}~{}\mathrm{d}x=\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa t^{3/2}x-2x^{2}}}{\sqrt{2\pi}}~{}\mathrm{d}x=\int_{0}^{\infty}\frac{\mathrm{e}^{-2x^{2}}}{\sqrt{2\pi}}~{}\mathrm{d}x+o(1)=\frac{1}{4}+o(1).$
3.4.3. Proof of (3.32) and (3.34)
By (3.58) and (3.61), the proof of (3.32) and (3.34) reduces to two claims: As $t\to 0^{+}$, one has
(3.63) $\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}\mathbf{E}\big{[}\|L_{t}(|B^{x,x}_{t}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x=\tfrac{1}{2\kappa}+o(1),$
(3.64) $\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx-2x^{2}/t}}{\sqrt{2\pi t}}\mathbf{E}\big{[}\|L_{t}(|B^{x,-x}_{t}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x=o(1).$
For this purpose, it is useful to note that, by Brownian scaling and the change of variables $y\mapsto t^{1/2}y$,
(3.65) $\|L_{t}(|B^{x,x}_{t}|)\|_{2}^{2}=\int_{0}^{\infty}L^{y}_{t}(|B^{x,\pm x}_{t}|)^{2}~{}\mathrm{d}y=t\int_{0}^{\infty}L^{t^{-1/2}y}_{1}(|B^{t^{-1/2}x,\pm t^{-1/2}x}_{1}|)^{2}~{}\mathrm{d}y\\ =t^{3/2}\int_{0}^{\infty}L^{y}_{1}(|B^{t^{-1/2}x,\pm t^{-1/2}x}_{1}|)^{2}~{}\mathrm{d}y=t^{3/2}\|L_{1}(|B^{t^{-1/2}x,\pm t^{-1/2}x}|)\|_{2}^{2}.$
We begin with the proof of (3.63). An application of (3.65) yields
$\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}\mathbf{E}\big{[}\|L_{t}(|B^{x,x}_{t}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x=\int_{0}^{\infty}\frac{t\,\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi}}\mathbf{E}\big{[}\|L_{1}(|B^{t^{-1/2}x,t^{-1/2}x}_{1}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x.$
If we now apply the change of variables $x\mapsto t^{-1}x$, then this becomes
$\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}\mathbf{E}\big{[}\|L_{t}(|B^{x,x}_{t}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x=\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa x}}{\sqrt{2\pi}}\mathbf{E}\big{[}\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x.$
The proof of (3.63) can thus be broken down into two steps:
(3.66) $\lim_{t\to 0^{+}}\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa x}}{\sqrt{2\pi}}\mathbf{E}\big{[}\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x\\ =\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa x}}{\sqrt{2\pi}}\mathbf{E}\big{[}\|L_{1}(B^{0,0}_{1})\|_{2}^{2}\big{]}~{}\mathrm{d}x=\frac{\mathbf{E}\big{[}\|L_{1}(B^{0,0}_{1})\|_{2}^{2}\big{]}}{\sqrt{2\pi}\kappa}$
and
(3.67) $\mathbf{E}\big{[}\|L_{1}(B^{0,0}_{1})\|_{2}^{2}\big{]}=\tfrac{\sqrt{\pi}}{\sqrt{2}}.$
For (3.66), we first notice that for any $y>0$,
$L_{1}^{y}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)=L_{1}^{y}(B^{t^{-3/2}x,t^{-3/2}x}_{1})+L_{1}^{-y}(B^{t^{-3/2}x,t^{-3/2}x}_{1}).$
Moreover, if we use the coupling
$B^{t^{-3/2}x,t^{-3/2}x}_{1}=t^{-3/2}x+B^{0,0}_{1},$
then this becomes
(3.68) $L_{1}^{y}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)=L_{1}^{y-t^{-3/2}x}(B^{0,0}_{1})+L_{1}^{-y-t^{-3/2}x}(B^{0,0}_{1}).$
Therefore,
(3.69) $\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}=\int_{0}^{\infty}\left(L_{1}^{y-t^{-3/2}x}(B^{0,0}_{1})+L_{1}^{-y-t^{-3/2}x}(B^{0,0}_{1})\right)^{2}~{}\mathrm{d}y.$
For every fixed $x>0$, there exists some finite random $t_{0}$ small enough so that $B^{0,0}_{1}$’s path does not intersect $(-\infty,-t^{-3/2}x]$ for every $t<t_{0}$. Therefore,
$\lim_{t\to 0^{+}}\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}=\lim_{t\to 0^{+}}\int_{0}^{\infty}L_{1}^{y-t^{-3/2}x}(B^{0,0}_{1})^{2}~{}\mathrm{d}y=\|L_{1}(B^{0,0}_{1})\|_{2}^{2}\qquad\text{almost surely.}$.
Consequently, (3.66) follows from dominated convergence if we prove that
$\sup_{x>0}\mathbf{E}\left[\sup_{t>0}\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}\right]<\infty.$
This immediately follows by combining (3.67) with an application of the inequality $(z_{1}+z_{1})^{2}\leq 2(z_{1}^{2}+z_{2}^{2})$ to (3.69), the latter of which yields
$\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}\leq 2\int_{0}^{\infty}L_{1}^{y-t^{-3/2}x}(B^{0,0}_{1})^{2}+L_{1}^{-y-t^{-3/2}x}(B^{0,0}_{1})^{2}~{}\mathrm{d}y=2\|L_{t}(B^{0,0}_{1})\|_{2}^{2}.$
With this in hand, the only element left in the proof of (3.63) is (3.67). By Tonelli’s theorem, we can write
$\mathbf{E}\big{[}\|L_{t}(B^{0,0}_{1})\|_{2}^{2}\big{]}=\int_{\mathbb{R}}\mathbf{E}\big{[}L^{y}_{1}(B^{0,0}_{1})^{2}\big{]}~{}\mathrm{d}x.$
By [52, (1)], we have the joint density
(3.70) $\mathbf{P}\left[L_{1}^{y}(B^{0})\in\mathrm{d}\ell,B^{0}(1)\in\mathrm{d}x\right]=\frac{(|y|+|x-y|+\ell)\mathrm{e}^{-(|y|+|x-y|+\ell)^{2}/2}}{\sqrt{2\pi}}$
for all $\ell>0$ and $x,y\in\mathbb{R}$. If we evaluate this at $x=0$, further noting that $\mathbf{P}\left[B^{0}(1)\in\mathrm{d}0\right]=\Pi_{B}(1;0,0)=\frac{1}{\sqrt{2\pi}},$ then we get the conditional density
$\mathbf{P}\left[L_{1}^{y}(B^{0,0}_{1})\in\mathrm{d}\ell\right]=(2|y|+\ell)\mathrm{e}^{-(2|y|+\ell)^{2}/2}.$
Therefore, we obtain (3.67) through a sequence of two exact calculations:
$\mathbf{E}\left[L^{y}_{1}(B^{0,0}_{1})^{2}\right]=\int_{0}^{\infty}\ell^{2}\cdot(2|y|+\ell)\mathrm{e}^{-(2|y|+\ell)^{2}/2}~{}\mathrm{d}\ell=2\mathrm{e}^{-2y^{2}}-2\sqrt{2\pi}|y|\,\mathrm{erfc}(\sqrt{2}|y|),$
and then
$\int_{\mathbb{R}}2\mathrm{e}^{-2y^{2}}-2\sqrt{2\pi}|y|\,\mathrm{erfc}(\sqrt{2}|y|)~{}\mathrm{d}y=\tfrac{\sqrt{\pi}}{\sqrt{2}},$
where we recall that $\mathrm{erfc}(z)=1-\frac{2}{\sqrt{\pi}}\int_{0}^{z}\mathrm{e}^{-y^{2}}~{}\mathrm{d}y$ is the complementary error function.
Now that (3.63) is proved, the proof of (3.32) and (3.34) relies on establishing (3.64). By (3.65), this is equivalent to showing
$t\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx-2x^{2}/t}}{\sqrt{2\pi}}\mathbf{E}\big{[}\|L_{1}(|B^{t^{-1/2}x,-t^{-1/2}x}_{1}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x=o(1).$
If we apply the change of variables $x\mapsto t^{1/2}x$, then this becomes
$t^{3/2}\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa t^{3/2}x-2x^{2}}}{\sqrt{2\pi}}\mathbf{E}\big{[}\|L_{1}(|B^{x,-x}_{1}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x=o(1).$
As $\mathrm{e}^{-\kappa t^{3/2}x}\to 1$ as $t\to 0$, by dominated convergence, we get (3.63) if we show that
(3.71) $\int_{0}^{\infty}\frac{\mathrm{e}^{-2x^{2}}}{\sqrt{2\pi}}\mathbf{E}\big{[}\|L_{1}(|B^{x,-x}_{1}|)\|_{2}^{2}\big{]}~{}\mathrm{d}x<\infty.$
Using a similar argument as in (3.68), we get
$\|L_{1}(|B^{x,-x}_{1}|)\|_{2}^{2}=\int_{0}^{\infty}\Big{(}L^{y}_{1}(B^{x,-x}_{1})+L^{-y}_{1}(B^{x,-x}_{1})\Big{)}^{2}~{}\mathrm{d}y\\ \leq 2\int_{0}^{\infty}L^{y}_{1}(B^{x,-x}_{1})^{2}+L^{-y}_{1}(B^{x,-x}_{1})^{2}~{}\mathrm{d}y=2\|L_{t}(B^{x,-x})\|_{2}^{2}.$
If we additionally use the fact that self-intersection local time is invariant with respect to shifts in space, we get $\|L_{t}(B^{x,-x})\|_{2}^{2}=\|L_{t}(B^{0,-2x})\|_{2}^{2}$. Thus (3.71) reduces to
(3.72) $\int_{0}^{\infty}\frac{\mathrm{e}^{-2x^{2}}}{\sqrt{2\pi}}\mathbf{E}\big{[}\|L_{1}(B^{0,-2x}_{1})\|_{2}^{2}\big{]}~{}\mathrm{d}x<\infty.$
If we combine (3.70) with $\mathbf{P}[B^{0}(1)\in\mathrm{d}(-2x)]=\Pi_{B}(1;0,-2x)=\frac{\mathrm{e}^{-2x^{2}}}{\sqrt{2\pi}}$, then we get that
$\mathbf{P}\left[L_{1}^{y}(B^{0,-2x}_{1})\in\mathrm{d}\ell\right]=(|y|+|-2x-y|+\ell)\mathrm{e}^{-(|y|+|-2x-y|+\ell)^{2}/2}\mathrm{e}^{2x^{2}}.$
We then get (3.72) through a sequence of exact calculations: Firstly,
(3.73) $\mathbf{E}\big{[}L^{y}_{1}(B^{0,-2x}_{1})^{2}\big{]}=\int_{0}^{\infty}\ell^{2}\cdot(|y|+|-2x-y|+\ell)\mathrm{e}^{-(|y|+|-2x-y|+\ell)^{2}/2}\mathrm{e}^{2x^{2}}~{}\mathrm{d}\ell\\ =\mathrm{e}^{2x^{2}}\left(\sqrt{2\pi}(|2x+y|+|y|)\left(\mathrm{erf}\left(\frac{|2x+y|+|y|}{\sqrt{2}}\right)-1\right)+2\mathrm{e}^{-(|2x+y|+|y|)^{2}/2}\right),$
where $\mathrm{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}\mathrm{e}^{-y^{2}}~{}\mathrm{d}y$ denotes the error function. Secondly, integrating the second line of (3.73) with respect to $y$ on all of $\mathbb{R}$ yields
(3.74) $\mathbf{E}\big{[}\|L_{1}(B^{0,-2x}_{1})\|_{2}^{2}\big{]}=4\sqrt{2\pi}\mathrm{e}^{2x^{2}}x^{2}\mathrm{erf}(\sqrt{2}x)\\ +\tfrac{\sqrt{\pi}}{\sqrt{2}}\mathrm{e}^{2x^{2}}(4x^{2}+1)\mathrm{erfc}(\sqrt{2}x)+2(1-2\sqrt{2\pi}\mathrm{e}^{2x^{2}}x)x.$
Thirdly, integrating the right-hand side of (3.74) multiplied by $\frac{\mathrm{e}^{-2x^{2}}}{\sqrt{2\pi}}$ finally yields
$\int_{0}^{\infty}\frac{\mathrm{e}^{-2x^{2}}}{\sqrt{2\pi}}\mathbf{E}\big{[}\|L_{1}(B^{0,-2x}_{1})\|_{2}^{2}\big{]}~{}\mathrm{d}x=\tfrac{\sqrt{2}}{3\sqrt{\pi}}<\infty.$
This proves (3.72), thus concluding the proof of (3.32) and (3.34).
3.4.4. Proof of (3.35)
By (3.58), we can split (3.35) into two parts:
(3.75) $\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx}}{\sqrt{2\pi t}}\mathbf{E}\left[\int_{0}^{t}|B^{x,x}_{t}(s)|-x~{}\mathrm{d}s\right]~{}\mathrm{d}x=o(1),$
(3.76) $\int_{0}^{\infty}\frac{\mathrm{e}^{-\kappa tx-2x^{2}/t}}{\sqrt{2\pi t}}\mathbf{E}\left[\int_{0}^{t}|B^{x,-x}_{t}(s)|-x~{}\mathrm{d}s\right]~{}\mathrm{d}x=o(1).$
We begin with (3.75). If we couple $B^{x,x}_{t}=x+B^{0,0}_{t}$ and then distribute the $t^{-1/2}$ term in $\frac{1}{\sqrt{2\pi t}}$ inside the integral, then we get that (RHS is short for right-hand side)
[…]
Then, by a Brownian scaling and the changes of variables $s\mapsto t^{-1}s$ and $x\mapsto t^{-1/2}x$,
[…]
At this point, if we show that
(3.77) $\int_{0}^{\infty}\left|\mathbf{E}\left[\int_{0}^{1}|x+B^{0,0}_{1}(s)|-x~{}\mathrm{d}s\right]\right|~{}\mathrm{d}x=\int_{0}^{\infty}\left|\int_{0}^{1}\mathbf{E}\big{[}|x+B^{0,0}_{1}(s)|-x\big{]}~{}\mathrm{d}s\right|~{}\mathrm{d}x<\infty,$
then we get (3.75) (actually, the stronger statement RHS of (3.75) $=O(t^{3/2})$) by dominated convergence since $\mathrm{e}^{-\kappa t^{3/2}x}\uparrow 1$ as $t\to 0^{+}$. Toward this end, since $B^{0,0}_{1}(s)$ is Gaussian with mean zero and variance $s(1-s)$, a direct calculation yields
$\mathbf{E}\big{[}|x+B^{0,0}_{1}(s)|-x\big{]}=\sqrt{\frac{2s(1-s)}{\pi}}\mathrm{e}^{-x^{2}/2s(1-s)}-x\,\mathrm{erfc}\left(\frac{x}{\sqrt{2s(1-s)}}\right).$
Thus, by Jensen’s inequality (taking the absolute value inside the $\mathrm{d}s$ integral), Tonelli’s theorem (interchanging the $\mathrm{d}x$ and $\mathrm{d}s$ integrals), and direct calculations,
[…]
Thus proves (3.77), and therefore also (3.75).
We now move on to (3.76). For this, we use the coupling
$B^{x,-x}_{t}(s)=(1-\tfrac{s}{t})x-\tfrac{s}{t}x+B^{0,0}_{t}(s).$
By the reverse triangle inequality and the triangle inequality, for any $x>0$ and $s\in[0,t]$, one has
(3.79) $\big{|}|B^{x,-x}_{t}(s)|-x\big{|}=\big{|}|(1-\tfrac{s}{t})x-\tfrac{s}{t}x+B^{0,0}_{t}(s)|-x\big{|}\\ \leq\big{|}(1-\tfrac{s}{t})x-\tfrac{s}{t}x+B^{0,0}_{t}(s)-x\big{|}=|B^{0,0}_{t}(s)-\tfrac{2s}{t}x|\leq|B^{0,0}_{t}(s)|+\tfrac{2s}{t}x.$
Combining this with $\mathrm{e}^{-\kappa tx}\leq 1$, we get the upper bounds
[…]
$\leq\int_{0}^{\infty}\frac{\mathrm{e}^{-2x^{2}/t}}{\sqrt{2\pi t}}\cdot t\left(\mathbf{E}\left[\sup_{0\leq s\leq t}|B^{0,0}_{t}(s)|\right]+x\right)~{}\mathrm{d}x$
$=\frac{t}{4}\mathbf{E}\left[\sup_{0\leq s\leq t}|B^{0,0}_{t}(s)|\right]+\frac{t^{3/2}}{4\sqrt{2\pi}}$
$=\frac{t^{3/2}}{4}\left(\mathbf{E}\left[\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|\right]+\frac{1}{\sqrt{2\pi}}\right),$
where the last line follows by Brownian scaling. This proves (3.76).
3.4.5. Proof of (3.36)
By (3.58), (3.36) reduces to
(3.80) $\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\Pi_{B}(t;x,\pm x)\mathbf{E}\big{[}\mathfrak{L}^{0}_{t}(B^{x,\pm x}_{t})\big{]}~{}\mathrm{d}x=o(1).$
(Note that we use $B^{x,\pm x}_{t}$ instead of $|B^{x,\pm x}_{t}|$ above because both processes have the same local time at zero.) By Brownian scaling,
$\mathbf{E}\big{[}\mathfrak{L}^{0}_{t}(B^{x,\pm x}_{t})\big{]}=t^{1/2}\mathbf{E}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,\pm t^{-1/2}x}_{1})\big{]}$
Thanks to [52, (1)], for every $x,y\in\mathbb{R}$ and $\ell>0$, we have that
$\mathbf{P}[\mathfrak{L}^{0}_{1}(B^{x})\in\mathrm{d}\ell,B^{x}(1)\in\mathrm{d}y]=\frac{(|x|+|y|+\ell)}{\sqrt{2\pi}}\mathrm{e}^{-(|x|+|y|+\ell)^{2}/2}.$
If we now turn the above into a conditional density given $B^{x}(1)=y$, then we get
$\mathbf{P}\big{[}\mathfrak{L}^{0}_{1}(B^{x,y}_{1})\in\mathrm{d}\ell\big{]}=\Pi_{B}(1;x,y)^{-1}(|x|+|y|+\ell)\mathrm{e}^{-(|x|+|y|+\ell)^{2}/2}.$
By Brownian scaling,
$\Pi_{B}(t;x,y)=\frac{\mathrm{e}^{-(x-y)^{2}/2t}}{\sqrt{2\pi t}}=t^{-1/2}\frac{\mathrm{e}^{-(t^{-1/2}x-t^{-1/2}y)^{2}/2}}{\sqrt{2\pi}}=t^{-1/2}\Pi_{B}(1;t^{-1/2}x,t^{-1/2}y);$
therefore, we obtain that
$\Pi_{B}(t;x,y)t^{1/2}\mathbf{P}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}y}_{1})\in\mathrm{d}\ell\big{]}=(|t^{-1/2}x|+|t^{-1/2}y|+\ell)\mathrm{e}^{-(|t^{-1/2}x|+|t^{-1/2}y|+\ell)^{2}/2}.$
With $y=\pm x$ for $x>0$, this then simplifies to
(3.81) $\Pi_{B}(t;x,\pm x)t^{1/2}\mathbf{P}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,\pm t^{-1/2}x}_{1})\in\mathrm{d}\ell\big{]}=(2t^{-1/2}x+\ell)\mathrm{e}^{-(2t^{-1/2}x+\ell)^{2}/2}.$
Thus, an exact calculation yields
(3.82) $\Pi_{B}(t;x,\pm x)t^{1/2}\mathbf{E}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}\pm x}_{1})\big{]}\\ =\int_{0}^{\infty}\ell\cdot(2t^{-1/2}x+\ell)\mathrm{e}^{-(2t^{-1/2}x+\ell)^{2}/2}~{}\mathrm{d}\ell=\frac{\sqrt{\pi}}{\sqrt{2}}\mathrm{erfc}\left(\sqrt{2}t^{-1/2}x\right).$
Consequently, another exact calculation yields
[…]
We now conclude the proof of (3.80) thanks to the asymptotic
$\frac{\sqrt{\frac{\pi}{2}}\left(1-e^{\kappa^{2}t^{3}/8}\text{erfc}\left(\frac{\kappa t^{3/2}}{2\sqrt{2}}\right)\right)}{\kappa t}=O(t^{1/2})\qquad\text{as }t\to 0^{+},$
which follows from the product of the Taylor expansions $\mathrm{erfc}(z)=1-\frac{2z}{\sqrt{\pi}}+O(z^{2})$ and $\mathrm{e}^{z}=1+O(z)$ as $z\to 0$.
3.4.6. Proof of (3.37)
With (3.31)–(3.36) proved, the only estimate left to prove in the paper is (3.37). If we combine $\Pi_{X}(t;x,x)\leq\frac{2}{\sqrt{2\pi t}}$ and (3.62), then (3.37) reduces to
(3.83) $t^{-1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\Bigg{[}\Big{(}\big{|}{\textstyle\int_{0}^{t}}\big{(}|B^{x,x}_{t}(s)|-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(|B^{x,x}_{t}|)\|_{2}+\mathfrak{L}^{0}_{t}(B^{x,x}_{t})\Big{)}^{2}\\ \cdot\exp\Bigg{\{}\mathfrak{c}\Big{(}\big{|}{\textstyle\int_{0}^{t}}\big{(}|B^{x,x}_{t}(s)|-x\big{)}~{}\mathrm{d}s\big{|}+\|L_{t}(|B^{x,x}_{t}|)\|_{2}^{2}+\mathfrak{L}^{0}_{t}(B^{x,x}_{t})\Big{)}\Bigg{\}}\Bigg{]}~{}\mathrm{d}x=o(1),$
where there is no absolute value on $B^{x,x}_{t}$ in the boundary local time $\mathfrak{L}^{0}_{t}$, since $B^{x,x}_{t}$ and $|B^{x,x}_{t}|$ both have the same local time at zero. If we use the coupling $B^{x,x}_{t}=x+B^{0,0}_{t}$ and the reverse triangle inequality, then we get
$\big{|}{\textstyle\int_{0}^{t}}\big{(}|B^{x,x}_{t}(s)|-x\big{)}~{}\mathrm{d}s\big{|}\leq{\textstyle\int_{0}^{t}}\big{|}|x+B^{0,0}_{t}(s)|-x\big{|}~{}\mathrm{d}s\leq t\sup_{0\leq s\leq t}|B^{0,0}_{t}(s)|.$
If we combine this with an application of Hölder’s inequality, then (3.83) reduces to
$t^{-1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\Bigg{[}\Big{(}t\sup_{0\leq s\leq t}|B^{0,0}_{t}(s)|+\|L_{t}(|B^{x,x}_{t}|)\|_{2}+\mathfrak{L}^{0}_{t}(B^{x,x}_{t})\Big{)}^{4}\Bigg{]}^{1/2}\\ \cdot\mathbf{E}\Bigg{[}\exp\Bigg{\{}2\mathfrak{c}\Big{(}t\sup_{0\leq s\leq t}|B^{0,0}_{t}(s)|+\|L_{t}(|B^{x,x}_{t}|)\|_{2}^{2}+\mathfrak{L}^{0}_{t}(B^{x,x}_{t})\Big{)}\Bigg{\}}\Bigg{]}^{1/2}~{}\mathrm{d}x=o(1).$
If we then apply a Brownian scaling, then we get the further reduction
(3.84) $t^{-1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\Bigg{[}\Big{(}t^{3/2}\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|+t^{3/2}\|L_{1}(|B^{x_{t},x_{t}}_{1}|)\|_{2}+t^{1/2}\mathfrak{L}^{0}_{1}(B^{x_{t},x_{t}}_{1})\Big{)}^{4}\Bigg{]}^{1/2}\\ \cdot\mathbf{E}\Bigg{[}\exp\Bigg{\{}2\mathfrak{c}\Big{(}t^{3/2}\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|+t^{3/2}\|L_{1}(|B^{x_{t},x_{t}}_{1}|)\|_{2}^{2}+t^{1/2}\mathfrak{L}^{0}_{1}(B^{x_{t},x_{t}}_{1})\Big{)}\Bigg{\}}\Bigg{]}^{1/2}~{}\mathrm{d}x=o(1),$
where we denote $x_{t}=t^{-1/2}x$ for simplicity.
By Hölder’s inequality,
$\sup_{x>0,~{}t\in(0,1]}\mathbf{E}\Bigg{[}\exp\Bigg{\{}2\mathfrak{c}\Big{(}t^{3/2}\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|+t^{3/2}\|L_{1}(|B^{x_{t},x_{t}}_{1}|)\|_{2}^{2}+t^{1/2}\mathfrak{L}^{0}_{1}(B^{x_{t},x_{t}}_{1})\Big{)}\Bigg{\}}\Bigg{]}^{1/2}$
(3.85) $\leq\mathbf{E}\left[\mathrm{e}^{6\mathfrak{c}\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|}\right]^{1/6}\sup_{z>0}\mathbf{E}\left[\mathrm{e}^{6\mathfrak{c}\|L_{1}(|B^{z,z}_{1}|)\|_{2}^{2}}\right]^{1/6}\sup_{z>0}\mathbf{E}\left[\mathrm{e}^{6\mathfrak{c}\mathfrak{L}^{0}_{1}(B^{z,z}_{1})}\right]^{1/6}.$
Since $|B^{0,0}_{1}|$ is a Bessel bridge of dimension 1, the first expectation in the second line of (3.85) is finite thanks to the Gaussian tail bound for Bessel bridge maxima in [40, Remark 3.1]. The third expectation in the second line of (3.85) is finite thanks to [28, Lemma 5.8]. Finally, if we use the bound $(z_{1}+z_{2})^{2}\leq 2(z_{1}^{2}+z_{2}^{2})$, then we see that
$\|L_{1}(|B^{z,z}_{1}|)\|_{2}^{2}=\int_{0}^{\infty}L^{y}_{1}(|B^{z,z}_{1}|)^{2}~{}\mathrm{d}y=\int_{0}^{\infty}\Big{(}L^{y}_{1}(B^{z,z}_{1})+L^{-y}_{1}(B^{z,z}_{1})\Big{)}^{2}~{}\mathrm{d}y\\ \leq 2\int_{0}^{\infty}L^{y}_{1}(B^{z,z}_{1})^{2}+L^{-y}_{1}(B^{z,z}_{1})^{2}~{}\mathrm{d}y=2\|L_{1}(B^{z,z}_{t})\|_{2}^{2},$
hence
(3.86) $\sup_{z>0}\mathbf{E}\left[\mathrm{e}^{6\mathfrak{c}\|L_{1}(|B^{z,z}_{1}|)\|_{2}^{2}}\right]^{1/6}\leq\sup_{z>0}\mathbf{E}\left[\mathrm{e}^{12\mathfrak{c}\|L_{1}(B^{z,z}_{1})\|_{2}^{2}}\right]^{1/6}<\infty,$
where the last inequality is from [28, Lemma 5.11]. In summary, (3.85) is finite, which means that we can now reduce (3.84) to
(3.87) $t^{-1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\Bigg{[}\Big{(}t^{3/2}\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|+t^{3/2}\|L_{1}(|B^{x_{t},x_{t}}_{1}|)\|_{2}^{2}+t^{1/2}\mathfrak{L}^{0}_{1}(B^{x_{t},x_{t}}_{1})\Big{)}^{4}\Bigg{]}^{1/2}~{}\mathrm{d}x=o(1).$
Since there exists some $c>0$ such that $(z_{1}+z_{2}+z_{3})^{4}\leq c(z_{1}^{4}+z_{2}^{4}+z_{3}^{4})$ for all $z_{i}\geq 0$ (by Jensen’s inequality), and $\sqrt{z_{1}+z_{2}+z_{3}}\leq\sqrt{z}_{1}+\sqrt{z}_{2}+\sqrt{z}_{3}$ for all $z_{i}\geq 0$, we can further reduce (3.87) to
(3.88) $t^{-1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\Bigg{(}\mathbf{E}\Big{[}t^{6}\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|^{4}\Big{]}^{1/2}+\mathbf{E}\Big{[}t^{6}\|L_{1}(|B^{x_{t},x_{t}}_{1}|)\|_{2}^{8}\Big{]}^{1/2}\\ +\mathbf{E}\big{[}t^{2}\mathfrak{L}^{0}_{1}(B^{x_{t},x_{t}}_{1})^{4}\big{]}^{1/2}\Bigg{)}=o(1).$
The first line of (3.88) vanishes thanks to the combination of three facts: Firstly,
$t^{-1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}(t^{6})^{1/2}~{}\mathrm{d}x=\frac{t^{3/2}}{\kappa}=o(1).$
Secondly,
$\mathbf{E}\Big{[}\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|^{4}\Big{]}<\infty$
thanks to the fact that $|B^{0,0}_{1}|$ supremum has Gaussian tails ([40, Remark 3.1]). Thirdly, it follows from (3.86) that
$\sup_{z>0}\mathbf{E}\big{[}\|L_{1}(|B^{z,z}_{1}|)\|_{2}^{8}\big{]}<\infty.$
Thus, we now only need to show that the second line of (3.88) vanishes, which we simplify as as follows:
(3.89) $t^{1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}x}_{1})^{4}\big{]}^{1/2}~{}\mathrm{d}x=o(1).$
Toward this end, noting that $\kappa t\mathrm{e}^{-\kappa tx}\mathrm{d}x$ is a probability measure on $[0,\infty)$ and that the square root function is concave, it follows from Jensen’s inequality that
$t^{1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}x}_{1})^{4}\big{]}^{1/2}~{}\mathrm{d}x\leq\frac{t^{1/2}}{\kappa t}\left(\int_{0}^{\infty}\kappa t\,\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}x}_{1})^{4}\big{]}~{}\mathrm{d}x\right)^{1/2}.$
If we now group together the powers of $t$ in the above, we can reduce (3.89) to
(3.90) $\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}x}_{1})^{4}\big{]}~{}\mathrm{d}x=o(1).$
If we combine (3.81) with the fact that $\Pi(t;x,x)t^{1/2}=\frac{1}{\sqrt{2\pi}}$, then we get from an exact calculation that
$\mathbf{E}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}x}_{1})^{4}\big{]}=\int_{0}^{\infty}\ell^{4}\cdot\sqrt{2\pi}(2t^{-1/2}x+\ell)\mathrm{e}^{-(2t^{-1/2}x+\ell)^{2}/2}~{}\mathrm{d}\ell\\ =8\sqrt{2\pi}e^{-2x^{2}/t}(t+2x^{2})t^{-1}-8\pi(3tx+4x^{3})\mathrm{erfc}\left(\tfrac{\sqrt{2}x}{\sqrt{t}}\right)t^{-3/2}.$
If we integrate the second line above multiplied by $\mathrm{e}^{-\kappa tx}$, an exact calculation yields
(3.91) $\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\big{[}\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}x}_{1})^{4}\big{]}~{}\mathrm{d}x\\ =\frac{192\pi\mathrm{e}^{\kappa^{2}t^{3}/8}\mathrm{erfc}\left(\frac{\kappa t^{3/2}}{2\sqrt{2}}\right)}{\kappa^{4}t^{11/2}}-\frac{192\pi}{\kappa^{4}t^{11/2}}+\frac{96\sqrt{2\pi}}{\kappa^{3}t^{4}}-\frac{24\pi}{\kappa^{2}t^{5/2}}+\frac{8\sqrt{2\pi}}{\kappa t}.$
To see why this vanishes, we combine the Taylor expansions (as $z\to 0$)
$\mathrm{e}^{z}=1+z+O(z^{2})\qquad\text{and}\qquad\mathrm{erfc}(z)=1-\tfrac{2}{\sqrt{\pi}}z+\tfrac{2}{3\sqrt{\pi}}z^{3}+O(z^{5}),$
which yields the $t\to 0^{+}$ asymptotics
$\mathrm{e}^{\kappa^{2}t^{3}/8}=1+\tfrac{\kappa^{2}t^{3}}{8}+O(t^{6})\qquad\text{and}\qquad\mathrm{erfc}\left(\frac{\kappa t^{3/2}}{2\sqrt{2}}\right)=1-\tfrac{\kappa t^{3/2}}{\sqrt{2\pi}}+\tfrac{\kappa^{3}t^{9/2}}{24\sqrt{2\pi}}+O(t^{15/2}).$
If we take a product of the above, and then multiply the result by $\frac{192\pi}{\kappa^{4}t^{11/2}}$, then we get
$\frac{192\pi\mathrm{e}^{\kappa^{2}t^{3}/8}\mathrm{erfc}\left(\frac{\kappa t^{3/2}}{2\sqrt{2}}\right)}{\kappa^{4}t^{11/2}}=\frac{192\pi}{\kappa^{4}t^{11/2}}-\frac{96\sqrt{2\pi}}{\kappa^{3}t^{4}}+\frac{24\pi}{\kappa^{2}t^{5/2}}-\frac{8\sqrt{2\pi}}{\kappa t}+O(t^{1/2}).$
If we put this into (3.91), then we obtain (3.89). This finally concludes the proof of (3.37), and thus of Proposition 3.16. Consequently, Theorem 3.1 is now proved.
Appendix A Heuristic 2.2
Recall the definitions of $Z_{\beta,n}$ and $H_{n}(\mathcal{B}^{\beta,n})$ in (2.2) and (2.3). Straightforward calculations using (2.2) reveal that
(A.1) $\mathbf{E}\big{[}H_{n}(\mathcal{B}^{\beta,n})\big{]}=-\tfrac{1}{n}\tfrac{\partial}{\partial\beta}\log Z_{\beta,n}\qquad\text{and}\qquad\mathbf{Var}\big{[}H_{n}(\mathcal{B}^{\beta,n})\big{]}=\tfrac{1}{n^{2}}\tfrac{\partial^{2}}{\partial^{2}\beta}\log Z_{\beta,n}.$
In the special case where $V(x)=x^{2}/2$, a classical calculation using a Selberg integral [55] yields the explicit formula $Z_{\beta,n}=Z_{G\beta E,n}$, where we define
$Z_{G\beta E,n}=(2\pi)^{n/2}(n\beta/2)^{-\beta n^{2}/4+(\beta/4-1/2)n}\frac{\prod_{j=1}^{n}\Gamma(1+j\beta/2)}{\Gamma(1+\beta/2)^{n}}.$
Asymptotics using several applications of the Mathematica ”Limit” function reveal that, as $n\to\infty$, one has
$-\tfrac{1}{n}\tfrac{\partial}{\partial\beta}\log Z_{G\beta E,n}=\tfrac{3}{8}n-\tfrac{1}{2}\log n-\tfrac{1}{4}\left(1+\log(\beta^{2}/4)-2\tfrac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}\right)+o(1)$
and $\tfrac{1}{n^{2}}\tfrac{\partial^{2}}{\partial^{2}\beta}\log Z_{G\beta E,n}=o(1)$. Combining this with (A.1), we get that when $V(x)=x^{2}/2$, the limit (2.4) holds. Using the well-known identity between the polygamma functions and the Hurwitz zeta function (e.g., [1, 6.4.10]), we have that for $\beta>0$,
$\frac{\mathrm{d}}{\mathrm{d}\beta}\left(\log(\beta^{2}/4)-2\frac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}\right)=\frac{2}{\beta}-\sum_{k=0}^{\infty}\frac{1}{(k+1+\beta/2)^{2}}>\frac{2}{\beta}-\int_{-1}^{\infty}\frac{1}{(k+1+\beta/2)^{2}}~{}\mathrm{d}k=0,$
where the inequality follows from the integral test since $k\mapsto(k+1+\beta/2)^{-2}$ is strictly decreasing. This implies that $\beta\mapsto\log(\beta^{2}/4)-2\tfrac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}$ is invertible for $\beta\in(0,\infty)$.

### v2

@@
     Let $\mathbb{F}_{1}=\mathbb{R}$, $\mathbb{F}_{2}=\mathbb{C}$, and $\mathbb{F}_{4}=\mathbb{H}$.
-If $r>1$ and $\beta\in\{1,2,4\}$, then let $W_{\beta}$ be a standard $r$-dimensional matrix Brownian motion on $\mathbb{F}_{\beta}$ (i.e., with GOE, GUE, or GSE increments depending on whether $\beta=1,2,4$; see [8, Page 2730]), and define
+If $r>1$ and $\beta\in\{1,2,4\}$, then let $W_{\beta}$ be a standard $r$-dimensional matrix Brownian motion on $\mathbb{F}_{\beta}$ (i.e., with GOE, GUE, or GSE increments depending on whether $\beta=1,2,4$; see [BloemendalVirag2, Page 2730]), and define
     (1.2) $\mathcal{H}_{\theta}=-\tfrac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+rx+\sqrt{2}W_{\beta}^{\prime}(x).$
@@
     (1.3) $\begin{cases}f(i,0)=0&\text{if }w_{i}=\infty\text{ (Dirichlet)},\\ f^{\prime}(i,0)=w_{i}f(i,0)&\text{if }w_{i}\in\mathbb{R}\text{ (Robin)},\end{cases}\qquad 1\leq i\leq r.$
-The SAOs were introduced by Edelman-Sutton [24], Ramírez-Rider-Virág [53], and Bloemendal-Virág [7, 8] to describe the edge fluctuations of various point processes of high interest in mathematical physics and statistics.
-That is, depending on the choice of $\theta$, the eigenvalues of $\mathcal{H}_{\theta}$ either describe the edge fluctuations of a wide class of beta-ensembles with so-called ”regular” external potential [4, 5, 11, 19, 44, 49, 53, 56], or the soft-edge fluctuations of critically-spiked Wishart or Gaussian invariant ensembles with real, complex, or quaternion entries [2, 7, 8, 45, 50, 59].
+The SAOs were introduced by Edelman-Sutton [EdelmanSutton], Ramírez-Rider-Virág [RamirezRiderVirag], and Bloemendal-Virág [BloemendalVirag1, BloemendalVirag2] to describe the edge fluctuations of various point processes of high interest in mathematical physics and statistics.
+That is, depending on the choice of $\theta$, the eigenvalues of $\mathcal{H}_{\theta}$ either describe the edge fluctuations of a wide class of beta-ensembles with so-called ”regular” external potential [BekermanFigalliGuionnet, Bekerman, BourgadeErdosYau, DeiftGioev, KrishnapurRiderVirag, PasturShcherbina, RamirezRiderVirag, Shcherbina], or the soft-edge fluctuations of critically-spiked Wishart or Gaussian invariant ensembles with real, complex, or quaternion entries [BaikBenArousPeche, BloemendalVirag1, BloemendalVirag2, Mo, Peche, Wang].
 In this paper, we are interested in the following question regarding the SAOs:
@@
 There exists a deterministic function $\mathcal{T}$ (which can be written as an explicit limit; see (3.5) for the details) such that for every $\theta\in\Theta$, one has
-(1.5) $\mathcal{T}(\Lambda^{\theta})=r_{0}+1/\beta\qquad\text{almost surely.}$.
+(1.5) $\mathcal{T}(\Lambda^{\theta})=r_{0}+1/\beta\qquad\text{almost surely.}$
 1.2. Applications
@@
 That is, the observation that an impressive amount of information about a differential operator’s domain can be inferred from its eigenvalues alone.
-One of the most well-known manifestation of this phenomenon is the Weyl law [60] and its various extensions: If $\big{\{}\lambda_{k}(M,b):k\geq 0\big{\}}$ are the eigenvalues of the Laplacian on some compact manifold $M$ subject to some boundary condition $b$ (if $\partial M\neq\varnothing$), then $M$’s dimension, $M$’s volume, $M$’s boundary area, and the boundary condition $b$ (at least Dirichlet versus Neumann) can in many cases be determined by the asymptotics of the eigenvalue counting function
+One of the most well-known manifestation of this phenomenon is the Weyl law [Weyl] and its various extensions: If $\big\{\lambda_{k}(M,b):k\geq 0\big\}$ are the eigenvalues of the Laplacian on some compact manifold $M$ subject to some boundary condition $b$ (if $\partial M\neq\varnothing$), then $M$’s dimension, $M$’s volume, $M$’s boundary area, and the boundary condition $b$ (at least Dirichlet versus Neumann) can in many cases be determined by the asymptotics of the eigenvalue counting function
 $\lambda\mapsto\big|\big\{k\geq 0:\lambda_{k}(M,b)\leq\lambda\big\}\big|.$
@@
 Thus, the proof of Theorem 1.2 relies on the observation that $r_{0}+1/\beta$ can be recovered from precise asymptotics of the $\Lambda_{k}^{\theta}$’s as $k\to\infty$.
-That being said, instead of recovering $r_{0}+1/\beta$ using the asymptotics of the eigenvalue counting function, we do so via the exponential traces $\mathrm{Tr}\big{[}\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big{]}=\sum_{k=0}^{\infty}\mathrm{e}^{-t\Lambda^{\theta}_{k}/2}$ for small $t>0$ (which is also widely used in spectral geometry; see, e.g., [38, Chapter 3] and references therein).
+That being said, instead of recovering $r_{0}+1/\beta$ using the asymptotics of the eigenvalue counting function, we do so via the exponential traces $\mathrm{Tr}\big[\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big]=\sum_{k=0}^{\infty}\mathrm{e}^{-t\Lambda^{\theta}_{k}/2}$ for small $t>0$ (which is also widely used in spectral geometry; see, e.g., [Gilkey, Chapter 3] and references therein).
 More specifically, our proof of Theorem 1.2 relies on showing that
@@
 In particular, the fact that any sequence $\mathfrak{o}_{t_{1}},\mathfrak{o}_{t_{2}},\mathfrak{o}_{t_{3}},\ldots$ such that $t_{n}\to 0^{+}$ has a bounded variance and vanishing covariance means that $r_{0}+1/\beta$ can be recovered from (1.6) by the law of large numbers, that is, by averaging the traces $\mathrm{Tr}\big[\mathrm{e}^{-t\mathcal{H}_{\theta}/2}\big]$ over a sequence of $t$’s that vanishes fast enough; see (3.5) for details.
-The main technical input in the proof of (1.6) consists of the recently-developed semigroup theory for the multivariate SAOs [29], i.e., the case $r>1$ (which is essential for our application of Theorem 1.2 to the spiked models in Corollary 2.14).
-This development extended prior results on the semigroup theory of the scalar SAOs [32, 39], i.e., the case $r=1$.
+The main technical input in the proof of (1.6) consists of the recently-developed semigroup theory for the multivariate SAOs [RigidityMSAO], i.e., the case $r>1$ (which is essential for our application of Theorem 1.2 to the spiked models in Corollary 2.14).
+This development extended prior results on the semigroup theory of the scalar SAOs [GaudreauLamarreShkolnikov, GorinShkolnikov], i.e., the case $r=1$.
 1.4. Questions of Optimality
@@
 That said, it can be shown that this cannot be achieved with the method used in this paper.
-More specifically, it is known that for some cases of $\theta$ (notably, $\theta=(1,2,\infty)$; see [31, Proposition 2.27]), the random remainder term $\mathfrak{o}_{t}$ in (1.6) satisfies
+More specifically, it is known that for some cases of $\theta$ (notably, $\theta=(1,2,\infty)$; see [RigidityFK, Proposition 2.27]), the random remainder term $\mathfrak{o}_{t}$ in (1.6) satisfies
 $\liminf_{t\to 0^{+}}\,\mathbf{Var}[\mathfrak{o}_{t}]>0.$
@@
 We adopt the convention that the points in $\mathcal{B}^{\beta,n}$ are in increasing order. The constant $Z_{\beta,n}>0$, called the partition function, ensures that (2.2) integrates to one.
-The question of whether the temperature of infinite Gibbs point processes can be recovered from a single sample has attracted some attention in the literature; see, e.g., [20, Section 5.4] and [22] and references therein.
+The question of whether the temperature of infinite Gibbs point processes can be recovered from a single sample has attracted some attention in the literature; see, e.g., [Dereudre, Section 5.4] and [DereudreLavancier] and references therein.
 In the case of a finite configuration, such as (2.2), this is of course not possible due to the mutual absolute continuity of the particles’ joint densities with different $\beta$’s.
 However, the problem becomes interesting if one asks about recovering $\beta$ as the number of particles grows to infinity.
-For instance, several prior works have studied the consistency of maximum likelihood and pseudolikelihood estimators (MLEs and MPLEs) of temperature in spin glasses in the large-dimensional limit; see, e.g., [6, 15, 17, 33] and references therein.
+For instance, several prior works have studied the consistency of maximum likelihood and pseudolikelihood estimators (MLEs and MPLEs) of temperature in spin glasses in the large-dimensional limit; see, e.g., [BhattacharyaMukherjee, Chatterjee, CometsGidas, GhosalMukherjee] and references therein.
 In view of extending this to beta-ensembles, one may then ask:
@@
 To the best of our knowledge, this has not been studied in the literature.
-That being said, one of the main hurdles in applying MLEs to Gibbs point processes consists of computing the partition function (e.g., [61, Section 1]).
-Therefore, given that $Z_{\beta,n}$ is very well understood for beta-ensembles—see [9, 10] and references therein—we expect that a similar strategy could be used to solve Question 2.1.
-To make this more concrete, consider the following heuristic, (which is informal as the asymptotics therein were calculated using Wolfram Mathematica; see Appendix A):
+That being said, one of the main hurdles in applying MLEs to Gibbs point processes consists of computing the partition function (e.g., [Younes, Section 1]).
+Therefore, given that $Z_{\beta,n}$ is very well understood for beta-ensembles—see [BorotGuionnet1, BorotGuionnet2] and references therein—we expect that a similar strategy could be used to solve Question 2.1.
+To make this more concrete, consider the following heuristic (which is informal as the asymptotics therein were calculated using Wolfram Mathematica; see Appendix A):
 Heuristic 2.2 (Informal).
@@
 As argued in Appendix A, the limit (2.4) relies on an asymptotic analysis of the partition function.
-Thus, we expect that (2.4) could be extended to more general $V$’s using the asymptotic expansions of general beta-ensemble partition functions in [9, (1-3)] and [10, (1.14)].
+Thus, we expect that (2.4) could be extended to more general $V$’s using the asymptotic expansions of general beta-ensemble partition functions in [BorotGuionnet1, (1-3)] and [BorotGuionnet2, (1.14)].
 2.1.2. New Result
 A common feature of the statistics discussed above (i.e., MLEs, MPLEs, or the total energy in (2.3)) is that they use every point in the configuration $\mathcal{B}^{\beta,n}_{1}\leq\cdots\leq\mathcal{B}^{\beta,n}_{n}$ as an input.
-In sharp contrast to this, Theorem 1.2 implies that the temperature of many beta-ensembles can be recovered from a much more parsimonious asymptotic, which only depends on the vanishing proportion of particles that contribute to edge scaling limits (as the vast majority of particles instead contribute to the bulk scaling limits; e.g., [57, 58]).
+In sharp contrast to this, Theorem 1.2 implies that the temperature of many beta-ensembles can be recovered from a much more parsimonious asymptotic, which only depends on the vanishing proportion of particles that contribute to edge scaling limits (as the vast majority of particles instead contribute to the bulk scaling limits; e.g., [ValkoVirag, ValkoVirag2]).
 This is based on the edge universality of beta-ensembles, part of which can be informally summarized as follows:
@@
 Let $V$ be a regular one-cut potential.
-That is, $V$ has an equilibrium measure $\mu_{V}$ supported on a compact interval $[\mathfrak{a},\mathfrak{b}]$, and $\mu_{V}$’s density vanishes like a square root at $\mathfrak{a}$ and $\mathfrak{b}$.
+That is, $V$ has an equilibrium measure $\mu_{V}$ supported on a compact interval $\mathfrak{a},\mathfrak{b}]$, and $\mu_{V}$’s density vanishes like a square root at $\mathfrak{a}$ and $\mathfrak{b}$.
 Under additional technical conditions on $V$, for any edge $\mathfrak{e}\in\{\mathfrak{a},\mathfrak{b}\}$, there exists scaling constants $\mathfrak{s}_{n}=O(n^{2/3})$ such that
 (2.5) $\lim_{n\to\infty}\mathfrak{s}_{n}\big(\mathfrak{e}-\mathcal{B}^{\beta,n}\big)=\Lambda^{(1,\beta,\infty)}\qquad\text{in distribution}.$
-See [4, 11, 19, 44, 49, 53, 56] for the details, as well as [5] for a multi-cut extension of Theorem 2.3, whose statement we omit for simplicity.
+See [BekermanFigalliGuionnet, BourgadeErdosYau, DeiftGioev, KrishnapurRiderVirag, PasturShcherbina, RamirezRiderVirag, Shcherbina] for the details, as well as [Bekerman] for a multi-cut extension of Theorem 2.3, whose statement we omit for simplicity.
 Given that $r_{0}=0$ when $\theta=(1,\beta,\infty)$, we thus obtain the following corollary from Theorem 1.2:
@@
 }|\mathcal{B}^{\beta,n}_{k}-\mathfrak{e}|\leq\varepsilon_{n}}\mathrm{e}^{-t_{n}\,\mathfrak{s}_{n}(\mathfrak{e}-\mathcal{B}^{\beta,n}_{k})/2}-\sum_{k=0}^{\infty}\mathrm{e}^{-t_{n}\Lambda^{(1,\beta,\infty)}_{k}/2}\right|=0$
-for $\varepsilon_{n},t_{n}=o(1)$. A similar result already exists in the special case where $V(x)=x^{2}/2$ when $t_{n}$ is replaced by a fixed $t>0$; see the proof of [39, Corollary 2.10].
+for $\varepsilon_{n},t_{n}=o(1)$.
+A similar result already exists in the special case where $V(x)=x^{2}/2$ when $t_{n}$ is replaced by a fixed $t>0$; see the proof of [GorinShkolnikov, Corollary 2.10].
 2.2. Detecting Critical Signals in Spiked Models
 2.2.1. Background
-In the early 2000’s, Johnstone introduced the spiked Wishart ensemble [42, Pages 301–304] as a means of modeling the effect of low-rank signals on the spectral statistics of covariance estimators in high dimension.
-Some years later, Péché [50] introduced a closely related additive spiked ensemble to study low-rank perturbations of self-adjoint matrices.
-In their simplest incarnation (i.e., rotationally-symmetric noise), these models are defined as follows: Let $r\in\mathbb{N}$, $\beta=\{1,2,4\}$, and $\ell=(\ell_{r-1},\ldots,\ell_{0})\in[0,\infty)^{r}$ be fixed, assuming that $\ell_{r-1}\leq\cdots\leq\ell_{0}$.
+In the early 2000’s, Johnstone introduced the spiked Wishart ensemble [Johnstone, Pages 301–304] as a means of modeling the effect of low-rank signals on the spectral statistics of covariance estimators in high dimension.
+Some years later, Péché [Peche] introduced a closely related additive spiked ensemble to study low-rank perturbations of self-adjoint matrices.
+In their simplest incarnation (i.e., rotationally-symmetric noise), these models are defined as follows: Let $r\in\mathbb{N}$, $\beta\in\{1,2,4\}$, and $\ell=(\ell_{r-1},\ldots,\ell_{0})\in[0,\infty)^{r}$ be fixed, assuming that $\ell_{r-1}\leq\cdots\leq\ell_{0}$.
   1. (1)
@@
 Let $r\in\mathbb{N}$ and $\beta=\{1,2,4\}$ be fixed. In the case of the Wishart ensemble, we assume that $p=p(n)$ depends on $n$ in such a way that
-(2.7) $p(n)=\gamma\big{(}1+o(1)\big{)}\qquad\text{as }n\to\infty$
+(2.7) $p(n)=\gamma n\big(1+o(1)\big)\qquad\text{as }n\to\infty$
 for some $\gamma\in(0,\infty)$.
@@
 Are there asymptotics of $\mathcal{S}^{\beta,\ell,n}$ (as $n\to\infty$) from which any information about the signal $\ell$ can be recovered almost surely?
-The first papers in this direction were by Baik-Ben Arous-Péché [2] and Péché [50].
+The first papers in this direction were by Baik-Ben Arous-Péché [BaikBenArousPeche] and Péché [Peche].
 Among other things, these results showed that the almost-sure asymptotic detectability of each signal $\ell_{i}$ depends on its size relative to the threshold
@@
   1. (1)
-    The supercritical case, where $\ell_{i}>\tau$ for at least one $10\leq i\leq r-$ 1;
+    The supercritical case, where $\ell_{i}>\tau$ for at least one $0\leq i\leq r-1$;
   2. (2)
@@
 To the best of our knowledge, the current state of the art for this theory can be briefly summarized as follows. Firstly, we state a result for the supercritical case:
-Theorem 2.7 ([2, 3, 14, 50]).
+Theorem 2.7 ([BaikBenArousPeche, BaikSilverstein, CapitaineDonatiMartinFeral, Peche]).
 Let $\mathfrak{e}$ be the top edge of the limiting empirical spectral distribution (ESD) of the unperturbed Wishart or Gaussian invariant model, i.e.,
@@
 For any fixed $\ell\in[0,\infty)^{r}$ and $i\geq 0$, one has
-(2.9) $\lim_{n\to\infty}\mathcal{S}^{\ell,\beta,n}_{n-i}=\begin{cases}\mathcal{O}(\ell_{i})&\text{if $i\leq r-1$ and $\ell_{i}>\tau$},\\ \mathfrak{e}&\text{if $i\leq r-1$ and $\ell_{i}\leq\tau$, or if $i>r-1$}\end{cases}\qquad\text{almost surely}.$
+[…]
 Remark 2.8.
@@
 Secondly, we state a result regarding the impossibility of detecting subcritical signals almost-surely using the eigenvalues $\mathcal{S}^{\beta,\ell,n}$ alone:
-Theorem 2.9 ([41, 46, 47, 48]).
+Theorem 2.9 ([JungChungLee, MontanariReichmanZeitouni, OnatskiMoreiraHallin, OnatskiMoreiraHallin2]).
 Let $\beta=1$ (i.e., $\mathbb{F}_{\beta}=\mathbb{R}$).
@@
 Despite the impossibility result in Theorem 2.9, one could ask if there nevertheless exist tests that distinguish between the hypotheses
-(2.10) $\boldsymbol{H}_{0}\text{ : "no perturbation"}\qquad\text{versus}\qquad\boldsymbol{H}_{\ell}\text{ : "perturbation $=\ell$"}$ ”
+[…]
 with nontrivial statistical power when $\ell\neq 0$ is subcritical.
-The short answer to this question is yes; see, e.g., [25, 41, 43, 47, 48] and references therein for a non-exhaustive sample of this very sizeable literature.
+The short answer to this question is yes; see, e.g., [ElAlaouiKrzakalaJordan, JungChungLee, JohnstoneOnatski, OnatskiMoreiraHallin, OnatskiMoreiraHallin2] and references therein for a non-exhaustive sample of this very sizeable literature.
 A unifying feature of all these results is that the statistical tests that achieve optimal or close-to-optimal power in distinguishing (2.10) (e.g., likelihood ratios or linear spectral statistics) all use the full spectrum $\mathcal{S}^{\beta,\ell,n}_{1}\leq\cdots\leq\mathcal{S}^{\beta,\ell,n}_{n}$ as an input.
-In the words of Dobriban [23]:
+In the words of Dobriban [Dobriban]:
 All eigenvalues matter to achieve sharp detection of weak principal components in high-dimensional data.
 To give a simple example of how such results look like, consider the following:
-Theorem 2.11 ([25, 41]).
+Theorem 2.11 ([ElAlaouiKrzakalaJordan, JungChungLee]).
 Suppose that $\beta=1$, and that we consider the Gaussian invariant model $Y_{\beta,\ell,n}$ (hence $\tau=1$).
@@
 The LRT minimizes (2.11) over all measurable functions of $\mathcal{S}^{\beta,\ell,n}$. Moreover,
-(2.12) $\lim_{n\to\infty}\mathcal{E}_{n}(\ell)=\mathrm{erfc}\left(\tfrac{r}{4}\sqrt{-\log(1-\lambda)}\right)\in(0,1),$
+(2.12) $\lim_{n\to\infty}\mathcal{E}_{n}(\lambda)=\mathrm{erfc}\left(\tfrac{r}{4}\sqrt{-\log(1-\lambda)}\right)\in(0,1),$
 where we use $\mathrm{erfc}(z)=\frac{2}{\sqrt{\pi}}\int_{z}^{\infty}\mathrm{e}^{-t^{2}}~\mathrm{d}t$ to denote the complementary error function.
 Thirdly, we did not find any works in the literature that directly address Question 2.6 in the critical case $\ell_{r_{\mathrm{c}}-1},\ldots,\ell_{0}=\tau+o(1)$.
-However, as pointed out in [51, Appendix A], consistent estimators of $\ell$ in this case can be proved to exist by examining the asymptotic power of subcritical tests when $\ell_{i}\to\tau^{-}$ (e.g., (2.12) goes to zero when $\ell_{i}=\lambda\to 1^{-}$).
+However, as pointed out in [PerryWeinBandeiraMoitra, Appendix A], consistent estimators of $\ell$ in this case can be proved to exist by examining the asymptotic power of subcritical tests when $\ell_{i}\to\tau^{-}$ (e.g., (2.12) goes to zero when $\ell_{i}=\lambda\to 1^{-}$).
 Thus, the answer to Question 2.6 is positive in the critical case.
@@
 This is based on the following:
-Theorem 2.13 ([7, 8]; see also [2, 45, 50, 59]).
+Theorem 2.13 ([BloemendalVirag1, BloemendalVirag2]; see also [BaikBenArousPeche, Mo, Peche, Wang]).
 Let $r\in\mathbb{N}$ and $\beta=\{1,2,4\}$ be fixed, let $\gamma$ and $p=p(n)$ be as in (2.7) in the Wishart case, and suppose that $\ell=\ell(n)\in[0,\infty)^{r}$ is either fixed or depends on $n$.
@@
 (2.13) $w_{i}=\begin{cases}\displaystyle\tfrac{1}{\sqrt{\gamma}}\left(1+\tfrac{1}{\sqrt{\gamma}}\right)^{-2/3}\lim_{n\to\infty}n^{1/3}\big(\sqrt{\gamma}-\ell_{i-1}\big)&\text{(Wishart)},\\ \displaystyle\lim_{n\to\infty}n^{1/3}(1-\ell_{i-1})&\text{(Gaussian invariant)},\end{cases}$
-exists and is an element of (,]($\infty$,$\infty$]( - ∞ , ∞ ]; then let $w=(w_{1},\ldots,w_{r})$. If we define
+exists and is an element of $(-\infty,\infty]$; then let $w=(w_{1},\ldots,w_{r})$. If we define
 $\mathfrak{s}_{n}=\begin{cases}\tfrac{1}{\sqrt{\gamma}}\left(1+\tfrac{1}{\sqrt{\gamma}}\right)^{-4/3}n^{2/3}&\text{(Wishart)},\\ n^{2/3}&\text{(Gaussian invariant)},\end{cases}$
@@
 for fixed $t>0$ has not been proved for any spiked model when $r_{0}>0$, even in the simplest case of the scalar SAO (i.e., $r=r_{0}=1$).
-Indeed, using current techniques, such a result would first require proving a seemingly difficult strong invariance principle for reflected random walks; see [27, Conjectures 2.23 and 6.11].
+Indeed, using current techniques, such a result would first require proving a seemingly difficult strong invariance principle for reflected random walks; see [ConvergenceToSpiked, Conjectures 2.23 and 6.11].
 2.3. Rigidity
 2.3.1. Background
-In the last 15 years, a sizeable literature on ”rigidity” properties of point processes has been developing, starting with pioneering works by Ghosh and Peres [34, 37].
-We refer to [18, 35] and references therein for a survey.
+In the last 15 years, a sizeable literature on ”rigidity” properties of point processes has been developing, starting with pioneering works by Ghosh and Peres [Ghosh, GhoshPeres].
+We refer to [Coste, GhoshLebowitz] and references therein for a survey.
 In this literature, rigidity refers to the observation that for certain strongly-correlated point processes, some features of the point configuration inside a bounded set is almost-surely determined by the configuration of points outside that set.
@@
 Example 2.17.
-The Ginibre point process was shown to be number rigid in [37].
+The Ginibre point process was shown to be number rigid in [GhoshPeres].
 Example 2.18.
 The $\alpha$-Gaussian analytic function ($\alpha$-GAF) zero point process (i.e., the zeros of the random entire function $z\mapsto\sum_{k\geq 0}\frac{g_{k}}{(k!)^{\alpha/2}}z^{k}$, where $g_{k}$ are i.i.d.
-complex Gaussians) were shown to be $k$-moment rigid for all $0\leq k\leq\lfloor 1/\alpha\rfloor$ in [36, 37].
+complex Gaussians) were shown to be $k$-moment rigid for all $0\leq k\leq\lfloor 1/\alpha\rfloor$ in [GhoshKrishnapur, GhoshPeres].
 Remark 2.19.
-Rigidity was in part introduced to help distinguish structured point processes that are otherwise difficult to tell apart. As stated in [37, Page 1794]:
+Rigidity was in part introduced to help distinguish structured point processes that are otherwise difficult to tell apart. As stated in [GhoshPeres, Page 1794]:
 While a simple visual inspection suffices to (heuristically) distinguish a sample of the Poisson process from that of either the Ginibre or the GAF zero process (of the same intensity), the latter two are hard to set apart between themselves.
@@
   1. (1)
-    For the Ginibre process, it was shown in [37] that the number of points is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
+    For the Ginibre process, it was shown in [GhoshPeres] that the number of points is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
   2. (2)
-    For the $\alpha$-GAF zeros, it was shown in [26, 37] that the $k^{\mathrm{th}}$ moments for $0\leq k\leq\lfloor 1/\alpha\rfloor$ is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
-A formal version of these optimality results can be stated using the notion of tolerance; see, e.g., [36, Definitions 2 and 3].
+    For the $\alpha$-GAF zeros, it was shown in [GangopadhyayGhoshTan, GhoshPeres] that the $k^{\mathrm{th}}$ moments for $0\leq k\leq\lfloor 1/\alpha\rfloor$ is the only information about the configuration inside a bounded set $B$ that can be obtained from the configuration outside $B$.
+A formal version of these optimality results can be stated using the notion of tolerance; see, e.g., [GhoshKrishnapur, Definitions 2 and 3].
 Going back to our main topic, the number rigidity of $\Lambda^{\theta}$ was proved in a sequence of three papers.
-Namely, first for $\theta=(1,2,\infty)$ in [12], then for $\theta=(1,\beta,w)$ with any choice of $\beta>0$ and $w\in(-\infty,\infty]$ in [30], and finally for all $\theta\in\Theta$ in [29].
+Namely, first for $\theta=(1,2,\infty)$ in [Bufetov16], then for $\theta=(1,\beta,w)$ with any choice of $\beta>0$ and $w\in(-\infty,\infty]$ in [RigiditySAO], and finally for all $\theta\in\Theta$ in [RigidityMSAO].
 2.3.2. New Results
-Theorem 1.2 refines current results on $\Lambda^{\theta}$’s rigidity in two directions: On the one hand, while the results of [29, 30] proved the existence of a single function $\mathcal{M}$ such that
-(2.16) $\Lambda^{\theta}\cap B=\mathcal{M}(\Lambda^{\theta}\cap B^{c})\qquad\text{almost surely}$
+Theorem 1.2 refines current results on $\Lambda^{\theta}$’s rigidity in two directions: On the one hand, while the results of [RigidityMSAO, RigiditySAO] proved the existence of a single function $\mathcal{M}$ such that
+(2.16) $|\Lambda^{\theta}\cap B|=\mathcal{M}(\Lambda^{\theta}\cap B^{c})\qquad\text{almost surely}$
 for every bounded Borel set $B$, they did not provide an explicit construction of $\mathcal{M}$.
@@
 To the best of our knowledge, the optimality of number rigidity for $\Lambda^{\theta}$ (i.e., tolerance) has not been proved.
-That said, we expect that such a result could be proved using the technology developed in [26].
-(The mutual absolute continuity of Palm measures of the Airy process—i.e., $\Lambda^{(1,2,\infty)}$—with the same number of fixed components proved in [13] serves as further evidence of this point.)
+That said, we expect that such a result could be proved using the technology developed in [GangopadhyayGhoshTan].
+(The mutual absolute continuity of Palm measures of the Airy process—i.e., $\Lambda^{(1,2,\infty)}$—with the same number of fixed components proved in [Bufetov18] serves as further evidence of this point.)
 In this context, our main result can be viewed as a first step toward developing statistics that distinguish between different instances of SAO spectra, which goes beyond the rigidity/tolerance hierarchy:
 Corollary 2.22.
-If $\theta=(r,\beta,w)$ and $\tilde{\theta}=(\tilde{r},\tilde{\beta},\tilde{w})$ are such that $r_{0}+1/\beta\neq\tilde{r}_{0}+1/\tilde{\beta}$, then $\Lambda^{\theta}$ and $\Lambda^{\tilde{\theta}}$ are mutually singular.
+[…]
 2.3.3. Open Problems
@@
 Open Problem 2.24.
-Find a function $\Psi$ on the parameter space $\Theta$ such that $\Lambda^{\theta}$ and $\Lambda^{\tilde{\theta}}$ are mutually singular if and only if $\Psi(\theta)\neq\Psi(\tilde{\theta})$.
-Once $\Psi$ is found, characterize the possible relationships between the laws of $\Lambda^{\theta}$ and $\Lambda^{\tilde{\theta}}$ when $\Psi(\theta)=\Psi(\tilde{\theta})$ (e.g., equality, mutual absolute continuity, etc.).
+[…]
 2.4. Cancellation of Subcritical Coordinates
 2.4.1. Background
-In [8, Page 2734], Bloemendal and Virág made the curious observation that any given spiked Wishart or Gaussian invariant model converges to an infinite number of SAOs with different parameters: Looking back at Theorem 2.13, if we add a trivial component to the spike $\ell=(\ell_{r-1},\ldots,\ell)$, thus transforming it into ~$\tilde{\ell}=(0,\ell_{r-1},\ldots,\ell_{0})$, then the matrix model remains unchanged.
+In [BloemendalVirag2, Page 2734], Bloemendal and Virág made the curious observation that any given spiked Wishart or Gaussian invariant model converges to an infinite number of SAOs with different parameters: Looking back at Theorem 2.13, if we add a trivial component to the spike $\ell=(\ell_{r-1},\ldots,\ell_{0})$, thus transforming it into $\tilde{\ell}=(0,\ell_{r-1},\ldots,\ell_{0})$, then the matrix model remains unchanged.
 However, the parameter $r$ in the limit (2.14) increases by one, and the added component has a Dirichlet boundary condition (since $\ell_{r}=0$ corresponds to $w_{r+1}=\infty$ in (2.13)).
 If we combine this observation with the fact that the SAOs are obviously invariant with respect to permutations of their components (as their noises are induced by GOE, GUE, and GSE matrices), then we obtain the following from Theorem 2.13:
-Corollary 2.25 ([8]).
+Corollary 2.25 ([BloemendalVirag2]).
 Let $\beta\in\{1,2,4\}$, and let $\theta=(r,\beta,w),\tilde{\theta}=(\tilde{r},\beta,\tilde{w})\in\Theta$ be such that
@@
 where $\mathfrak{o}_{t}$ satisfy (1.7) and (1.8). With this in hand, we can obtain necessary conditions on the parameter $\eta$ in order for the cancellations in Corollary 2.25 to occur.
-More specifically, let $\tilde{\theta}=(\tilde{r},\beta,\tilde{w})$ be related to $\theta$ in such a way that (2.18) and (2.19) hold.
-Then, consider $\eta=(\kappa,\sigma,\upsilon)$ such that, similarly to the SAOs,
+[…]
 (2.21) $\kappa=\kappa(r)\text{ possibly depends on }r,$
@@
 (3.2) $\mathbf{Cov}\left[\mathrm{Tr}\big[\mathrm{e}^{-s\widehat{H}_{\theta,\eta}}\big],\mathrm{Tr}\big[\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big]\right]\leq C\left(\frac{\min\{s,t\}}{\max\{s,t\}}\right)^{1/4}\qquad\text{ for every }s,t\in(0,1].$
-The covariance bound in Theorem 3.1 was proved in [29, (8.4)].
+The covariance bound in Theorem 3.1 was proved in [RigidityMSAO, (8.4)].
 Consequently, we only need to prove (3.1).
 The remainder of this section is organized as follows: Before we prove (3.1), in Section 3.1 we show how Theorem 3.1 is used to construct the function $\mathcal{T}$ in (1.5), and thus prove Theorem 1.2.
-Then, in Section 3.2, we state a Feynman-Kac formula in proved in [29] for the expected trace $\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]$.
+Then, in Section 3.2, we state a Feynman-Kac formula proved in [RigidityMSAO] for the expected trace $\mathbf{E}\left[\mathrm{Tr}\big[\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big]\right]$.
 Finally, in Sections 3.3 and 3.4, we use the Feynman-Kac formula in question to prove the asymptotics in (3.1).
@@
 Definition 3.2.
-Let $c_{1},c_{2}>0$ be any fixed constants. Define the vanishing sequence of positive numbers $01\geq t_{1}>t_{2}>\cdots>$ 0 as
+Let $c_{1},c_{2}>0$ be any fixed constants. Define the vanishing sequence of positive numbers $1\geq t_{1}>t_{2}>\cdots>0$ as
 (3.3) $t_{n}=\mathrm{e}^{-c_{1}n},\qquad n\geq 1,$
@@
 By a trivial rearrangement of the terms in (3.5), we therefore obtain that $\mathcal{T}(\Lambda^{\theta})=r_{0}+1/\beta$ almost surely if we can prove that
-$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}-\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}\right]\right)=0\qquad\text{almost surely.}$.
+$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\mathrm{Tr}\big[\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big]-\mathbf{E}\left[\mathrm{Tr}\big[\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big]\right]\right)=0\qquad\text{almost surely.}$
 Toward this end, by Markov’s inequality, we get that for every $m\in\mathbb{N}$ and $\varepsilon>0$,
@@
 where $t_{n}$ and $N_{m}$ are defined as in (3.3) and (3.4) respectively. Moreover, for any $k\geq 0$, one has
-$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\mathrm{e}^{-t_{n}\Lambda^{\theta}_{k}/2}=1\qquad\text{almost surely.}$.
+$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\mathrm{e}^{-t_{n}\Lambda^{\theta}_{k}/2}=1\qquad\text{almost surely.}$
 Consequently, if we define the map
@@
 3.2. Proof of (3.1) Step 1. Expected Trace Formula
-The first step in our proof of (3.1) is to state the Feynman-Kac formula for $\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]$ obtained in [29, Theorem 4.9], which is what we will use to calculate the leading orders of the expected trace.
+The first step in our proof of (3.1) is to state the Feynman-Kac formula for $\mathbf{E}\left[\mathrm{Tr}\big[\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big]\right]$ obtained in [RigidityMSAO, Theorem 4.9], which is what we will use to calculate the leading orders of the expected trace.
 This is stated as Lemma 3.13 below.
@@
 $\Pi_{B}(t;x,y)=\frac{\mathrm{e}^{-(x-y)^{2}/2t}}{\sqrt{2\pi t}},\qquad t>0,~x,y\in\mathbb{R}.$
-We denote $B^{x}=(B|B(0)=x)$ and $B^{x,y}_{t}=(B|B(0)=x\text{ and }B(t)=y)$. Then, let $X=|B|$ be a reflected standard Brownian motion on $[0,\infty)$, and denote $X$’s transition kernel by
+We denote $B^{x}=(B|B(0)=x)$ and $B^{x,y}_{t}=(B|B(0)=x\text{ and }B(t)=y)$. Then, let $X=|B|$ be a reflected standard Brownian motion on $0,\infty)$, and denote $X$’s transition kernel by
 $\Pi_{X}(t;x,y)=\frac{\mathrm{e}^{-(x-y)^{2}/2t}}{\sqrt{2\pi t}}+\frac{\mathrm{e}^{-(x+y)^{2}/2t}}{\sqrt{2\pi t}},\qquad t>0,~x,y\geq 0.$
@@
 Finally, we use the shorthands $\mathfrak{L}_{t}^{0}(Z)=\mathfrak{L}_{[0,t)}^{0}(Z)$ and $L_{t}^{y}(Z)=L_{[0,t)}^{y}(Z)$.
-If we were only dealing with scalar-valued Schrödinger operators on $[0,\infty)$, then these would be the only processes that are needed to state the Feynman-Kac formula.
+If we were only dealing with scalar-valued Schrödinger operators on $0,\infty)$, then these would be the only processes that are needed to state the Feynman-Kac formula.
 However, since we are dealing with the vector-valued setting, we need more tools.
@@
 $\langle f,g\rangle=\int_{0}^{\infty}f(x)g(x)~\mathrm{d}x\qquad\text{and}\qquad\|f\|_{2}=\sqrt{\langle f,f\rangle}.$
-We denote $\mathcal{A}=\{1,\ldots,r\}\times[0,\infty)$, and we let $\mu$ denote the measure on $\mathcal{A}$ obtained by taking the product of the counting measure on $\{1,\ldots,r\}$ and the Lebesgue measure on $[0,\infty)$.
+We denote $\mathcal{A}=\{1,\ldots,r\}\times[0,\infty)$, and we let $\mu$ denote the measure on $\mathcal{A}$ obtained by taking the product of the counting measure on $\{1,\ldots,r\}$ and the Lebesgue measure on $0,\infty)$.
 Given $f:\mathcal{A}\to\mathbb{F}_{\beta}$, we denote
@@
 Let $t>0$ and $x\geq 0$, let $n\geq 2$ be an even integer, and let $q\in\mathcal{P}_{n}$.
-Given a realization of $X^{x,x}_{t}$, we let $\mathrm{si}_{t,n,q,X^{x,x}_{t}}$ denote the unique random Borel probability measure on $[0,t)^{n}$ such that for every $s_{1},t_{1},\ldots,s_{n},t_{n}\in[0,t)$ with $s_{k}<t_{k}$ for $k=1,\ldots,n$, one has
+Given a realization of $X^{x,x}_{t}$, we let $\mathrm{si}_{t,n,q,X^{x,x}_{t}}$ denote the unique random Borel probability measure on $0,t)^{n}$ such that for every $s_{1},t_{1},\ldots,s_{n},t_{n}\in[0,t)$ with $s_{k}<t_{k}$ for $k=1,\ldots,n$, one has
 (3.8) $\mathrm{si}_{t,n,q,X^{x,x}_{t}}\big([s_{1},t_{1})\times[s_{2},t_{2})\times\cdots\times[s_{n},t_{n})\big)\\ =\frac{1}{\|L_{t}(X^{x,x}_{t})\|_{2}^{2n}}\prod_{\{\ell_{1},\ell_{2}\}\in q}\left\langle L_{[s_{\ell_{1}},t_{\ell_{1}})}(X^{x,x}_{t}),L_{[s_{\ell_{2}},t_{\ell_{2}})}(X^{x,x}_{t})\right\rangle$
@@
 respectively denote $X^{x,x}_{t}$’s boundary local time at zero and local time collected on the intervals of $[0,t]$ where the jump process $\widehat{U}^{i}_{t}$ is equal to $j$.
-We note that since $[u,v)\mapsto\mathfrak{L}^{0}_{[u,v)}(X^{x,x}_{t})$ and $[u,v)\mapsto L^{y}_{[u,v)}(X^{x,x}_{t})$ induce measures on $[0,\infty)$, it is clear that
+We note that since $u,v)\mapsto\mathfrak{L}^{0}_{[u,v)}(X^{x,x}_{t})$ and $u,v)\mapsto L^{y}_{[u,v)}(X^{x,x}_{t})$ induce measures on $0,\infty)$, it is clear that
 (3.14) $\sum_{j=1}^{r}\mathfrak{L}^{(j,0)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=\mathfrak{L}_{t}(X^{x,x}_{t})\quad\text{and}\quad\sum_{j=1}^{r}L^{(j,y)}_{t}(\widehat{U}^{i}_{t},X^{x,x}_{t})=L^{y}_{t}(X^{x,x}_{t}).$
@@
 $J_{k}=\big(M^{i}_{k-1},M^{i}_{k}\big),\qquad k\geq 1.$
-Then, we let $J_{k}^{*}=\big{(}M^{i}_{k},M^{i}_{k-1}\big{)}$ denote these same jumps in reverse order.
-We define the combinatorial constant $\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})$ as follows:
+Then, we let $J_{k}^{*}=\big(M^{i}_{k},M^{i}_{k-1}\big)$ denote these same jumps in reverse order. We define the combinatorial constant $\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})$ as follows:
   1.   (a)
@@
     (3.15) $\mathfrak{D}_{t}(\widehat{p},\widehat{U}^{i}_{t})=2^{-\widehat{N}(t)/2}\sum_{m\in\mathcal{B}^{0,0}_{N(t)},~m\text{ respects }(\widehat{p},J)}(-1)^{f(m,\widehat{p},J)}.$
-We are now finally in a position to state our Feynman-Kac formula for the trace expectation, which is a special case of [29, Theorem 4.9]:
+We are now finally in a position to state our Feynman-Kac formula for the trace expectation, which is a special case of [RigidityMSAO, Theorem 4.9]:
 Lemma 3.13.
@@
 Remark 3.14.
-Theorem 4.9 in [29] provides a formula for any joint moment of the form $\mathbf{E}\big{[}\mathrm{Tr}\big{[}\mathrm{e}^{-t_{1}\widehat{H}}\big{]}\cdots\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\widehat{H}}\big{]}\big{]}$ for general vector-valued Schrödinger operators $\widehat{H}=-\frac{1}{2}\Delta+V+\xi$ on any one-dimensional domain $I\subset\mathbb{R}$.
-The statement in Lemma 3.13 can be recovered from the latter by taking $n=1$ and $t=t_{1}$, then considering what is called ”Case 2” in [29] (i.e., an operator on the half-line), and then taking the potential function $V(i,x)=\kappa x$.
+Theorem 4.9 in [RigidityMSAO] provides a formula for any joint moment of the form $\mathbf{E}\big[\mathrm{Tr}\big[\mathrm{e}^{-t_{1}\widehat{H}}\big]\cdots\mathrm{Tr}\big[\mathrm{e}^{-t_{n}\widehat{H}}\big$$\big]$ for general vector-valued Schrödinger operators $\widehat{H}=-\frac{1}{2}\Delta+V+\xi$ on any one-dimensional domain $I\subset\mathbb{R}$.
+The statement in Lemma 3.13 can be recovered from the latter by taking $n=1$ and $t=t_{1}$, then considering what is called ”Case 2” in [RigidityMSAO] (i.e., an operator on the half-line), and then taking the potential function $V(i,x)=\kappa x$.
 Remark 3.15.
@@
 3.3.3. Four or More Jumps
-We now discuss $T_{\geq 4}(t)$. For this purpose, we begin with the observation that the combinatorial constant introduced in Definition 3.12 has a very simple upper bound (see [29, (4.2)]):
+We now discuss $T_{\geq 4}(t)$. For this purpose, we begin with the observation that the combinatorial constant introduced in Definition 3.12 has a very simple upper bound (see [RigidityMSAO, (4.2)]):
 $|\mathfrak{C}_{t}(\widehat{p},\widehat{U}^{i}_{t})|\leq 1.$
@@
 If we replace $F(X^{x,x}_{t})$ by $F(X^{x,x}_{t})\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}$, then (3.58) can be further simplified.
-Indeed, given that $B^{x,-x}_{t}$ must obviously reach zero as it travels from $x$ to $-x$, one has $\mathbf{1}_{\{\mathfrak{L}^{0}(|B^{x,-x}_{t}|)=0\}}=0$.
+Indeed, given that $B^{x,-x}_{t}$ must obviously reach zero as it travels from $x$ to x-x, one has $\mathbf{1}_{\{\mathfrak{L}^{0}(|B^{x,-x}_{t}|)=0\}}=0$.
 Thus, in that case (3.58) simplifies to
@@
   1. (1)
-    By the strong Markov property and the symmetry of Brownian motion about 00, we notice that
+    By the strong Markov property and the symmetry of Brownian motion about 0, we notice that
     $\big(|B^{x,x}_{t}|~\big|~\exists s\leq t:~B^{x,x}_{t}(s)=0\big)\stackrel{{\scriptstyle\mathrm{distr.}}}{{=}}|B^{x,-x}_{t}|$
-    (i.e., we can couple the two processes above by reflecting the path of $B^{x,x}_{t}$ after its first passage to zero, thus making it end at $-x$ instead of $x$).
+    (i.e., we can couple the two processes above by reflecting the path of $B^{x,x}_{t}$ after its first passage to zero, thus making it end at x-x instead of $x$).
   2. (2)
-    By using the joint density of the running minimum and the current value of a Brownian motion (e.g., [54, Chapter III, Exercise 3.14]), we obtain that
+    By using the joint density of the running minimum and the current value of a Brownian motion (e.g., [RevuzYor, Chapter III, Exercise 3.14]), we obtain that
     $\mathbf{P}\big[\exists s\leq t:~B^{x,x}_{t}(s)=0\big]=\mathrm{e}^{-2x^{2}/t}.$
@@
 Finally, while (3.58) and (3.61) are useful for exact calculations (and thus necessary to obtain the explicit leading order terms in our asymptotics), simpler expressions can be used when we are only interested in upper estimates.
-For this purpose, we invoke [28, (5.9)], which states that for any nonnegative functional $F$, one has
+For this purpose, we invoke [GaudreauLamarreEJP, (5.9)], which states that for any nonnegative functional $F$, one has
 (3.62) $\mathbf{E}\big[F(X^{x,x}_{t})\big]\leq 2\mathbf{E}\big[F(|B^{x,x}_{t}|)\big].$
@@
 For every fixed $x>0$, there exists some finite random $t_{0}$ small enough so that $B^{0,0}_{1}$’s path does not intersect $(-\infty,-t^{-3/2}x]$ for every $t<t_{0}$. Therefore,
-$\lim_{t\to 0^{+}}\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}=\lim_{t\to 0^{+}}\int_{0}^{\infty}L_{1}^{y-t^{-3/2}x}(B^{0,0}_{1})^{2}~{}\mathrm{d}y=\|L_{1}(B^{0,0}_{1})\|_{2}^{2}\qquad\text{almost surely.}$.
+$\lim_{t\to 0^{+}}\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}=\lim_{t\to 0^{+}}\int_{0}^{\infty}L_{1}^{y-t^{-3/2}x}(B^{0,0}_{1})^{2}~\mathrm{d}y=\|L_{1}(B^{0,0}_{1})\|_{2}^{2}\qquad\text{almost surely.}$
 Consequently, (3.66) follows from dominated convergence if we prove that
@@
 $\mathbf{E}\big[\|L_{t}(B^{0,0}_{1})\|_{2}^{2}\big]=\int_{\mathbb{R}}\mathbf{E}\big[L^{y}_{1}(B^{0,0}_{1})^{2}\big]~\mathrm{d}x.$
-By [52, (1)], we have the joint density
+By [Pitman, (1)], we have the joint density
 (3.70) $\mathbf{P}\left[L_{1}^{y}(B^{0})\in\mathrm{d}\ell,B^{0}(1)\in\mathrm{d}x\right]=\frac{(|y|+|x-y|+\ell)\mathrm{e}^{-(|y|+|x-y|+\ell)^{2}/2}}{\sqrt{2\pi}}$
@@
 (3.77) $\int_{0}^{\infty}\left|\mathbf{E}\left[\int_{0}^{1}|x+B^{0,0}_{1}(s)|-x~\mathrm{d}s\right]\right|~\mathrm{d}x=\int_{0}^{\infty}\left|\int_{0}^{1}\mathbf{E}\big[|x+B^{0,0}_{1}(s)|-x\big]~\mathrm{d}s\right|~\mathrm{d}x<\infty,$
-then we get (3.75) (actually, the stronger statement RHS of (3.75) $=O(t^{3/2})$) by dominated convergence since $\mathrm{e}^{-\kappa t^{3/2}x}\uparrow 1$ as $t\to 0^{+}$.
-Toward this end, since $B^{0,0}_{1}(s)$ is Gaussian with mean zero and variance $s(1-s)$, a direct calculation yields
+then we get (3.75) (actually, the stronger statement RHS of (3.75) $O(t^{3/2})$) by dominated convergence since $\mathrm{e}^{-\kappa t^{3/2}x}\uparrow 1$ as $t\to 0^{+}$.
+Toward this end, since $B^{0,0}_{1}(s)$ is Gaussian with mean zero and variance s(1s)s(1-s), a direct calculation yields
 $\mathbf{E}\big[|x+B^{0,0}_{1}(s)|-x\big]=\sqrt{\frac{2s(1-s)}{\pi}}\mathrm{e}^{-x^{2}/2s(1-s)}-x\,\mathrm{erfc}\left(\frac{x}{\sqrt{2s(1-s)}}\right).$
@@
 $\mathbf{E}\big[\mathfrak{L}^{0}_{t}(B^{x,\pm x}_{t})\big]=t^{1/2}\mathbf{E}\big[\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,\pm t^{-1/2}x}_{1})\big]$
-Thanks to [52, (1)], for every $x,y\in\mathbb{R}$ and $\ell>0$, we have that
+Thanks to [Pitman, (1)], for every $x,y\in\mathbb{R}$ and $\ell>0$, we have that
 $\mathbf{P}[\mathfrak{L}^{0}_{1}(B^{x})\in\mathrm{d}\ell,B^{x}(1)\in\mathrm{d}y]=\frac{(|x|+|y|+\ell)}{\sqrt{2\pi}}\mathrm{e}^{-(|x|+|y|+\ell)^{2}/2}.$
@@
 (3.85) $\leq\mathbf{E}\left[\mathrm{e}^{6\mathfrak{c}\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|}\right]^{1/6}\sup_{z>0}\mathbf{E}\left[\mathrm{e}^{6\mathfrak{c}\|L_{1}(|B^{z,z}_{1}|)\|_{2}^{2}}\right]^{1/6}\sup_{z>0}\mathbf{E}\left[\mathrm{e}^{6\mathfrak{c}\mathfrak{L}^{0}_{1}(B^{z,z}_{1})}\right]^{1/6}.$
-Since $|B^{0,0}_{1}|$ is a Bessel bridge of dimension 1, the first expectation in the second line of (3.85) is finite thanks to the Gaussian tail bound for Bessel bridge maxima in [40, Remark 3.1].
-The third expectation in the second line of (3.85) is finite thanks to [28, Lemma 5.8].
+Since $|B^{0,0}_{1}|$ is a Bessel bridge of dimension 1, the first expectation in the second line of (3.85) is finite thanks to the Gaussian tail bound for Bessel bridge maxima in [GruetShi, Remark 3.1].
+The third expectation in the second line of (3.85) is finite thanks to [GaudreauLamarreEJP, Lemma 5.8].
 Finally, if we use the bound $(z_{1}+z_{2})^{2}\leq 2(z_{1}^{2}+z_{2}^{2})$, then we see that
@@
 (3.86) $\sup_{z>0}\mathbf{E}\left[\mathrm{e}^{6\mathfrak{c}\|L_{1}(|B^{z,z}_{1}|)\|_{2}^{2}}\right]^{1/6}\leq\sup_{z>0}\mathbf{E}\left[\mathrm{e}^{12\mathfrak{c}\|L_{1}(B^{z,z}_{1})\|_{2}^{2}}\right]^{1/6}<\infty,$
-where the last inequality is from [28, Lemma 5.11]. In summary, (3.85) is finite, which means that we can now reduce (3.84) to
+where the last inequality is from [GaudreauLamarreEJP, Lemma 5.11]. In summary, (3.85) is finite, which means that we can now reduce (3.84) to
 (3.87) $t^{-1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\Bigg[\Big(t^{3/2}\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|+t^{3/2}\|L_{1}(|B^{x_{t},x_{t}}_{1}|)\|_{2}^{2}+t^{1/2}\mathfrak{L}^{0}_{1}(B^{x_{t},x_{t}}_{1})\Big)^{4}\Bigg]^{1/2}~\mathrm{d}x=o(1).$
@@
 $\mathbf{E}\Big[\sup_{0\leq s\leq 1}|B^{0,0}_{1}(s)|^{4}\Big]<\infty$
-thanks to the fact that $|B^{0,0}_{1}|$ supremum has Gaussian tails ([40, Remark 3.1]). Thirdly, it follows from (3.86) that
+thanks to the fact that $|B^{0,0}_{1}|$ supremum has Gaussian tails ([GruetShi, Remark 3.1]). Thirdly, it follows from (3.86) that
 $\sup_{z>0}\mathbf{E}\big[\|L_{1}(|B^{z,z}_{1}|)\|_{2}^{8}\big]<\infty.$
@@
 (3.89) $t^{1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\big[\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}x}_{1})^{4}\big]^{1/2}~\mathrm{d}x=o(1).$
-Toward this end, noting that $\kappa t\mathrm{e}^{-\kappa tx}\mathrm{d}x$ is a probability measure on $[0,\infty)$ and that the square root function is concave, it follows from Jensen’s inequality that
+Toward this end, noting that $\kappa t\mathrm{e}^{-\kappa tx}\mathrm{d}x$ is a probability measure on $0,\infty)$ and that the square root function is concave, it follows from Jensen’s inequality that
 $t^{1/2}\int_{0}^{\infty}\mathrm{e}^{-\kappa tx}\mathbf{E}\big[\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}x}_{1})^{4}\big]^{1/2}~\mathrm{d}x\leq\frac{t^{1/2}}{\kappa t}\left(\int_{0}^{\infty}\kappa t\,\mathrm{e}^{-\kappa tx}\mathbf{E}\big[\mathfrak{L}^{0}_{1}(B^{t^{-1/2}x,t^{-1/2}x}_{1})^{4}\big]~\mathrm{d}x\right)^{1/2}.$
@@
 (A.1) $\mathbf{E}\big[H_{n}(\mathcal{B}^{\beta,n})\big]=-\tfrac{1}{n}\tfrac{\partial}{\partial\beta}\log Z_{\beta,n}\qquad\text{and}\qquad\mathbf{Var}\big[H_{n}(\mathcal{B}^{\beta,n})\big]=\tfrac{1}{n^{2}}\tfrac{\partial^{2}}{\partial^{2}\beta}\log Z_{\beta,n}.$
-In the special case where $V(x)=x^{2}/2$, a classical calculation using a Selberg integral [55] yields the explicit formula $Z_{\beta,n}=Z_{G\beta E,n}$, where we define
+In the special case where $V(x)=x^{2}/2$, a classical calculation using a Selberg integral [Selberg] yields the explicit formula $Z_{\beta,n}=Z_{G\beta E,n}$, where we define
 $Z_{G\beta E,n}=(2\pi)^{n/2}(n\beta/2)^{-\beta n^{2}/4+(\beta/4-1/2)n}\frac{\prod_{j=1}^{n}\Gamma(1+j\beta/2)}{\Gamma(1+\beta/2)^{n}}.$
 Asymptotics using several applications of the Mathematica ”Limit” function reveal that, as $n\to\infty$, one has
-$-\tfrac{1}{n}\tfrac{\partial}{\partial\beta}\log Z_{G\beta E,n}=\tfrac{3}{8}n-\tfrac{1}{2}\log n-\tfrac{1}{4}\left(1+\log(\beta^{2}/4)-2\tfrac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}\right)+o(1)$
+$\tfrac{1}{n}\tfrac{\partial}{\partial\beta}\log Z_{G\beta E,n}=\tfrac{3}{8}n-\tfrac{1}{2}\log n-\tfrac{1}{4}\left(1+\log(\beta^{2}/4)-2\tfrac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}\right)+o(1)$
 and $\tfrac{1}{n^{2}}\tfrac{\partial^{2}}{\partial^{2}\beta}\log Z_{G\beta E,n}=o(1)$.
 Combining this with (A.1), we get that when $V(x)=x^{2}/2$, the limit (2.4) holds.
-Using the well-known identity between the polygamma functions and the Hurwitz zeta function (e.g., [1, 6.4.10]), we have that for $\beta>0$,
+Using the well-known identity between the polygamma functions and the Hurwitz zeta function (e.g., [AbramowitzStegun, 6.4.10]), we have that for $\beta>0$,
 $\frac{\mathrm{d}}{\mathrm{d}\beta}\left(\log(\beta^{2}/4)-2\frac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}\right)=\frac{2}{\beta}-\sum_{k=0}^{\infty}\frac{1}{(k+1+\beta/2)^{2}}>\frac{2}{\beta}-\int_{-1}^{\infty}\frac{1}{(k+1+\beta/2)^{2}}~\mathrm{d}k=0,$